A Selected Bibliography of Publications by, and about, Leo Szilard

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

09 July 2021 Version 2.93

Title word cross-reference

+ [ABF^{+}43, CL18, LC02, Lan04c, Lan12b]. $1 [Duf46]. $20 [Per62b]. $27.50 [Cas93]. $35.00 [Bus93]. U^{238} [Tur46].

//www.store.aetv.com/= [Wac06].

1 [Szi54a]. 11th [KNSB54]. ’15 [Lan15]. 17th [Lan96b]. 1930 [Cho10]. 1930/41 [Fer68, Fer71]. 1933 [MWW11]. 1933-1980s [MWW11]. 1939 [Dan96, Szi63e]. 1942 [Ano11, CMS42, Lau46, MS42, MLC^{+}42]. 1945 [CD48, Dan95, Lan95a, Szi76b, Pal10]. 1952 [KNSB54]. 1963 [Szi87]. 1964 [E.01, Rab64]. 1969 [ZB69]. 1971 [Jas93]. 1979 [All80]. 1980s [MWW11, BB12]. 1981 [Fel81]. 1998 [Mar98b]. 1st [Ano87a].

20 [Cas93]. 2002 [Blo04]. 20th [Pri03]. 21st [Kel02]. 25 [Szi63e].

387 [Pei43].

60th [Lan05a].

737 [Fri74].

888-423-1212 [Wac06].

9 [Szi44c].

= [Wac06].

A&E [Wac06]. A-387 [Pei43]. A-Bomb [Ano61g, Ano64a, Ano64b, Wei61, Ano60b, Lan99b, Rob14c]. A-Scientist [Rob6x]. A. [Ano55a]. ABM [Kub98]. ABM-Vertrages [Kub98]. Abolishing [Weh61]. abrogate [Szi60e]. Absorption [AFS39, BJ41, Szi35a, Cha11, Sch00a]. Abstracts [ZB69]. Academic [BB12, MWW11]. Academy [Szi59b]. Acceleration [Szi28]. Accelerators [Tel00, Tel98a]. achievements [Gab74]. acid [NS54]. Action [Szi55d, Woo62, Szi41d]. Active [Szi39c]. activist [Bes85, Bes93, Win98]. acts [Ano61g]. Adaptation [Szi60c]. Addendum [Szi54a]. Address [Ano84a]. Administrators [STS^{+}49]. Adventure [ND98]. adventurous [ABES12]. AEC [Jas93, Szi49a, Szi63a]. Affairs [GR63]. After [Smi65a, Mar94b, Pod08]. again [Cso98a, Cso98b]. Against [Ano61g, Ano62a, Szi76b, Bes89, Wal97, Jor79]. Age [Ack82, Ano62a, GR63, Lan98i, Lea59, Szi60d, Wat93, Wei79, Ber87, EE72, Mar96, Boy85, Lau46]. Ageing [Szi59c]. agents [Pal05]. agents-clusters [Pal05]. aggregates [Szi42g]. Aging [ZBZ^{+}09, Szi55e, Szi59a, Szi63b]. Agnew [Ano12]. ago [Ano90]. Agony [Fen78]. Agreements [Szi55d]. ahead [Ano62e, Bro54]. Air [FS58a, FS61]. Air-cooled [FS61]. Al [OSY44]. Alarm [Rid84, Yor75]. Albert [Ein39, Nef11, Sch12a, SE45]. Alexander [LSW58]. Alfred [Bus93]. All-Out [Mab62, Oli62]. Allen [Zuc88]. Along [Ano62d]. Also [Kat75]. altri [Szi62d]. America [Cos84, FB69, Smi65b, Smi70, Szi63i, Tel02, Wac06, Jor79]. American [Kub98, Ano76, Bad05, Boy85, CHW91a, Cha99, Lan12a, LSW58, Szi61f, Val06]. americana [Val06, Szi60f]. Amerikanisch [Kub98]. Amerikanisch-Sowjetische [Kub98]. amino [NS54]. among [Lew94, Szi98b, Szi98a]. analyses [Gou96]. analysis [Cha11, Del98, Sch00a]. Anarchistic [Jor79]. Anatomy [Bus93]. ancestral [Szi53a]. Andrew [WS92, WS03]. angestammten [Szi53a]. Angolból [Wig92]. Animal [Szi60a]. Anniversary [Lan05a]. annual [KNSB54]. Anti [NS52]. Anti-mutagens [NS52]. Antibody [Szi60b, Szi64c]. anticarcinogenesis [vBDL96]. Anticipating [Gus12]. antimutagenesis [vBDL96]. Anwendung [Szi26]. any [Gwe63]. Anything [BD02]. Apparatus [Szi39a, Szi55e]. application [Szi26]. Applications [Szi34b, Szi78b]. approach [FGMvH14]. Approximate [WWCS42]. Area [Szi42e, DS57]. Arising [Szi34b, Szi78b]. Arms [Kub98, Lan96c, Lan97b, Lan09a, Pea89, Per62b, Per62a, Zuc88, Far13b, HGS87, Szi46a, Szi07]. Arms-Control [Pea89]. Arnyekban [Lan97d]. Art [Lan86a, Lan09c]. Arthur [Duf46]. article [Ano50a]. Articles [GR63]. Arvet [Tib98]. Ascent [Bro73a]. Asks [Ano61a]. assisted [CDS12]. Asynchronous [Szi34a]. Ateisten [Kle87]. atheist [Kle87, Kle90]. Atom [Ano55b, Ano61d, Ano62a, Duf46, HH74, Jas93, Mab62, Scu07b, See03, CHW91a, GBL05, Ber98]. atomeromutípusokról [Rós98]. Atomic [Ack82, Ano61a, Ano62b, Ano84b, Ano17, Aro50, Ber76, Ber03, Bla46, Can00, Cas93, Com56, Duf46, FR13b, Fie07, Fin60, GR63, Jen11, Kel04, Lan85, Lan94a, Lan99a, Nef11, PT02, Rei84, Ste75, Swi45, Szi45a, Szi47a, Szi47e, STS^{+}49, Szi50a, Szi60f, Szi72, Szi76b, Szi78a, Tur06, Tur46, Vol63, Wei79, WHF^{+}96, Yav78, Amr59, Amr60, EE72, Fei66, GS50, Gwe63, Lan01d, MW46a, MW07, O^{+}01, Rai02, Ree15, Szi85, Van03, Wal97, Boy85, Lau46]. Atomic-Test [Fin60]. atomica [Szi85]. Atomphysiker [Her76]. Atomreaktor [Lan01d]. Áttekintés [Rós98]. Attempt [Szi45b]. Attempts [Lan96d, Yor75]. Attention [Ken54]. átvette [PM98]. audience [Ano12]. Auffindung [MS25]. August [Dan96, Lan95a, MLC^{+}42, Szi87]. Auschwitz [Pri03, Pri03]. Ausdehnung [Szi25]. Auslösung [BLW^{+}34a]. Author [Lan93c, Lan93b, Lan93a]. authors [SH57, SHG58]. Autoren [Ano02]. Autorinnen [Ano02]. Avert [Mab62, Oli62, Szi46a, Szi07]. Avoid [Rob6x]. Award [Lan93h, Nef97, Bak98, Jea09, Kel02, Pod08]. Awarded [Smi66].

B [Szi64d]. Background [Ano85, Ano95, Szi72]. backward [KK11]. Bacteria [NS49, NS50b, NS51a, Szi60a, Szi62h, NS54]. Bacterial [NS51a, FS55b]. Baiting [LHN02, Lan04a]. Ban [Fin60, Lan90]. Baptism [Ste75]. Barak [Wac06]. Barbarei [Pri03]. barbarism [Pri03]. Bares [Ano62b]. barriers [GBL05]. Bars [Hig60]. based [Pal05, Szi41d]. Basis [Szi60b, Szi64f]. bathtub [Low93]. Bauhaus [Pri03]. be [Gwe63, Sch15, Szi41d, ABF^{+}43]. beam [Szi29d]. beams [RS31]. became [Ree15]. Bee [Szi65]. Beeinflussung [RS31]. Been [BD02]. Before [Rab46]. Beginning [HA62, MCW47, Yav78]. Beginnings [Haw13]. behalf [BL81]. Behind [Bet93, Hawxx, Her93, Lan93d, Lan93h, Lan96c, Lan13, Lan14a, Lew94, Per93, Ter93, Bad94, Hew94a, Hew94b, LS92, LS94, Lan97a]. being [Gra94, Gra96]. Beings [Szi64h, Szi29a]. Bekecsinalo [Lan97c]. Bela [Bad94, Bet93, Han93, Low93, Per93, Sch93b]. Beller [Mar98a]. Berlin [Szi60g]. Bernard [Fri74, Gab73, Gol74, Wei74, Ano93a]. Beryllium [BLW^{+}34a, BLW^{+}34b, SC34b, Szi41a, ABF^{+}47, MS41b]. Beschleunigung [Szi28]. beszélgetés [PM98]. Better [Zuc88]. Between [Lan00b, CGS35, Gou96, Sen21, SE45, WWCS42, SH57, SHG58, Val06]. Beyond [EN99]. Bi [BFAS47, SBFA47, SBFA48]. bibliography [Inf70]. Biochemical [Ank60]. Biographical [Liv60, Wig69]. Biographie [Nef05, Nef06]. Biography [Bad94, Bet93, Her93, Hew94a, Hew94b, Lan88, Lan95b, Lan96f, Lan13, Lew94, Nef05, Nef06, PL88, Per93, Sch93b, Sch93a, Ter93, LS92, LS94, Nef07]. biológiában [ND98]. Biological [Ano64d]. Biologist [Lan14b]. Biology [BD02, DS57, Lan92b, CSW92, ND98]. Birth [Anoxx, Lau46, Mar98b, Szi60d, Wat93]. bis [Ano02]. bismuth [FS42a, Szi42f]. bodies [Szi41d]. Bohr [Duf46, BS50, Sch00b]. Bomb [AB96, Ano60a, Ano61b, Ano61d, Ano61g, Ano64a, Ano64b, Ano84b, Arn50b, Aro50, Ber76, BBSS50, Bet93, Boy85, Duf46, Fen78, Fie07, Gol92c, HH74, Hawxx, Her93, Kel04, Ken54, Lan86b, Lan90, Lan93d, Lan93g, Lan95c, Lan95d, Lan95e, Lan96d, Lan96e, Lan98g, Lan99a, Lan99d, Lan00a, Lan01b, Lan01e, Lan03, Lan09a, Lan11c, Lan12c, Lan13, Lan14a, Log93, MD67, Nef11, Rab60, Rei84, Swi45, Szi47e, STS^{+}49, Szi60f, Szi63d, Szi76b, Ter93, Wei61, WHF^{+}96, Wol67, Lan15, Amr59, Amr60, Ano60b, Bul79, Eli50, Far13a, Far13b, Fei66, FGMvH14, GS50, Gwe63, Lan89a, LS92, LS94, Lan97a, Lan99b, MW46a, MW07, O^{+}01, Pow93, Rai02, Ree15, Rob14c, Sch15, Szi85, Van03, Wei98, Lan01c, Bad94, Ber93, Bet93, Duf46, Hew94a]. Bomb [Hew94b, Per93, Swe93, Cas93, Lew94]. bomba [Lan01e, Szi85, Wei98]. Bombarded [AFH39, SC34a]. Bombenabwurfs [Lan03]. Bombes [Lan86b]. Bombowa [Lan01c]. Bombs [Can00, Szi47a, Szi78a, Szi45a, Wal97]. Book [Bad94, BB12, Ber93, Bet93, Bus93, Cas93, Fri74, Gol74, Han93, Hew94a, Hew94b, Jas93, Lau94, Lew94, Low93, Mal08, McC65, MCW47, Pal10, Sch93b, Sch93a, Swe93, Wei74]. Books [Ano85, Ano95, Gab73, Hut61]. Booksigning [Lan93b]. boron [ABF^{+}43, ABF^{+}47]. boundary [WWCS42]. Box [Wac06]. Branch [Szi34b, Szi78b]. Brasilia [Pri03, Pri03]. Brass [LHN02, Lan04a]. breaker [Log93]. breaking [Par01]. Breeder [Szi45c]. brief [Tel98b]. Briggs [Szi39b]. Brillouin [Gou96]. bring [Yor75]. Brings [Lan64]. Britain [Far13b, Rid84]. broadcast [Swi45]. Broadway [Sch00b]. bubble [Cha11]. Budapest [Mar98b, Rad98, Sza87, Sza98]. budapesti [Sza98]. Build [Bet93, Gol92c, Lan90, Lan99a]. Builder [Wei61]. Builders [MD67]. Building [Lan09a]. Bulletin [Ano93a, GR63, Ano90]. Bumblebee [Per93]. Bumbling [Lan89a]. Burlington [Wac06]. Bush [Gol92c].

C [Swe93]. C. [Ano55a]. Caffeine [Szi52a]. calculations [FS42b]. Called [Szi60f, Szi85]. Calling [Szi47b]. Came [Ano95, Rei95, Sch15]. Campaign [Ano61a, Las72]. campaigned [Bes89]. Campus [Woo62]. Can [Szi46a, Szi50a, Szi07, LTD02, Szi41d]. Canfield [DS57]. cans [OSY44]. Cape [Cas93]. Capital [Fin60, Weh61]. Capture [GS37, Szi41a, MS41b]. CARA [Lan08b, Lan11e]. Carbon [Szi41c, Szi46b]. Career [Lan89d]. Case [Gru83, Ken54, Szi72]. Cash [Lan64]. Cass [DS57]. Cassidy [Swe93]. casting [Szi26]. Cause [See95]. caused [MS41a]. CBE [Blo04]. Cells [BD02, MSHP06, Szi60a, Szi62h]. Cent [Smi62]. Centenary [Mar98b]. Central [Ano52, Szi49c, Szi82, Szi03]. Century [Lan99a, Har06, Puc60, Har10, Kel02, Pri03]. Ceremony [Fin60]. CF338 [Pei43]. CH [Blo04]. Chain [FS41, FS58b, Szi34c, Szi41b, Szi41c, Szi79, And73, Ano11, Mar94b, SF44, Szi41d, Szi46b]. chain-reacting [Szi41d]. Challenge [Lau94, Bro54]. Chalmers [Ber98]. Chalmers-jelenség [Ber98]. Chance [Hig60, Ber87]. changé [VB15]. Changed [Wac06, Har06, Mar06, VB15]. changer [Ree15]. Changes [Oli62]. Changing [Jas93]. Channel [Wac06]. chapters [HA62]. Charter [Arn55]. chat [PM98]. Chemical [SC34a, Szi35b, Szi36]. chemically [NS51a]. Chemistry [Sch15, Ber98]. Chemostat [NS50a, NS50b, NS51a, NS54]. chess [Kle98]. chiama [Szi85]. Chicago [FHN^{+}45, CD48, Pri95]. Chromatin [Szi62h]. Churchill [Far13a, Far13b]. Circulated [Szi76b]. Circumstances [Lan99d, Lan03]. Cities [Szi61h]. Citizen [Lan86a, Zuc88]. city [Kle87, Kle90]. Civilian [Lan09b]. classical [MP10]. Climate [Gol92c]. Close [Ano61c]. Close-up [Ano61c]. closer [Har10]. closures [OSY44]. cloud [Bes93, Win98]. clusters [Pal05]. Cobalt [Arn50b]. Cold [Nef97, ZB69, Pod08, Sen21, AB96, Bes89, Kub98, Lan08a]. Collected [Fri74, Gab73, Gol74, Wei74, FS87, Wig96b]. College [KNSB54]. collision [SZ41]. Collisions [CGS35]. Columbia [Fer55, Szi41d]. Come [STS^{+}49, Szi63i, Szi60e]. Comic [Lan98c, Lan98d, Lan98b, Lan98j]. command [PM98]. Commandments [Bro78]. commemoration [Tel98b]. Comment [Szi47c]. Commentary [BL81]. Comments [KG12a, Pei43]. Committee [ACF^{+}42, Szi59b, FHN^{+}45]. Communications [Szi71, Zei70]. Communicators [Kub98]. Community [Hab69, Jor79]. companion [Cha99]. comparison [ABF^{+}43, ABF^{+}47]. Compensation [Szi62h]. Competition [Ano98a]. Competitive [Szi63b]. composed [Szi46b]. Compton [Duf46, Szi42a]. computational [Pal05]. Concept [MSHP06]. Concern [Fin60]. concerning [Bro54]. Condition [FS41, Bro54]. conditions [FS55b, WWCS42]. Conference [Mar98b, Osb97, Ano57]. Conferences [Kub98]. conflict [BO76]. Confronts [Ano76]. Connected [Nag81]. Connections [Lan93i, Lan99e]. Conscience [Ano98b, Cof64, Lan98h, Viz98]. consciousness [Scu07b, Szi85]. Consequences [Moo92]. considerations [LZL^{+}12]. conspiracies [Cas93]. Consultant [Lan93d]. Containing [Szi52a, Jas93]. Contribution [Szi41c, SW98]. contributions [Mar98b]. Control [KG12b, Lan09a, Lan09b, Pea89, Szi41b, Szi50a, Szi60a, HGS87, KG12a, Kub98, Zuc88]. controversy [BL96]. conversation [Wol67]. Conversations [Szi45b]. Cooled [FS58a, Szi45c, ACF^{+}42, FS61, LCM^{+}42, Szi42f]. Cooling [MP10, Szi42d, FS42b, Szi42b, Szi42c]. Cooperation [Lan01b, Lan01c, PT02]. Copenhagen [Sch00b]. copier [Sch12b]. core [Szi44c]. corpuscles [Szi28]. Corpuscular [Szi29d]. Correction [Ano50a, Ano94]. Correspondence [BS50, Sta99, SE45, SF44, SH57, SHG58, SK63, SWS80, WS78]. coscienza [Szi85]. Cosmic [Lan98c, Lan98d, Lan98b, Lan98j]. Cost [Arn50a, GBL05, Kra21]. Council [Oli62, Weh61, Lan64, Szi62b]. countering [Kle05b]. Couple [Lan00a, Lan01e]. Craters [Szi71]. create [DS57]. creation [Lan12a, Kle05a]. Creative [Szi72]. Criminal [Szi49b]. Crisis [Szi60g]. Critical [Ber76, FS41]. criticism [Szi50c]. Critics [Jas93]. Crusade [Pea89, Szi47b, Zuc88, HGS87]. crystals [MS26]. családfája [Cze98]. csille [ÉK98]. Cuba [DFF^{+}63]. Culture [Boy85, Lap68, Lap69]. Current [BD02]. currents [Szi26]. Curriculum [Szi64a]. cycle [Del98]. Cycles [Szi46d, CDS12]. Cyclotron [Man53].

D [Mal08]. dagens [Tib98]. Daniel [Jor79]. Danilo [Bes89, Bes93, Win98]. Danish [MW46b, Szi64b]. Dannen [Ano76]. dans [Dem15]. daughter [PM98]. dauphins [Szi62e]. David [Swe93]. Davis [Szi50b]. Dawn [Ack82, Boy85]. day [HHW99]. Days [Lan08b, Lan11e, Wac06]. Death [Lan05b, Kle98]. Debate [Ano46, ST60]. Dec. [Lau46]. December [Ano11]. decipherability [Man65]. Decision [Gol92c, Amr59, Amr60]. decisions [Str63]. Decoration [Gro46]. Decrease [Szi64h]. deep [KK12]. defeat [Ols63]. Defence [BB12, MWW11]. Defense [Zuc88]. Define [Szi45b]. Delano [Ein39]. delfinek [Szi05]. Delfinernes [Szi64b]. delfini [Szi62d, Szi04]. delfinjei [Her98]. delights [Ano12]. Delphine [Szi63c, Szi81]. Demon [Ehr67, KG12b, Ber90, CDS12, Tam98, KG12a, Kra21, Scu07a, VJ11, Dev02, Dev03, EN98, EN99, Lef02, Rex17, Scu07b, Zur03]. démona [Tam98]. Demons [Ben87]. denial [BL96]. Dennis [All80]. Denver [Ano61b]. Dependence [Szi60d]. Description [NS50a]. Design [ACF^{+}42, Del98]. designated [DS57]. designed [Dan97b]. Desirable [ST60]. desperately [Gwe63]. destruction [Wal97]. Detection [SC34b, Szi34c, MS25]. detectors [FSS46, FSS47]. deterrence [Fel78]. Deterrent [Szi64e]. deuterium [FS50]. Develop [WHF^{+}96]. Developer [Ano61d]. development [Gab57]. device [FS55b]. Dexer [Duf46]. Dialogue [Lan00b, Zsa76]. Diary [Szi50b]. Did [STS^{+}49, Jen11, Ols63]. Didn’t [Wei94]. Dies [Ano64a, Ano64b]. difference [Sen21, Hew94a]. Different [NS51b]. diffusion [WWCS42]. digitize [Rob14d]. díj [Bak98]. díjas [Zsa76]. Dilemma [Szi62b, Fel78]. Dimensions [Lan00b]. Diplomacy [Lan86a]. Diplons [CGS35]. Directed [Wac06]. Disarmament [Ano61a, See95, Szi55a, ST60, Szi60i, Szi61g, FGMvH14, Pod08, Fow55]. disarmed [Szi61a]. Discharge [Szi29b, Szi29c]. Disclose [Fin60]. Discovery [BHT86, CHW91b, CHW91a]. Discussion [LCM^{+}42]. Disputes [Ken54]. Dissent [Pri95]. Divergent [Szi46b]. Divine [Lau94]. division [CMS42, MS42]. divisions [MLC^{+}42]. do [Pol19, Szi52b, Gwe63]. Doan [Szi42e]. documentary [CHW91a]. Dolci [Bes89, Bes93, Win98]. Dolphin [Cle08, Her98]. Dolphins [Szi60c, Szi62d, Szi62e, Szi63c, Szi81, Szi04, Szi61j, Szi64b, Szi92, Szi05, Szi60c, Hut61, Far92, McC65]. Dominance [Szi62g, Szi63f]. Doomsday [Smi07, Mal08]. Dosage [Szi62h]. dossier [Szi85]. Double [Fra09, Pal10]. Down [Lan90]. Dr [Ano64d, Mal08]. Dr. [Ano84a, Chi60, Gro46, Irw61, Ken54, Mar48, Smi60, Smi07, Szi42e, Szi47c, Szi50b]. Draft [Szi87, Szi41c]. Drama [Her76, Ken54]. Dream [Lan83, Mal08, Bal05, Smi07]. dreamers [Poh62, VB15]. Drive [Lew61]. drop [FS42b]. Dropped [Lan95c, Lan95e, Lan96e, Lan12c, Lan15]. due [Wei93]. Duo [Ano76]. durch [BLW^{+}34a, MS26, RS31]. during [Bro54]. Dynamic [Ano76]. Dynamics [AB96, KNSB54].

Early [Boy85, Mar68, Mar94a]. Economy [Szi46d]. Ecstasy [Fen78]. Edited [Duf46, BB12, Gab73, Jor79, Wei74, Zuc88]. Editor [Fow55, Moo92, Szi55c, DFF^{+}63, Har65, Szi60g, Szi61f, Szi61c, Szi61d, Szi61e]. Editorial [Ano66]. Editors [Bro85, Mar48, Szi47c, Szi61f]. eds [Fri74]. Edward [Zuc88, BO76, Har10, Hor97, Szi87]. Effect [Hew94a, SC34a, Ber98, LL12, MS25]. Effektes [MS25]. Efforts [Lan90]. einem [Szi29a]. einfacher [MS25]. Eingriffen [Szi29a]. Einstein [Ano76, Wac06, Cha11, Cho10, Dan96, Dan97a, Dan97b, Del98, Ein39, Lan02a, Lan05b, Lan06a, Lan06b, Lan11b, Lan11a, Nef05, Nef06, Nef07, Nef11, PB79, RTS^{+}65, RTS^{+}69, Sch12a, Sen21, Sta99, SE45]. Eisenhower [Lau94]. Electrical [KG12b, Szi26]. Electrodynamic [SE28]. electrographite [BJ41]. Electron [BLW^{+}34b, Haw13, Gab57, Gab74, Mar68, Mar94a, RS31]. elektrischer [Szi26]. Elektrographit [BJ41]. Elektronenmikroskops [Gab57]. Elektronenstrahlen [RS31]. Element [SC34a, Szi58b, WSC59, Szi41f]. Elemente [BLW^{+}34a]. Elements [Szi35b, Szi36, Szi39c, SWC59, BLW^{+}34a, SZ41]. életmuve [Csi98]. elimination [Szi44a]. Elusive [Smi60]. Embellishments [Kra93]. Emerging [Gus12]. Emigration [Rid84]. Emission [SZ39, ZS39, SFA^{+}43]. emlékeim [Sil98]. emléktáblájának [Cso98a, Cso98b]. encounters [Kle87, Kle90]. End [Ano61a, MCW47, Per62b, Per62a, Yav78, Fei66, OSY44]. ending [CMS42, MS42, MLC^{+}42]. endorsed [All01]. Energy [BO76, Jen11, KG12a, Lan85, Seg85, Szi34c, Szi41a, Szi50a, Tur06, Tur46, Wei79, Fer55, GBL05, Lau46, MS41b]. Engine [Bri56b, Bri62b, Bri04b, Bri13b, Dev03, KSDU11, KG12b, Scu07b, BS95, CDS12, GBL05, HS98, KG12a, LZL^{+}12, Mag96, MP10, Par01, VJ11, Dev02, Zur03]. Engineering [MLC^{+}42]. Engines [Ben87, KK11, KK12, LL12]. England [Bes89, Bes93, Win98]. English [Wig92]. Enrico [Seg70, Tel02, Tur06]. Entropieverminderung [Szi29a]. Entropy [Lub87, Par01, Szi64h, BS95, Gou96, Scu07b, Szi29a, Szi76a]. Entwicklungsgeschichte [Gab57]. Environment [Jas93]. Enzyme [Szi60h, Szi62h]. èOsztöndíj [CST01, BST03, BSST04, CST02]. epic [Bou13]. epiphany [Far13c, Far13d]. epistemological [Gou96]. Equality [Szi50d]. Equation [Lan05b, WWCS42]. Erasmus [Mag98]. erasure [HS98, Kra21]. Erinnerungen [Szi54a, Szi53a]. Ernest [Danxx]. erste [Lan01d]. Erzählungen [Szi63c, Szi81]. Érzékeny [Szi98b]. Erzeugung [BLW^{+}34a]. escape [Mar06]. essäer [Kle05a]. Essay [Bus93]. essays [Kle05a]. est [Dem15]. establishment [Ols63]. estinto [Szi03]. Eszter [Wig92, Wig92]. ethical [Tib98]. Ethics [Lan11c]. etiska [Tib98]. Eugene [Hor97, Zsa76, WS92, Wig96b, WS03, Wol67]. Europe [Aro50, Fer68, Fer71, FB69, Szi61i]. European [Lan12a]. Eva [PM98]. Évával [PM98]. Even [Wei94]. Events [Szi60i]. Ever [Fra05]. Evolution [Lan05d, Lan06c, Pal05]. Examination [ACF^{+}42]. example [All01]. Examples [FS42b]. Excerpts [Ano79a, Ano79b]. Excitation [GHS39]. excited [GS37]. exile [Fra09, Pal10]. Exiles [Fra05]. Exorcist [EN98, EN99]. Expected [STS^{+}49]. Expelled [MP01]. Experiences [Cos84, Gou96]. Experiment [Szi39b, Lau46]. Experiments [NS49, NS50b, NS51a, Szi34c, NS54]. Expert [Ano62d, Ano62c, Mab62, Poh62]. explains [Sen21]. Explosive [Wal04, Szi41d]. expulsion [Szi41d]. extension [Szi25]. extinct [Szi03]. extrana [Lan01e]. Extraordinary [Gib19]. Extrusion [Szi44a].

Face [Szi49d]. Faces [Rot73]. Factal [Rip98]. Factor [MB00]. Facts [Ano79c, Ano79d, BBSS50, Hug80, Phi80, Sar82, Szi49d, Ano79a, Ano79b, SWS80, WS78, Jor79]. Faculty [Lan93e]. Failed [WHF^{+}96]. Fallout [O^{+}01, Rai02]. Fame [Rip96]. Familie [Szi53a]. family [Cze98, Szi53a]. Farm [Ber93]. Fashioned [Duf46]. Fast [FS41, SZ39, Szi41c, Szi45c, SBFA48, BFS47, FSS46, FSS47]. Father [Ano62a, HH74, Szi60d, Rob14c]. Fe [BFAS47, SBFA47, SBFA48]. Fearing [Lan09a]. Features [BC12]. Februar [Szi54a]. February [All80, Blo04, Far13c, Far13d, Mar98b, Szi54a]. Feeder [MSHP06]. felavatása [Cso98a, Cso98b]. Feld [Fri74, Gab73, Gol74, Wei74, Ano93a, Pei43]. felelossége [NS98]. Fellowships [Szi49a, Szi63a]. Ferdinand [Blo04]. Fermi [Tel02, And74b, And74a, Ano55a, Bye02, Bye04, Cro04, Lan92a, Lan01d, Seg70, SC34a, Tur06, Wei94]. Feyerabend [Jor79, Jor79]. Fiction [Cle08, Far92, AJ82, Szi81]. fields [RS31]. Fight [Szi49a, Szi63a]. File [Les94, McM94, Szi85]. Film [MCW47]. First [Ano55a, Ano11, Bet93, Can00, Ken54, Szi55b, Wei61, Far13b, Lan01d, Lau46]. fissile [FGMvH14]. Fission [BHT86, CHW91b, CHW91a, MS41a, SFA^{+}43]. Five [Wol67, Har06]. Fizikaverseny [Ano98a]. Fizikus [Lan97c]. fled [Mar06]. Flimsy [McM94]. flow [FS42b]. fluctuation [Szi22]. Fluid [Nag81, SE28]. Fly [Szi62h]. Focus [Wal04]. Focuses [Ken54]. Foray [Far92]. ford [Wig92]. Foreign [Vol63]. Foreword [Duf46, Per93]. forgotten [Sch12b]. Form [Weh61]. Formation [Szi60a, Szi60b, Szi64c]. Formen [Szi26]. Formula [Rob6x]. Formulate [Szi55d]. forróatom [Ber98]. forróatom-kémia [Ber98]. forskning [Tib98]. Forum [Lan15]. forward [KK11]. foundation [Kub98, Szi62f]. Founded [Lan64]. founders [Bou13]. Founding [Lan08b, Lan11e]. Four [Ano79d, Bes89, Fin60, Lan93i, Lan99e, Win98, Bes93, VB15, Lan11d]. fra [Val06]. från [Tib98]. France [Bes89, Bes93, Win98]. Franck [Rab64, Pri95]. Frank [Pal10]. Franklin [Ein39]. Frederick [Gus12]. Free [Jor79, Szi46d]. French [Dem15, PT02, Szi62e, VB15]. Frequency [SG55]. Fridge [Lan11b, Sen21]. Friedrich [ABES12]. friends [RTS^{+}65, RTS^{+}69]. Fruit [Szi62h]. Fuel [Szi58b, SWC59, WSC59]. Full [Duf46, Weh61, MW46a, MW07]. Full-Time [Weh61]. fullkomlighet [Kle05a]. fully [LZL^{+}12]. Fundamental [DS57, Szi72]. funny [AJ82]. Future [Ano95, Bla46, Rei95, Szi39b, Bro54, Bro85, BM98, GC02, Wei98].

G [See03, Zuc88]. Gable [Szi62f]. Gabor [All80]. Gaia [Mar98c]. Galactosidase [Szi64d]. gallery [SE97]. game [Ree15]. game-changer [Ree15]. Gamma [GS37, SC34b]. Gebote [Bro78]. Gene [Ano76]. General [Ber03, DS57]. generation [LM97]. Genesis [Bou13, MSHP06, Fer55]. Genetic [NS51a]. génies [VB15]. Geniĭ [Lan05c]. Genius [Her93, LS92, Lan93d, Lan93e, Lan93f, LS94, Lan95b, Lan98i, Lan13, Lew94, Ros64, Ter93, Lan97d, Bad94, Bet93, Hew94b, Lan05c, Han93, Hew94a, Lew94, Low93, Per93, Sch93b, Sch93a]. Geniuses [Lan93g, VB15]. Genotype [NS51b]. geopolitical [Ree15]. Georg [ABES12]. Gerda [Dan15]. German [Ber93, Cas93, Swe93, BJ41, BLW^{+}34a, Bro78, Gab57, Her76, Kub98, Lan86b, Lan01d, Lan03, MS25, MS26, Nef05, Nef06, Ols63, Pow93, Pri03, RS31, Szi22, Szi25, Szi26, Szi28, Szi29a, Szi29d, Szi53a, Szi54a, Szi63c, Szi81]. Germany [Pal10, AB96, Fra09, WHF^{+}96]. Gertrud [Gab73, Gol74, Jor79, Zuc88, Fel81, Hawxx]. Gertrude [Fri74, Wei74]. geschrieben [Szi54a]. Gets [Ano55a, Dan97b]. Giants [MD67]. Gießen [Szi26]. Gift [MP01]. Give [BFM^{+}49]. Giving [Ber87, Wei93]. global [Ree15]. Go [Fra05]. Goal [Ano62c, Per62b]. goddess [Mar98c]. golden [Mar96]. Goldhaber [Lan01a]. Goodman [Wac06]. Gorky [Sak83]. got [Ols63]. Grand [Szi03, Szi49c, Szi82, Szi03, Ano52]. grant [Rob14b]. granted [Cho10]. Graphite [SWC59]. Great [Ano46, Amr59, Amr60, Lau46, Mar06, Ols63, Sch12b]. greatness [Rob14a]. Greb [Zuc88]. Grid [Lan86b]. Grief [Szi63i]. grips [Yor75]. Ground [Ken54]. Group [Kub98, Szi55d, CMS42]. Groupp [Weh61]. Groves [Ber03, Gol92a, Gol92b]. growing [FS55b, NS51a]. growth [KNSB54]. Grunde [Lan03]. Grundlegung [Kub98].

H [Duf46, Szi42a, Ano60a, Eli50, Ken54]. H-Bomb [Ano60a, Ken54]. H. [See03]. halandó [Mar98c]. Hall [Rip96, Ber93]. halted [Dan01]. Handbuch [Ano02]. hangja [Szi05]. Harbor [ZB69]. hard [BLW^{+}34a]. harmful [Ano55c]. Harold [Ano12]. harte [BLW^{+}34a]. hatása [Csi98]. Hats [LHN02, Lan04a]. Hawkins [Zuc88]. Hazatér [Mag98]. hazatért [Mar98d]. He-cooled [ACF^{+}42]. heady [Scu07b]. health [DS57, DS57]. heart [Mil05]. Heat [Bri56b, Bri62b, Bri04b, Bri13b, KK12, Kra21, Mag96, Sch00a]. Heavy [Szi58a, SZ41]. Heisenberg [Sch00b, Kra93, Pow93, Swe93, Wei93, Ber93, Cas93, Swe93]. held [Ano55c, KNSB54]. Helen [Zuc88]. heliga [Kle87]. helium [FS42b, LCM^{+}42]. helped [Gwe63]. Her [Mur73, PM98]. Here [Wei61]. Herewith [Ols63]. Heritage [Pri03, Tib98]. Herkunft [Ano02]. hidden [Far13a]. high [Kra21]. Highlights [BD02]. Hilberry [Szi42e]. Hintergrunde [Lan03]. Hiroshima [Pri03, BL96, Lan95c, Lan99d, Lan03, Lan05a, Pri03, Rab46, Smi65a, Szi85]. histoire [VB15]. Historian [Lan02b]. Historical [BD02, Gou96, Rex17, Rip98, Seg85]. History [Ano11, Ber93, Cho10, Gab74, Jor79, Swe93, Swi45, Szi47e, Wac06, WP85, Amr59, Amr60, BL96, CHW91a, Cas93, Cha99, ÉK98, Far13a, Gab57, Mar68, Mar94a, Pow93, VB15]. Hitler [Mar06, MP01]. Holds [Ano55a]. Hollywood [Lan04c, Yav78]. holy [Kle87, Kle90]. Home [Wac06, Cso98a, Cso98b, Mag98, Mar98d]. homogeneous [AACZ09]. Homologous [Szi62c]. Honoring [Fin60, Kon98]. Hope [Lan98f, Lan99c, Lan08b, Lan11e, Smi65b, Smi70]. Hot [Rob6x, Ber98, Sen21]. Hours [Tel83]. House [Duf46, Las72, Szi45b]. Houtermans [ABES12]. hozzájárulása [SW98]. Human [Ano61e, Duf46, Yor75]. Humanities [Lan00b]. Hungarian [Mar94b, Pal10, Ano98a, Bak98, BM98, Csi98, Cso98a, Cso98b, Cze98, ÉK98, Fra09, Tam98, Kle98, Her98, Jon10, Ber98, Kon98, Lan93g, Lan97c, Lan97d, Lan98j, Mag98, Mar96, Mar98c, Mar98d, NS98, ND98, PM98, Rad98, Rós98, Rot98, SW98, Sil98, Sza98, Szi98b, Szi98c, Szi05, Tel98b, Wei98, Wig92, Zsa76]. Hungarians [Lan11d]. Hungary [ÉK98]. hur [Tib98]. Hydrogen [Ano50b, Arn50b, BBSS50, Ken54]. Hypothesis [ZBZ^{+}09]. Hysteria [Ano50b].

iced [Dan97b]. Ideas [Lan92a, Pla58, Gab74, Hor97, ND98]. Identical [NS51b]. If [Ano62c]. II [Wig96b, Ano79b, Bad05, EN99, Fei66, Lan93b, NS54]. Illustrious [Fer68, Fer71]. I’m [Ano61f, Kub98, Szi54a, Ano61c]. imaginary [Gou96]. Immediate [Szi39b]. immigrants [Fer68, Fer71]. Impact [Cos84, Csi98]. Implement [Mar48]. Implications [BS95]. Important [BD02]. Impossible [Lan09c]. Improved [Haf96]. Improvements [Szi35b, Szi36]. Inactivated [NS49]. including [Szi64a]. Income [Ano61a]. Induced [BLW^{+}34b, GHS39, SC35, NS51a]. Inducing [SC34b]. Inducted [Rip96]. Induction [Szi61b, Szi64d]. Industrial [Szi34b, Szi78b]. Inelastic [BFS47, BFAS47, SBFA47, SBFA48, SZ41]. Inevitable [Ano62c]. Infeld [Inf70, Sta99]. Influence [Eck64, Sal64, RS31]. influential [Har10]. információ [Tam98]. informal [SK63]. Information [Bri56a, Bri62a, Bri04a, Bri13a, Dev02, Dev03, KK11, BS95, Tam98, Gou96, HS98, KK12, Scu07b]. Informed [Bri56b, Bri62b, Bri04b, Bri13b]. ingen [MW46b]. initiating [FS50]. Inquiry [Gib19, Bro54]. inspection [Szi46a, Szi07]. Inspektionssystem [Szi46c]. Instantaneous [SZ39]. Institute [Ano64d, DS57, Bou13, DS57]. institutes [DS57]. Intellectual [Per93, Fer68, Fer71, FB69]. Intellectuals [Ano61g, Bes89, Bes93, SE97, Win98]. Intelligence [Szi72]. Intelligent [Szi64h, Szi29a, Szi76a]. intelligenter [Szi29a]. Interaction [SZ39]. interdependent [DS57]. Interdisciplinarity [Mar98a]. interface [WWCS42]. interference [Szi76a]. Interment [Lan98e]. Intermittently [Szi57b]. International [Jea09, Szi50a]. Intervention [Szi64h, Szi29a]. Interview [PL88]. Interviewer [Lan93d]. Introduction [Duf46, Rab60]. Invades [Szi47f]. invent [Gwe63]. invented [Sch12b]. Inventing [Gol92c]. Invention [Ano55a]. Inventions [Ano55b, Dan01, Sch12b]. Inventor [Ano84b, Dan98, Tel98a, Tel00, Rip96]. Inverse [Gro96]. Ionizing [Szi61b]. Irradiating [BD02]. Irradiator [WSCF61]. iskolái [Rad98]. Isotope [SC34a]. isotopes [Lan01a]. Israel [PT02]. Israeli [PT02]. Issued [Ano55a]. Issues [Ber76]. istenno [Mar98c]. Italian [Szi62d, Szi85, Szi03, Szi04, Val06]. Italy [Bes89, Bes93, Win98].

J [Jas93, Jor79, Szi39b]. J. [O^{+}01, Rai02]. Jacketed [SWC59, WSC59]. Jahre [Szi54a]. Jahrhundert [Ano02]. Jahrhunderts [Pri03]. James [Rab64, Sch12b]. Jan [Ano76]. January [Szi63e]. Japan [Ano60b, Szi76b, Wal97]. jelenség [Ber98]. Jeno [Zsa76]. Jewish [Pal10, Fra09]. Jews [Mar06]. John [Wac06, Zuc88, Hor97]. Johnson [KG12b]. Joliot [BL81]. Jonas [Per93]. Joseph [Sza87]. Journal [Vol63]. Journey [FR13b]. jövo [BM98]. jövoje [Wei98]. József [Sza98]. Judging [Har10]. jüdischer [Ano02]. July [CD48, Dan95, Lan96b, Szi76b]. June [All80, KNSB54]. Just [Cay62].

Kalandozás [ND98]. Kalinga [Smi66]. Kalten [Kub98]. Kan [Szi46c]. Kaprustning [Szi46c]. Kármán [Hor97]. Kathlerine [Duf46]. keepers [Van03]. Keeping [Lub87]. kémia [Ber98]. Kennedy [Szi61f]. Kettenreaktion [Her76]. Kevles [Jor79]. Key [Ken54]. kisebbség [Szi98b]. Knopf [Bus93, Cas93]. Knopf/Cape [Cas93]. Knowledge [Jor79]. Komikus [Lan98j]. Kommunikatoren [Kub98]. Konferenzen [Kub98]. Korpuskeln [Szi28]. Korpuskularstrahlrohre [Szi29d]. kozmikus [Lan98j]. Krieg [Kub98]. Kristallen [MS26]. különbözo [Rós98].

L [Szi63e]. Lab [Pri95]. Laboratory [FHN^{+}45, Lan93i, Lan99e]. Lactose [ZB69]. Landauer [EN99]. Landed [Lan93i, Lan99e]. Landmarks [Wal04]. Lanouette [Bad94, Bet93, Han93, Hew94a, Hew94b, Low93, Per93, Sch93b, Lan88, Lan93a, PL88]. lányával [PM98]. Large [Szi39b, Szi55e]. Large-Scale [Szi39b]. last [ÉK98]. Later [Ano61i]. Laughing [AJ82]. Laura [BB12]. laureate [Zsa76]. Law [Bus93, Ben87, Lef02]. Lawrence [Danxx]. leaders [Ols63]. Leakage [WSCF61]. Learned [BD02]. Learning [BB12, MWW11]. leaving [Szi60c]. Leben [Szi53a, Szi54a]. Lecture [Lan93b, Haf96, Mar98a, Nef97]. lectures [Mar98b]. Lectureship [Jea09, Kel02, Pod08]. Left [Cay62]. Legacy [And74b, Kel04, Lan08a, BM98]. Legend [Irw61]. Leo [Rot48, Ano64d, Ano84a, Ano98a, Bad94, Bak98, Bes93, Bet93, Bro78, Fri74, Gab73, Gol74, Gro46, Hew94a, Hew94b, Hut61, Kon98, Lan96a, Lan98d, LHN02, Lew94, Per93, Rab64, Sch93b, Sch93a, Wei74, Win98, Zuc88, BST03, BSST04, CST01, CST02, Wig92, Ano61c, Ano61f, Ano62e, Ano64a, Ano64b, Ano64c, Ano73, Ano79a, Ano79b, Ano79c, Ano79d, Ano84b, Ano87a, Ano93b, Ano94, Ano98c, Ano05, Ano08, Asi6x, Ber87, Bes85, Bes89, BS50, Bro73b, Bro78, Can07, CHW91b, Cle08, Cof64, Csi98, Cso98a, Cso98b, Cze98, Dan98, Danxx, Dan15, Dol61, E.01, Eck64, EE72, Far92, FS87, Fel84, Fen78, FR13a, Fra05, Ger98, Ger99, Gra94, Gra96, Gru83, Gwe63, HH74, HGS87, Hawxx, Her93, Hor97, Hug80, Jea09, Jor79, Kel02, Kle05b]. Leo [Lan86a, Lan88, Lan89d, Lan90, LS92, Lan92b, Lan93a, Lan93h, LS94, Lan94b, Lan94c, Lan95b, Lan97b, Lan97c, Lan97d, Lan98a, Lan98c, Lan98b, Lan98e, Lan98f, Lan98h, Lan98i, Lan98j, Lan99b, Lan99c, Lan04a, Lan04b, Lan06c, Lan08b, Lan10, Lan11e, Lan13, Lan14a, Lan14b, Liv60, Maa04, Man65, Mar98a, Mar98b, Mar98d, Nef97, O^{+}01, PL88, Pea89, Per62b, Phi80, Pla58, Pod08, Puc60, Rad98, Rai02, Rip98, Rob14a, Ros64, Rot73, Rot98, Sal64, Sar82, Sco71, See95, She94, Shi64, Sza87, Sza98, SE45, Szi76b, SWS80, Ter93, Tho04, Tib98, WS78, Wei98, Wig69, Wig92, Wig96a, Ano98a, Ano98b, Bak98, Csi98, Cso98a, Cso98b, Cze98, Lan98j, Mar98d, Rad98, Rot98, SW98, Sza98, Wei98]. Leónak [Kon98]. Leopold [Inf70]. Leorol [Lan89b]. Leóval [Kle98]. Lesart [Kra93]. Leslie [Ber03]. Letter [Arn55, DFF^{+}63, Ein39, Fow55, Har65, Lan02a, Lan06a, Lan06b, Mar48, SG55, Szi47d, Szi55c, Szi60g, Szi61f, Szi61c, Szi61d, Szi61e, Szi63e, Wei94, Wei93, Dem15, Szi98c, Wac06]. Letters [Kat75, Szi55d]. Levél [Szi98c]. Leverett [ACF^{+}42]. Lewis [Szi63e]. Liberated [SC34b]. Liberating [See03]. Liberation [BLW^{+}34b, Szi34c]. lie [Bro54]. Life [Fen78, Har65, Lan94b, Lan98i, Lan05b, Ler60, Swe93, ABES12, BO76, Csi98, Kle05a, Szi53a, Szi54a]. Light [Boy85, NS49]. Light-Reactivation [NS49]. Like [BC12]. Lilienthal [Szi50c]. Limitations [Bus93]. Linking [Bro85]. Links [Szi79]. Liquid [Szi45c, FS42a]. list [Szi64a]. Literature [Can00]. Livable [Ger98, Ger99, Pea89, Zuc88, HGS87]. Live [Gwe63, Rab60, Szi60f, Szi63d]. Liveable [Oli62]. lives [Szi53a]. livets [Kle05a]. Lobbies [Ano62a]. Lobby [Ano61d, Ano61g, Ano61h, Lew61, Per62b, Weh61]. Lobbying [Lan64]. Logic [KG12b]. London [Fri74]. Long [Szi60f, Szi64f]. Long-Range [Szi60f]. Long-Term [Szi64f]. look [Har10]. Looking [Ano61c, Ano61f]. Lost [Dan15]. Louise [Bes89, Bes93, Win98]. Love [Dan15]. lunaires [VB15]. Lunar [Szi71, VB15]. Lunch [Lan93c]. Lyman [Szi39b].

M.I.T. [CMS42]. Mabley [Mab62]. Machine [BC12, Lan83]. Machines [SE28]. macroscopic [Par01]. Made [Hew94a, MB00]. Maggio [Wac06]. Magnetfelder [RS31]. magnetic [FS42a, RS31]. Magnetohydrodynamics [PB79]. magyarországi [ÉK98]. make [Gwe63]. maker [Log93]. Making [Kel04]. Mammalian [Szi62h, Szi55e]. Mammals [Szi61b]. Man [Bad94, Bet93, Bro73a, Hawxx, Her93, Hew94a, Hew94b, Lan93h, Lan13, Lan14a, Ler60, Lew94, Per93, Szi62c, Szi62g, Szi63f, Ter93, Bro54, LS92, LS94, Lan97a, RTS^{+}65, RTS^{+}69, Szi55e]. Manhattan [FHN^{+}45, Gru83, HHW99, Kel04, Lan90, Lan93i, LLGS95, Lan99e, Lan02c, Lan08a, Lan11d, Ree15, Sch15]. Many [Lan14b, Lew94]. March [Ano17]. Margin [Lan98f, Lan99c, Lan08b, Lan11e]. Mark [Szi62f]. Market [Ano61c, Lan10, Szi46d, Ano61f]. Marks [BB12, Sch00b]. Martian [Lan04b]. Martians [Har06, Lan93i, Lan99e, Mar94b, Mar96]. Marvels [Lan02c]. mass [FS50]. Massive [WSCF61]. Masters [Duf46]. material [FGMvH14]. Mathematicians [Rid84]. Maverick [Gru83]. Max [Blo04]. Maxwell [CDS12, Dev02, Dev03, EN98, EN99, Ehr67, Gou96, KG12b, KG12a, Kra21, Lef02, Rex17, Scu07a, Scu07b, VJ11, Zur03]. May [Blo04, Szi63i]. Mean [Jen11]. Meaning [Duf46, Hor97, MW46a, MW07]. Means [BLW^{+}34b]. meant [She85]. Measurement [Cha11, Dev02, Dev03, KG12b, Lub87]. Measurements [Zur03]. mechanical [BS95]. Mechanism [Ank60]. Mechanisms [NS51a, vBDL96]. med [Tib98]. media [WWCS42]. medicine [All01]. Medium [Gro96]. Meet [Wie65]. Meeting [Szi53c, ZB69, LSW58]. meetings [SK63]. Meets [Rei84]. megatons [GC02]. Megawatts [GC02]. megemlékezés [Tel98b]. meinem [Szi53a, Szi54a]. meinen [Szi54a]. meiner [Szi53a]. melting [Szi41e]. Members [Szi45b]. Memo [Szi61g]. Memoir [Shi64, Wig69, Bus93]. Memoirs [Szi39d, SSSS95]. Memorandum [DS57, FS41, LSW58, Szi34b, Szi39b, Szi41c, Szi41d, Szi42b, Szi42c, Szi42a, Szi60h, Szi64c, Szi64d, Szi78b, Szi42f]. Memorial [Ano98a, Lan98e]. Memories [McM94, Szi53a, Szi54a]. Memory [Ank60, Szi64f, Szi64g]. Men [Mal08, Str63, Szi98a, Smi07, Szi98b, Van03]. Menace [Ano76, Yor75]. Merton [Tho04]. message [Sak83]. Metabolites [Szi51]. Metal [FS41, Szi45c]. Metallen [Szi26]. Metallurgical [FHN^{+}45]. Metals [SE28, Szi26]. Method [FS57a, Jor79, Szi52a, Szi57b, Szi57a, Szi26, Szi41d]. MGM [Rei84]. Michael [Gus12]. Microbial [Szi51]. microcanonical [MP10, VJ11]. microscope [Gab57, Gab74, Mar68, Mar94a]. Microscopy [Haw13]. Migration [Pal10, Fer68, Fer71, FB69]. migrations [Fra09]. Military [Gro46, Cha99]. million [Per62b]. Millions [MB00]. Minds [Szi53c]. Mined [Szi61h]. miner [ÉK98]. Minimal [Szi64e]. mining [ÉK98]. Minority [Sch05, Szi98a, Szi98b]. mins [Wac06]. Miracle [Rab47]. mistake [Ano60b]. mixture [Scu07b]. Mobile [HS11]. Mobilizer [Sch93b, Sch93a]. mode [Gra94, Gra96]. model [AACZ09, Pal05]. model-based [Pal05]. Modeling [VJ11]. Moderated [Szi58a, SWC59, Szi44c]. Modern [Anoxx, Bus93, Gib19, Jor79, Lan02c, PB79]. Modernität [Pri03]. modernity [Pri03]. Modified [Dev03]. molds [Szi26]. Molecular [Lan92b, Szi60b, Szi64f, CSW92]. monk [Tho04]. Month [Ano11, Cho10, CMS42, MS42, MLC^{+}42]. Moonshine [Jen11]. Moore [ACF^{+}42]. Morality [Lan11c]. mortal [Mar98c]. most [Har10]. möten [Kle87]. Motion [Lan04b, SMM^{+}60, SBSW60]. Motor [BC12]. Movement [SE28, Wau62, Rot98, Smi65b, Smi70]. Movie [Wac06]. mozgalom [Rot98]. Mr. [Mar48]. Muegyetemen [Sza98]. Multiparticle [CDS12]. Multiple [Fra05]. Muse [Can00]. mushroom [Bes93, Win98]. must [Pol19]. mutagens [NS52]. Mutations [NS50b, Szi61b, NS51a]. My [Szi49b, Szi48, Szi53a, Szi54a]. Mystery [FR13b, Scu07a]. mystics [Scu07a]. myth [Mar96].

Nachtrag [Szi54a]. Nádor [Sza98]. Nagasaki [Lan05a]. Narrative [Szi63c, Com56, Szi81]. Narrow [Lan98f, Lan99c, Lan08b, Lan11e]. National [GR63, Lan93i, Per62b, Lan99e]. Nature [Ano61e, ZBZ^{+}09, Szi41f, Szi59a]. Nazi [Lan98g, MP01, WHF^{+}96]. Nazism [Ols63]. neguentropy [Gou96]. Netz [Lan86b]. Neumann [Hor97, Lan93i, Lan99e]. Neutron [AFS39, FS41, SFA^{+}43, Szi45c, ABF^{+}43, ABF^{+}47, FSS46, FSS47]. Neutronen [BJ41, BLW^{+}34a]. Neutronic [FS55a, FS57a, FS57b, FS58a, FS61, Szi57b, Szi58a, Szi58b]. Neutrons [AFH39, BLW^{+}34b, CGS35, GHS39, GS37, SC34b, Szi35a, SC35, SZ39, Szi41a, Szi41c, SBFA48, Tur06, ZS39, BFS47, BJ41, BLW^{+}34a, MS41b, MS41a, SZ41]. Never [Ano95, Rei95]. News [Kra21]. Newspaper [Szi61f]. next [LM97]. Niels [Duf46, BS50]. nine [Mar06, Van03]. Nineteen [Ano61i]. No [Szi35b]. Nobel [Wol67, Zsa76]. Nobel-díjas [Zsa76]. Noise [KG12b]. None [Duf46, MW46a, MW46b, MW07]. nonisothermality [AACZ09]. nonproliferation [FGMvH14]. Note [Lan96b, Moo92]. Notes [Liv60]. November [Cho10]. Nuclear [Ano60a, Ano62d, Ano62c, Ano85, BHT86, Bad05, Can00, Far92, GHS39, Hig60, Jas93, Lan92b, Lan96c, Lan97b, Lan98i, Lan09a, Lan09b, LTD02, Pea89, PT02, Pod08, Seg85, Szi34c, Szi39a, Szi59b, Wat93, Wei79, Wig96b, Zuc88, All01, And73, Ano11, Ber87, CHW91a, Far13b, Far13c, Far13d, FGMvH14, Fer55, GC02, HGS87, Her76, Kle05b, Lan12a, Lif80, Mar94b, Pol19, Ree15, Rob14a, Rós98, SF44, SK63, Tho04, Her76]. nucleation [AACZ09]. Nucleus [FR13b]. nukleáris [BM98]. Numerical [HS98]. Nuremberg [Bus93, Bus93]. nyertesei [BST03, BSST04, CST01, CST02].

O [Danxx]. obituary [Ano64c]. Obligation [Kel02]. Observations [Szi46e]. Observer [Ano76]. Occasional [Szi62g, Szi63f]. October [CMS42]. Odd [Lan00a, Lan01e]. Offers [Rob6x]. Oh [Mil05]. OM [Blo04]. once [Ols63]. once-great [Ols63]. One [Gib19, Hew94a, MW46a, MW07, Scu07b, GBL05, Har10, LTD02, MW46b, Duf46]. one-atom [GBL05]. ont [VB15]. Opening [Sch00b]. Operating [Szi57b, DS57, FS57a, Szi59b]. Operation [Hig60]. Operon [ZB69]. Opinion [Gol92c, Woo62]. Oppenheimer [Gol92a, Ken54, O^{+}01, Rai02, Val06]. Opportunity [Rid84]. Oppose [Bet93, Szi76b]. optimal [CDS12]. organisms [Szi55e]. Origins [Pri95, CSW92]. öröksége [BM98]. Országos [Ano98a]. oscillation [Szi25]. Österreichischer [Ano02]. other [HA62, LTD02, SZ41, Szi61j, Szi62d, Szi92]. Our [Szi45b, And74b, And74a, Sch12b]. Outline [Jor79]. overtook [Far13b]. Ownership [Ano55a]. Oxford [Cha99].

P [Bes89, Bes93, WS92, WS03, Win98, Wol67, Szi44c]. P-9 [Szi44c]. P. [Mal08]. P.O [Wac06]. Pacific [Ken54]. Palatine [Sza87]. Pandora [Van03]. Papers [Ano98c, Ano05, KNSB54, Mur73, ZB69, FS87, Rob14d, Rob14b, Wig96b, Wei74, Fri74, Gab73, Gol74]. Parade [Sco71]. parallel [FS42b, Tho04]. parallel-flow [FS42b]. parancsnokságot [PM98]. Párbeszéd [Zsa76]. pareja [Lan01e]. Parity [LL12, Szi64e, Szi65]. Part [Ano79d, Ano79b, EN98, EN99, Wig96b]. Particles [Szi34a, MP10]. particularly [SE28]. past [Bro85]. Patent [Ano55a, Cho10, PB79, Tur06]. patently [Dan01]. Patents [Ano55b]. paths [Tho04]. Paul [BB12, Hor97, Jor79]. Pax [Szi60f]. Pay [Tur06]. Pb [BFAS47, SBFA47, SBFA48]. Peace [Ano61d, Ano61h, Ano62b, Fow55, Hig60, Lan02d, Lew61, LM97, Smi62, Szi55a, Szi55b, Vol63, Wau62, Wei61, Ber87, Bes93, Szi61a, Wei98, Win98]. Peacemaker [Lan96a, Lan97c, Lan14b]. Peaches [CL18, LC02, Lan04c, Lan12b]. People [Lew61, Per62b, Per62a]. Perceptible [Szi62g, Szi63f]. perfection [Kle05a]. peril [Smi65b, Smi70]. Periodical [Ano76]. Periodicals [Ano87b]. perpetual [Yor75]. Perpetuum [HS11]. Persecution [BB12, MWW11]. Personal [Bus93, Szi47e, Com56, Maa04, Sil98]. Personality [Sco71]. Perspective [Seg85]. Perspectives [Kel04, Zuc88]. Perutz [Blo04]. Petition [Dan95, LLGS95, Szi63g, Szi76b, Lan96b]. Phage [CSW92]. phänomenologischen [Szi25]. phase [LL12, Par01]. Phenomena [Nag81, Szi22, Szi25]. phenomenological [Szi25]. Phenotype [NS51b, Szi62g, Szi63f]. Philipsborn [Dan15]. Photo [Szi41a, MS41b]. Physical [Wal04]. Physicist [Ano64a, Ano64b, Dan15, Far92, Lan96a, Lan97c, Lan14b, Szi47f, ABES12, Her76, Rob14a, Rob14b, Seg70, Tho04]. Physicists [Bad05, Jor79, Rid84, Har06, Jon10, Kle05b]. Physics [Anoxx, Ano11, Can00, Cho10, Eck64, Fer55, Lan92b, Lan98f, Lan99c, Sal64, Szi34b, Csi98, Lan92a, Ree15, Szi78b, WP85, Wig96b]. pianeta [Szi03]. picture [Ano94, SMM^{+}60, SBSW60]. pile [Szi44b, Szi44c]. Pioneer [Ano55b, Ano61a, Ano61g, Ano62b, Rob14a, Rot98]. Placement [BB12, MWW11]. Plan [Ano62b, Mab62, Per62b, Per62a, Woo62]. planet [Szi03]. Plant [ACF^{+}42, Szi42d, LCM^{+}42, Szi42b, Szi42c]. plants [Rós98]. Platform [Szi45b]. plays [Kle98]. Plight [BB12, MWW11]. plutonium [MB00]. poisoning [Szi44b]. Polanyi [Gus12]. polarisierter [RS31]. Polarisierung [MS26]. polarization [MS26]. polarized [RS31]. policies [CHW91a]. Policing [Fin60]. Policy [AB96, Haf96, Lan15, Oli62]. Polish [Lan01c]. Political [FHN^{+}45, Szi61i, Woo62, O^{+}01, Rai02]. Politicians [Lan09c]. Politics [Aro96, Hab69, Lan64, Lan98f, Lan99c, Lan05d, Szi47f, Wie65, Bes89, Far13a, Lan06c]. Poll [CD48]. popular [Per62b]. populations [FS55b]. Portraits [SE97]. Portrévázlat [Lan89b]. Position [Szi47a, Szi78a, Szi45a]. Possibility [Szi39b, Szi60f, FS50]. Possible [Ank60, Ano61a, Szi34b, ST60, Szi78b]. Post [Lan92b]. Post-War [Lan92b]. Postscript [Hawxx]. Postwar [Szi47a, Szi60i, Szi78a, Szi45a]. potere [Val06]. powder [Szi41e]. Power [Ano85, Ano17, Lan89c, Lan99a, PT02, Szi42d, GC02, Lan12a, LTD02, LCM^{+}42, Rós98, Szi42b, Szi42c, Szi42f, Szi42g, Val06, Jor79]. Powers [Cas93, Swe93, Ber93, Kra93]. Pp [Bus93, Cas93, Duf46, Fri74]. Practice [MSHP06]. Preceded [Ken54]. predictions [Gwe63]. Preliminary [ABF^{+}43, ABF^{+}47, MS41b, MS41a, SZ41, Szi41e]. Preparation [Szi52a]. Prescience [Ano98b, Lan98h]. Present [Kat75, CHW91a]. Presented [ZB69, KNSB54]. President [Dan95, Ein39, LLGS95, Szi61f, Szi63g]. Press [Fri74, Vol63]. Pressure [Weh61, Cha11, FS42b, Sch00a]. Prevent [Szi55d, Pol19]. prevention [Lif80]. Princeton [Lan11a]. principle [Bro73b]. Prize [Smi66, Wol67]. Prize-winner [Wol67]. Problem [Bri56b, Bri62b, Bri04b, Bri13b, DS57, Fow55, Gro96, Szi55a]. Problems [FHN^{+}45, Szi59b]. Process [Haf96, Szi39c, Szi51, Szi55e, Szi44a, Szi59a, Szi63b]. processes [KK11, KNSB54]. Processive [BC12]. produced [Wac06]. Producing [Szi39c, Szi51]. Production [AFS39, AFH39, BLW^{+}34a]. Products [Szi52a]. professionals [Fra09, Pal10]. Professor [Szi42a, Zsa76]. Professzori [BST03, BSST04, CST01, CST02]. professzorral [Zsa76]. Program [Tur06]. programs [Gwe63]. Progress [Szi42e, Szi46e]. Project [FHN^{+}45, Fer55, Gru83, HHW99, Kel04, Lan90, Lan93i, LLGS95, Lan99e, Lan02c, Lan08a, Lan11d, Ree15, Sch15]. Prompt [Wal97]. Prophet [EE72]. Proposal [Chi60, DS57, Mar48, Szi44b]. Proposals [Szi60i]. Proposed [Szi55d, LSW58]. Proposes [Ano61d]. Protein [BC12]. Proteins [Szi60a]. provided [Lan01a]. Proving [Ken54]. Provision [Szi46e]. Provisional [Szi35b]. Public [Duf46, Haf96, Kel02, Lan15, DS57, MW46a, MW07]. Public-Policy [Haf96]. publications [Szi64a]. Pugwash [Kub98, Ano57, Ano87a, Kub98, Rot98]. Pugwash-Konferenzen [Kub98]. Pugwash-mozgalom [Rot98]. pump [Cha11, FS42a, Sch00a]. Pumpkin [Gus12]. pure [Mil05]. Pythagorean [Scu07a].

Qualified [Szi55d]. Quantum [GBL05, KSDU11, Zur03, BS95, CDS12, KK12, LZL^{+}12, LL12, Scu07a]. quantum-mechanical [BS95]. quatre [VB15]. quest [Com56, Mar94b]. question [Szi41d]. questions [And73]. Qui [Dem15, VB15].

R [Ber03, Jor79]. rêveurs [VB15]. Ra [ABF^{+}43]. Rabi [Zuc88]. Rabinowitch [Smi66]. racconti [Szi62d]. Race [Duf46, Lan96c, Lan97b, Lan99a, Per62b, Per62a, Far13b, Szi46a, Szi07]. racket [Szi48]. radiant [Mil05]. Radiation [Ano55c, BD02, Szi61b]. Radio [Szi39c]. Radio-Active [Szi39c]. Radioactive [SC34a, BLW^{+}34a]. Radioactivity [BLW^{+}34b, GHS39, SC34b, SC35]. radioaktiver [BLW^{+}34a]. Radium [Szi41a, ABF^{+}47, MS41b]. radon [ABF^{+}43, ABF^{+}47]. Raised [Szi59b]. raising [GBL05, Szi41d]. Rally [Ano61g]. randomness [Par01]. Range [Szi60f]. Rapporteur [Osb97]. rapporto [Szi03]. rates [NS54]. Ratio [Szi60d]. Rational [See95]. Raymond [Swi45]. Rays [BLW^{+}34b, GS37, SC34b, BLW^{+}34a, MS25, MS26]. reacting [Szi41d]. Reaction [BHT86, FS41, Szi41b, Szi41c, And73, Ano11, FS50, Her76, Mar94b, Szi46b]. Reactions [FS58b, Szi34c, SF44]. Reactivation [NS49]. Reactor [Ano55a, FS55a, FS57a, FS58a, FS61, Szi57b, Szi58a, Szi58b, FS57b, Lan01d]. Reactors [Szi57a, SWC59]. Ready [Fra05]. Real [Mal08, Smi07]. Realism [Bes93, Win98]. Realities [Har65]. reality [Hor97]. Really [Jen11]. Realm [Far92]. Reason [Lan99d, Lan03]. Rebels [Bes89]. Recall [Szi64g]. recast [Bes89]. Receive [Ano55b]. receives [Rob14b]. Recognition [Ano55a]. Recollections [Szi53b, Sil98, SWS80, WS78, WS92, WS03]. Recommendation [Gro46]. Reconsidering [Ber03]. reduction [Szi29a, Szi76a]. reflection [MS26]. Reflections [Lan98i, Kle87, Kle90, Tib98]. Reflexion [MS26]. reflexionen [Tib98]. Refrigerating [SE28]. Refrigeration [ES30, Cha11, Del98]. Refrigerator [Ano76, Cho10, Dan97b]. Refrigerators [Ano76, Dan97a]. Refugee [Cos84, Seg85, Jon10]. Refugees [BB12, MWW11]. Regime [MP01, KK12]. Region [Szi41a, MS41b]. Register [Ano98c]. Regulation [Jas93]. Reins [Gol92b]. relating [Szi35b, Szi36]. relations [Gou96]. Release [BLW^{+}34a]. released [Lau46]. Reluctant [HH74]. remembered [Cro04]. Remembering [Kel04]. remembers [Rob14c]. Remembrance [Ros64, Maa04]. Remind [Duf46]. Reminiscences [SW68, Szi69, Tel83]. reply [Szi50c]. Report [ACF^{+}42, Duf46, FHN^{+}45, FHN^{+}46, FHN^{+}63, Mab62, MLC^{+}42, Osb97, Sch05, Szi42e, Szi49c, Szi59b, Szi82, MS41b, MS41a, MW46a, MW07, Pei43, SZ41, Szi41e, Szi03, CMS42, MS42, Pri95]. Representatives [Szi45b]. Repression [Szi60h, Szi62h]. requirement [KG12a]. Rescue [PT02]. Research [Nag81, Szi72, DS57, Tib98, DS57]. Residual [Szi35a]. Resistive [Gro96]. Respondent [Lan00b]. Response [Lau94]. Responsibility [Bad05, Lan09a, Vol63, NS98]. Retrospectives [She85]. returned [Mar98d]. returns [Mag98]. Revelations [Ber93]. Review [Ano52, Bad94, BB12, Ber93, Bet93, Bus93, Cas93, Duf46, Fri74, Gol74, Han93, Hew94a, Hew94b, Jas93, Lau94, Lew94, Low93, Mal08, McC65, Pal10, Rex17, Sch93b, Vol63, Wac06, Wei74, Ols63, Wal04]. Reviews [Sch93a, Swe93]. Revising [Arn55]. Revisited [Lub87, ZBZ^{+}09, Ber90, Par01]. Revisiting [LZL^{+}12, Rob14a]. Revolution [Lan05d, Lan06c]. Rickover [Hew94a]. Rigden [Zuc88]. Right [Sch15]. Risk [Szi54b]. Road [Ano62d, Dan00, Szi62a]. Robert [O^{+}01, Rai02, Lau94]. Rocket [Szi60f]. Rockwell [Hew94a]. Role [Ken54, Kle05b]. Rollin [Ano62d]. Röntgenstrahlen [BLW^{+}34a, MS26, MS25]. Roosevelt [Dan96, Dem15, Ein39, Lan02a]. Roots [Mar98a, Pri95]. Rövid [Tel98b]. rpm [Wig92]. Rules [Ano60a]. Running [Lan90]. runs [Lau46]. Russia [Szi52b]. Russian [Lan05c, LSW58, Szi60i]. Russo [Szi60f]. Rüstungskontrolle [Kub98]. Rutherford [Jen11].

S [Zuc88]. Salk [Ano64d, Bou13, Per93, Rob14a]. Salvaging [Wei79]. Samuel [Jas93]. Satellite [Lau94]. Saturation [Szi64e, Szi65]. save [Gwe63]. Says [Ano62d, Ano62c]. Scale [Szi39b]. scattered [Ols63]. Scattering [SBFA48, BFS47, BFAS47, MS25, SBFA47]. Schawlow [Rip96]. Scholar [Ano61d]. Scholars [Cos84]. school [Rad98]. Schwankungserscheinungen [Szi22, Szi25]. Science [Ano46, Ano50b, Aro96, Bri56a, Bri62a, Bri04a, Bri13a, Cle08, Gib19, Gra94, Hab69, Haf96, Jea09, Lan92b, Lan00b, Lan05d, Swe93, Szi48, Szi72, Szi98a, Viz98, Wie65, AJ82, Far13a, Gra96, Lan06c, Mar96, O^{+}01, Puc60, Rai02, She94, Szi81, Szi98b, Val06, Har06, Jor79, Szi81]. Science-fiction-Erzählungen [Szi81]. Sciences [Szi46e]. Scientific [FS87, Fri74, Gab73, Gol74, Jor79, Wei74, Ols63, Wig96b, Ano76]. Scientist [Ano61c, Ano61d, Cof64, Fel84, Oli62, Ols63, Per62b, Rob6x, Vol63, Weh61, Zuc88, Bes85, NS98, Smi65b, Smi70, Ano61f]. Scientists [BFM^{+}49, CD48, Duf46, Fie07, Fin60, GR63, Jor79, Kel02, Lan64, LLGS95, Lan95d, Lan96c, Lan09a, Lan09c, Lan12a, MP01, Seg85, Ste75, STS^{+}49, Wea76, Yav78, Har10, LSW58, Sch12b, SK63, GR63, Jor79]. scienza [Val06]. Search [Haf96, Szi42e, Szi41f]. Seasons [Fel84]. Second [Ben87, Lef02]. secondary [Rad98]. Secrecy [Ano66]. Secret [Ber93, Swe93, Wal04, Amr59, Amr60, Cas93, Pow93, Wea76]. Secretary [FHN^{+}46, FHN^{+}63]. secure [Szi61a]. Security [Jea09, Ken54, Szi50d, Szi54b, Szi59b, LM97, Yor75]. Seeking [Lan10]. Seen [Sta99]. Sees [Ano61a, Hig60]. selected [SWS80, WS78]. selective [MS25]. selektiven [MS25]. self [Ano11]. self-sustained [Ano11]. Seminar [Lan93e, Lan94c]. Senate [Szi45b]. Sense [Fie07, Gwe63]. Sensitive [Szi98a, Szi98b]. Separation [SC34a]. September [ZB69, Far13c, Far13d, MS42]. Series [Lan93b]. seriously [Gwe63]. Servants [Jor79]. Service [Lan98e]. set [Rob14a]. settlement [Szi61i]. Seven [Tel83]. Sex [Szi60d, Szi62h]. Shadow [Lan98i, BL96]. Shadows [Bad94, Bet93, Han93, Her93, Hew94a, Hew94b, Lan93e, Lan93f, Lan95b, Lan13, Lew94, Per93, Ter93, LS92, LS94, Lan97d, Lan05c, Sch93a, Low93, Sch93b]. Shall [Szi49a, Szi49d, Szi63a]. Shaped [Gib19]. Shell [FS41]. Shield [Zuc88]. Short [Szi42f]. Shula [BB12]. si [Szi85]. Silard [Bad94, Per93, Sch93b, Low93, Han93]. similar [Szi55e]. simple [MS25]. simulations [HS98]. Single [Sch00a, Cha11]. Site [Szi42e]. Sites [Pri03]. Six [Gwe63]. Sixteen [Rot73]. Sixty [FR13b]. Skapelsens [Kle05a]. Slideshow [Dan98]. Slow [SZ39, Tur06]. slowing [Szi55e]. slugs [Szi44a, Szi44b]. Small [SMM^{+}60, SBSW60]. Smith [Mal08]. Smithsonian [BL96]. So-Called [Szi60f]. Social [Bad05, FHN^{+}45, Lan09a, NS98]. Society [Gib19, Szi61f, Szi72, KNSB54, Jor79]. Soddy [Gus12]. Somber [Chi60]. Some [Szi79, Ols63]. Sooner [STS^{+}49]. Sophistication [Lan64]. Sounders [Yor75]. Sources [Szi41a, ABF^{+}43, ABF^{+}47, MS41b]. South [Wac06]. Soviet [Lau94, Kub98, SSSS95, STS^{+}49, SK63]. Sowjetische [Kub98]. space [AJ82]. Spanish [Lan01e]. Speaks [Cle08]. Special [SSSS95]. Specific [Szi60a]. Specification [Szi35b]. Spencer [Jor79]. Spermatozoa [Szi63b]. Spherical [FS41]. Spies [Lan94a]. Spirit [Gib19]. Sponsorship [Szi55d]. Spontaneous [NS50b, NS51a]. spread [LTD02]. Spring [ZB69]. Sputnik [Lau94]. spymaster [SSSS95]. står [Tib98]. staden [Kle87]. Staff [Weh61]. Stage [Szi60f]. Stalemate [Szi60f]. Stalin [Mar48, Szi47d, Szi98c]. state [FS55b]. Statement [Ano57, Szi87]. States [Pal10, Dan95, Far13b, Fra09, Rid84, Szi45a, Szi47a, Szi63g, Szi78a]. Statesmen [Lan95d]. steady [FS55b]. stemme [Szi64b]. Step [Szi55b]. Stewardship [Lan97b]. Still [Cle08]. Stimme [Szi63c, Szi81]. Sting [Szi65]. Stockpile [Lan97b]. Stop [Lan96d, Szi60j]. storia [Val06]. stories [Szi61j, Szi62d, Szi92]. Story [Dan97b, Lan89c, Lau46, MP01, Ree15, RTS^{+}65, RTS^{+}69, Val06]. Strains [NS51b]. Strangelove [Mal08, Smi07]. strategies [Bes93, Win98]. Strauss [Szi63e]. Strength [Szi63b]. Striker [PM98, PM98]. Ströme [Szi26]. Strong [Wau62]. Studiengruppe [Kub98]. Studies [Ano64d, Sza87, FSS46, FSS47, Sza98, DS57]. Study [Kub98, Pal05]. Stumps [Wau62]. subjects [Szi76a]. Subversion [Lan05d, Lan06c]. Successful [Ken54]. Sudoplatov [Les94, McM94]. Suggested [Oli62, Szi34c]. Suggestions [Szi41f]. Suggests [Woo62]. Suicide [Arn50a]. sulla [Szi85]. sun [HHW99]. superweapon [Smi07, Mal08]. Suppression [Szi62c]. Surface [Nag81]. survey [Rós98]. Survive [Szi60f, Szi63d, GS50, Rab60]. sustained [And73, Ano11]. Swedish [Kle87, Kle05a, Tib98]. Swing [Swi45]. switch [Szi45d]. Sword [Zuc88]. symmetry [Par01]. symposium [KNSB54, Sch00b]. Synchronous [Szi34a]. synthesis [NS54]. System [Szi29a, Szi41c, Szi64h, Cha11, Szi46a, Szi76a, Szi07]. systems [Szi46b]. Szanton [WS92, WS03]. Személyes [Sil98]. Szilard [Rot48, Ano61f, Ano61g, Ano76, Ano98a, Bad94, Bak98, Bes89, Bes93, Bet93, Danxx, E.01, Fel81, Fri74, Gab73, Gol74, Hew94a, Hew94b, Jea09, Jor79, Kel02, Lan96a, Lan98d, Lan01d, LHN02, Lew94, Mar98b, Nef97, Per93, Pod08, Rab64, Weh61, Wei74, Wig92, Wig96a, Win98, Zuc88, Far13c, Far13d, Mar98b, All01, AACZ09, And74b, And74a, Ano50a, Ano55a, Ano61c, Ano62e, Ano64a, Ano64b, Ano64c, Ano64d, Ano73, Ano79a, Ano79b, Ano79c, Ano79d, Ano84a, Ano84b, Ano87a, Ano93b, Ano94, Ano98c, Ano05, Ano08, Asi6x, Bal05, Ber90, Ber87, Bes85, Bet93, BS95, BC12, BS50, Bri56b, Bri62b, Bri04b, Bri13b, Bro73b, Bro78, Bul79, Bye02, Bye04, CDS12, Can07, CHW91b, Cha11, Chi60, Cho10, Cle08, Cof64, Csi98, Cso98a, Cze98, Dan97a]. Szilard [Dan97b, Dan98, Dan01, Dan15, Dev02, Dev03, Dol61, EN98, EN99, Eck64, EE72, Far92, FS87, Fel84, Fen78, FR13a, Fra05, Gab73, GBL05, Ger98, Ger99, Tam98, Gol74, Gol92a, Gou96, Gra94, Gra96, Gro46, Gru83, Gwe63, Kle98, Haf96, HH74, Har65, HS98, HGS87, Hawxx, HS11, Her93, Her98, Hig60, Hor97, Hug80, Hut61, Irw61, KK11, KSDU11, KK12, KG12b, KG12a, Kle05b, Kon98, Lan64, Lan86a, Lan88, Lan89b, Lan89d, Lan90, LS92, Lan92a, Lan92b, Lan93a, Lan93h, Lan93i, LS94, Lan94b, Lan94c, Lan95b, Lan96f, Lan97b, Lan97c, Lan97d, Lan98a, Lan98c, Lan98b, Lan98e, Lan98f, Lan98h, Lan98i, Lan99b, Lan99c, Lan99e, Lan01d, Lan01a, Lan04a, Lan04b, Lan06c, Lan08b, Lan10, Lan11a, Lan11e, Lan13, Lan14a]. Szilard [Lan14b, Lew61, LZL^{+}12, Liv60, LL12, Lub87, Maa04, Mag96, Man65, MP10, Mar48, Mar98a, Mar98d, Mur73, Nag81, O^{+}01, Pal05, PL88, PM98, Par01, Pea89, Pei43, Per62b, Phi80, Pla58, PB79, Puc60, Rad98, Rai02, Rip96, Rip98, Rob14a, Rob14d, Ros64, Rot73, Rot98, Sal64, Sar82, Sch93b, Sco71, Scu07b, SW98, See95, She94, Shi64, Smi60, Sza87, Sza98, SE45, Szi47c, SMM^{+}60, SBSW60, Szi76b, SWS80, Tel98a, Tel00, Ter93, Tho04, Tib98, VJ11, Wau62, WS78, Wei98, Wig69, Woo62, Zei70, ZBZ^{+}09, Zur03, Ano98a, Ano98b, Bak98, BST03, BSST04, BM98, Csi98, Cso98a, Cso98b, Cze98, CST01, CST02, Tam98, Kle98, Her98, Ber98, Kon98, Lan98j, Mar98d, PM98, Rad98, Rot98, SW98, Sza98, Wei98, Wig92, Sch93a, Wei74]. Szilard-Machine-Like [BC12]. Szilardian [Lan05d]. Sztálinhoz [Szi98c].

T [Ano93a, Fri74, Gab73, Gol74, Wei74]. taken [Gwe63]. Takes [Gol92b]. Taking [Kra21]. Találkozások [Kle98]. Talk [Jea09, Pod08, Wei61, Ano12, Rob14a]. Talks [Lan93a]. tankar [Kle87]. tanulmányai [Sza98]. társadalmi [NS98]. Tartaglia [Gro96]. tasks [SSSS95]. Taylor [Bus93]. Technical [Sza87, Sza98]. Technique [SC34b]. Technological [CMS42, MS42, MLC^{+}42]. Technologies [Gus12]. Technology [Zuc88]. Telford [Bus93]. Tell [Wei94]. Teller [BO76, Har10, Hor97, Lan93i, Lan99e, Szi87, Zuc88]. Ten [Bro78, Wac06]. teni [Lan05c]. Term [Szi64f]. terminal [Szi03, Szi49c, Szi82, Szi03, Ano52]. Terrible [Duf46]. Test [Fin60, Ken54, Szi63h, MS25]. Their [Cos84, HHW99, Sta99, Bes93, Hor97, Lan93g, Lan12a, Szi52a, Win98]. them [Ols63]. Theodore [Hew94a, Hor97]. Theory [AB96, Bri56a, Bri62a, Bri04a, Bri13a, Jor79, Lea59, Szi59c, AACZ09, Gou96]. there [Ols63]. Therefor [Szi58b]. thermal [BJ41]. thermischer [BJ41]. Thermodyamics [Lef02]. Thermodynamic [Szi64h, Szi22, Szi29a, Szi76a]. thermodynamics [Szi25]. Thermodynamik [Szi25]. thermodynamischen [Szi22, Szi29a]. thermonuclear [FS50]. Thinking [Ano62e, Low93]. Thirty [Yor75]. Thomas [Ber93, Cas93, Swe93, Tho04]. Thompson [Bes89, Bes93, Win98]. thorium [Szi44b]. Thought [Boy85]. Threat [Duf46]. threats [Kle05b]. Three [And73, Lan96d]. threshold [FSS46, FSS47]. Tibor [Pal10]. Tiger [Gus12]. till [Tib98]. Time [And74b, And74a, Ano17, Weh61, KK11, Szi60e]. time-backward [KK11]. time-forward [KK11]. times [BO76, Arn55, SG55, Szi55d]. Tisztelet [Kon98]. Titanic [Jon10]. today [Tib98]. told [WS92, WS03]. tolerance [Bro73b]. took [PM98]. Topchiev [LSW58]. története [ÉK98]. Tóth [Wig92, Wig92]. Trackdown [Ols63]. Trade [Szi46d]. tragedy [Kle05a]. tragic [She94]. tragik [Kle05a]. Transformers [Szi34a]. transitions [LL12, Par01]. Transmutation [Szi35b, Szi36, Szi39a]. treatment [BS95]. Treaty [Kub98]. tree [Cze98]. tréfamester [Lan98j]. Trial [Szi49b]. Trials [Bus93]. Trinity [And74a]. truck [ÉK98]. True [MP01]. Truman [LLGS95, Wal97]. Trust [Mur73]. trying [Gwe63]. Tube [Szi29b, Szi29c]. Tubes [BLW^{+}34b, Szi29d]. tudós [NS98]. tudóstársadalomban [Szi98b]. turned [Szi45d]. twentieth [Har06, Puc60, Har10]. Twenty [Wol67].

Twenty-Five [Wol67]. Two [Fin60, Smi62, DS57, WWCS42]. types [Rós98].

U.N. [Arn55, DFF^{+}63]. U.S [Vol63]. U.S. [AB96, Ano61g, Ano62c, Szi76b, Tur06]. U.S.A [Bes89]. U238 [SFA^{+}43]. UCSD [Rob14a, Rob14d, Rob14c, Rob14b]. Ultimate [Lan99a, Hor97]. Ultra [NS49]. Ultra-Violet [NS49]. Uncertainties [Swe93]. Uncertainty [Swe93]. undgaa [Szi46c]. Unexpectedly [Wac06]. unique [Man65]. unit [Szi42f, Szi42g]. United [Dan95, Pal10, Szi63g, Far13b, Fra09, Rid84, Szi45a, Szi47a, Szi78a]. universe [Lau46, Sen21]. Universities [Ano66]. University [FHN^{+}45, Sza87, Fer55, Sza98, Szi41d]. unknown [Ss45]. Unloading [Szi57a]. Unmaking [FGMvH14]. Unravelling [FR13b]. Unreliable [Les94]. unwanted [SSSS95]. Unwinding [Nef97]. uránbányászat [ÉK98]. Uranium [AFS39, AFH39, CL18, FS41, LC02, Lan04c, Lan12b, SZ39, Szi41a, Szi41c, Szi42g, ZS39, ÉK98, MS41b, SZ41, Szi41e, Szi46b]. Urges [Ano61g, Lew61, Oli62]. USA [Bes93, Win98]. Use [Ano62c, FSS46, FSS47, Lan09a, Szi76b, Szi44b, Wal97]. Utilizing [KG12b]. utolsó [ÉK98]. Utopia [Win98, Bes93]. Utopian [Szi63c]. utopische [Szi63c]. utter [Wal97]. úttöroje [Rot98].

V [LSW58, Lan05c]. Vannevar [Gol92c]. Varied [Lan89d]. Various [Rós98]. ved [Szi46c]. Verden [MW46b]. Verfahren [Szi26]. Version [Ano79c, Ano79d, Hug80, Jor79, Phi80, Sar82, Ano79a, Ano79b, GBL05, Mar94b, SWS80, WS78]. Versuch [MS25]. Versus [Duf46]. Vertrages [Kub98]. vi [Szi46c]. Victory [Ano62c]. Video [Wac06]. View [Lan02b]. Vignette [Ano93b]. világbéke [Wei98]. Világhoz [SW98]. Violet [NS49]. Virus [NS51b]. Viruses [NS51a]. Visionary [Sch93b, Sch93a, Bes85, Rob14a]. Vitae [Szi64a]. voce [Szi62d, Szi04]. Voice [Dol61, Far92, Hut61, McC65, Szi60c, Szi62d, Szi62e, Szi63c, Szi81, Szi04, Mar94b, Szi61j, Szi92, Szi05, Szi64b]. voix [Szi62e]. Vol [Wig96b]. Volume [Gol74, Mar98b]. vom [Szi54a]. vs [Szi64e]. VT [Wac06].

W [Szi44b]. Walker [Jas93]. want [Szi52b]. Wants [Weh61]. War [Ano60a, Ano61a, Ano61e, Ano61g, Ano62a, Ano62d, Bad05, Ber93, Bus93, Dan00, Fei66, Hig60, Lan92b, Lan93b, Lan02d, Mab62, Nef97, Oli62, Rob6x, Swe93, Szi49b, Szi55d, Szi62a, Vol63, Weh61, Cas93, Far13a, Lif80, Pod08, Pol19, Pow93, Szi52b, Szi60e, AB96, Bes89, FHN^{+}46, FHN^{+}63, Kub98, Lan08a]. Warning [BFM^{+}49]. Wars [Aro96, SG55]. Wartime [STS^{+}49]. Wary [Fin60]. Was [Ano60b, Lan95c, Ros64]. Washington [Lan10]. Water [Szi44c, Szi58a]. Watt [Sch12b]. Way [Duf46]. Weapons [Ano62c, Bad05, Lan09b, Szi59b, GC02, Lan12a, Lap68, Lap69, LTD02]. Weart [Jor79]. Weindling [BB12]. Weiss [Bes89, Bes93, Fel81, Fri74, Gab73, Gol74, Hawxx, Jor79, Wei74, Win98, Zuc88]. weld [Szi44a]. Well [Bri56b, Bri62b, Bri04b, Bri13b, She85]. Well-Informed [Bri56b, Bri62b, Bri04b, Bri13b]. Wells [See03]. Weltkulturerbe [Pri03]. went [Ols63]. Werner [Swe93]. Wesen [Szi29a]. Where [Lan93i, Lan99e, Wie65, Ols63]. whether [Szi41d]. which [Lau46]. White [Las72]. Whittlesey [Duf46]. Who [Duf46, Sco71, Bes89, Dem15, Gwe63, Har06, Mar06, VB15, Dem15]. Whys [Ano61e]. Wiesner [SMM^{+}60, SBSW60]. Wife [Wei94]. Wigner [Hor97, Lan93i, Lan99e, WS92, Wig96b, WS03, Wol67, Zsa76]. William [Han93, Hew94a, Low93, Per93, Sch93b, Bad94, Bet93, Hew94b, Lan88, Lan93a, PL88]. Williams [KNSB54]. winner [Wol67]. Wintour [BB12]. Wisdom [Ano61c, Lan10, Ano61f]. Wit [Lan98c, Lan98d, Lan98b, Lan98j]. without [LTD02]. Witness [Les94, SSSS95]. Wizards [Zuc88]. women [HHW99]. Wonder [Fie07]. Work [Hig60, Lew61, Szi59b, Csi98]. Worked [Bet93]. Working [Rab47]. Works [Fri74, Gab73, Gol74, Wei74, FS87, Wig96b]. World [Ano62d, Duf46, Ger98, Ger99, GR63, Gwe63, Oli62, Pea89, Szi47a, Szi59b, Szi78a, Zuc88, Bes89, HGS87, Mar06, MW46a, MW46b, MW07, SMM^{+}60, SBSW60, Szi61a, Wei98, Yor75, Bad05, Fei66, Lan93b, Pri03, SW98]. Worlds [Lan14b]. Would [Weh61]. Wrath [EN98, EN99]. Writing [Lan96f]. writings [BL96]. written [Szi54a]. Wrong [Szi42h]. Wspolpraca [Lan01c]. www.store.aetv.com [Wac06].

x [Duf46, BLW^{+}34a, BLW^{+}34b, MS25, MS26]. X-Rays [BLW^{+}34b, BLW^{+}34a, MS25, MS26]. xii [Bus93]. XIV [EN98, EN99]. xxii [Fri74].

Year [FR13b, Szi59b]. Years [Ano61i, Cay62, Gwe63, Las72, Wol67, Ano90, Bro54, Mar94b, Rad98, Yor75]. Yield [Szi49a, Szi63a]. York [Bus93, Duf46, ZB69, Szi42e].

Zehn [Bro78]. Zeisel [PM98]. Zeisel-Striker [PM98]. Zerstreuung [MS25]. Zseni [Lan97d]. zum [Szi26]. zur [Kub98, MS25].

References

Altman:2009:THN

[AACZ09] I. S. Al’tman, I. E. Agranovskii, M. Choi, and V. A. Zagainov. To the theory of homogeneous nucleation: The nonisothermality of the Szilard model. Russian Journal of Physical Chemistry A, 83(1): 63–70, January 2009. CODEN RJPCBS. ISSN 0036-0244 (print), 1531-863X (electronic).

Alperovitz:1996:TCW

[AB96] Gar Alperovitz and Kai Bird. A theory of Cold War dynamics: U.S. policy, Germany, and the bomb. The History Teacher, 29(3):281– 300, May 1996. CODEN ???? ISSN 0018-2745 (print), 1945-2292

(electronic). URL http://www.jstor.org/stable/494546. Brief

mention of Szilard’s 1929 patent on the cyclotron, and the 1939 Szilard–Einstein letter to President F. D. Roosevelt, with reference to [LS92].

Amaldi:2012:ALF

[ABES12] Edoardo Amaldi, S. (Saverio) Braccini, Antonio Ereditato, and Paola Scampoli, editors. The adventurous life of Friedrich Georg Houtermans, physicist (1903–1966). Springerbriefs in Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-32854-7 (paperback), 3-642-32855-5 (e-book). ISSN 2191-5423. xiv + 152 pp. LCCN Q124.6-127.2; QC16.H688 A43 2012. URL http://www.sif.it/

attivita/saggiatore/recensioni/e_amaldi.

Ashkin:1943:PCR

[ABF^{+}43] J. Ashkin, S. Bernstein, B. Feld, H. Kubitschek, and Leo Szilard. Preliminary comparison of radon–boron and Ra + Be neutron sources. Report CP-412, US Atomic Energy Commission, Washington, DC, USA, January 19, 1943. Reprinted in [FS87, pages 292–294].

Ashkin:1947:PCR

[ABF^{+}47] J. Ashkin, S. Bernstein, B. Feld, H. Kubitschek, and Leo Szilard. Preliminary comparison of radon–boron and radium–beryllium neutron sources. Report MDDC-1436, US Atomic Energy Commission, Washington, DC, USA, October 27, 1947.

Allison:1942:RCE

[ACF^{+}42] Samuel K. Allison, C. M. Cooper, Enrico Fermi, Eugene P. Wigner, and Leo Szilard. Report of the committee for the examination of the Moore–Leverett design of a He-cooled plant. Report CE-324, US Atomic Energy Commission, Washington, DC, USA, 1942. Included are four pages that seem to have been written by Enrico Fermi, chairman of the committee. Undated: year chosen according to report number.

Ackland:1982:DAA

[Ack82] Len Ackland. Dawn of the atomic age. Chicago Tribune Magazine, ??(??):23, November 28, 1982.

Anderson:1939:PNU

[AFH39] H. L. Anderson, E. Fermi, and H. B. Hanstein. Production of neutrons in uranium bombarded by neutrons. Physical Review, 55(8):797–798, April 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v55/i8/p797_2.

Anderson:1939:NPA

[AFS39] H. L. Anderson, E. Fermi, and Leo Szilard. Neutron production and absorption in uranium. Physical Review, 56(3):284–286, August 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://adsabs.harvard.edu/abs/1939PhRv.
..56..284A;   http://prola.aps.org/abstract/PR/v56/i3/

p284_1. Reprinted in [FS87, pages 160–162].

Asimov:1982:LSF

[AJ82] Isaac Asimov and J. O. Jeppson. Laughing space: funny science fiction. Houghton Mifflin, Boston, MA, USA, 1982. ISBN 0-395-30519-5. xvi + 520 pp. LCCN PN6231.S42 L38.

Allibone:1980:DGJ

[All80] T. E. Allibone. Dennis Gabor. 5 June 1900–9 February 1979. Biographical Memoirs of Fellows of the Royal Society, 26(??):106–147, November 1980. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL http://www.jstor.org/stable/769781.

This memoir discusses Leo Szilard’s unsuccessful attempt to convince his close friend Dennis Gabor (Gábor Dénes in Hungarian) to build the electronic microscope that Szilard had invented.

Allen:2001:SEN

[All01] A. D. Allen. Szilard endorsed nuclear medicine by example. Physics Today, 54(2):82, February 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://adsabs.

harvard.edu/abs/2001PhT....54b..82A.

Amrine:1959:GDS

[Amr59] Michael Amrine. The great decision: the secret history of the atomic bomb. Putnam, New York, NY, USA, 1959. 251 pp. LCCN D767.2 .A5.

Amrine:1960:GDS

[Amr60] Michael Amrine. The great decision: the secret history of the atomic bomb. Heinemann, London, UK, 1960. 251 pp. LCCN D767.2 .A5 1960.

Anderson:1973:TQA

[And73] Herbert Anderson. Three questions about the sustained nuclear chain reaction. The University of Chicago Magazine, 65(5):3–7, March/April 1973. ISSN 0041-9508. URL http://library.ucsd.

edu/dc/object/bb3960604v.

Anderson:1974:AOTb

[And74a] Herbert L. Anderson. “All in our time”: Fermi, Szilard, and Trinity. Bulletin of the Atomic Scientists, 30(8):40–47, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also part 1 [And74b].

Anderson:1974:AOTa

[And74b] Herbert L. Anderson. “All in our time”: The legacy of Fermi and Szilard. Bulletin of the Atomic Scientists, 30(7):56–62, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Continued in [And74a].

Anker:1960:PBM

[Ank60] H. S. Anker. A possible biochemical mechanism for memory. Nature, 188(4754):938, ???? 1960. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Anonymous:2016:MTA

[Ano17] Anonymous. The march of time: Atomic power! Periscope Film LLC archive, April 17, 201617. URL https://www.youtube.com/

watch?v=iTJJ7tkD5L4.

Anonymous:1946:GSD

[Ano46] Anonymous. The great science debate. Fortune Magazine, 33(6): 117–??, June 1946.

Anonymous:1950:CSA

[Ano50a] Anonymous. Correction to Szilard article. Bulletin of the Atomic Scientists, 6(5):152, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:SHH

[Ano50b] Anonymous. Science: Hydrogen hysteria. Time Magazine, ??(??): ??, March 6, 1950. CODEN TYMEA9. ISSN 0040-781x (print),

2169-1665 (electronic). URL http://content.time.com/time/
magazine/article/0,9171,858694,00.html.

Anonymous:1952:RBC

[Ano52] Anonymous. [review: Grand Central Terminal]. The University of Chicago Magazine, ??(??):??, ???? 1952. ISSN

0041-9508. URL http://magazine.uchicago.edu/0206/class-
notes/ourpages.html.

Anonymous:1955:PIF

[Ano55a] Anonymous. Patent is issued on first reactor; Fermi–Szilard invention gets recognition — A. E. C. holds ownership. New York Times, page 19, May 19, 1955. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/113334584/. See entry [FS55a]

for the patent.

Anonymous:1955:PAI

[Ano55b] Anonymous. Pioneer atom inventions receive patents. The Science News-Letter, 68(9):134, August 27, 1955. CODEN SNLEAI. ISSN 0096-4018 (print), 2326-1285 (electronic). URL http://www. jstor.org/stable/3935856. Discusses US patents 2,714,577 and 2,714,668, applied for in 1945, but not awarded for a decade, and after Fermi’s death.

Anonymous:1955:RHH

[Ano55c] Anonymous. Radiation held harmful. New York Times, ??(??):6, May 14, 1955. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-

667X, 1553-8095. URL http://search.proquest.com/docview/
113179966.

Anonymous:1957:PCS

[Ano57] Anonymous. The Pugwash Conference: Statement. Bulletin of the Atomic Scientists, 13(7):249–250, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). “The release of the following statement with the three committee reports was approved by all the above-listed participants, with the exception of Dr. Foster, who objected to the Report of Committee II and Dr. Szilard, who objected to those of Committees I and II.”.

Anonymous:1960:HBR

[Ano60a] Anonymous. The H-bomb: Rules for nuclear war. Newsweek, ?? (??):88–??, March 7, 1960. ISSN 0028-9604.

Anonymous:1960:WBJ

[Ano60b] Anonymous. Was A-bomb on Japan a mistake? U.S. News and World Report, ??(??):68–71, August 15, 1960. CODEN XNWRAV. ISSN 0041-5537. Interview with Leo Szilard.

Anonymous:1961:APA

[Ano61a] Anonymous. Atomic pioneer asks campaign to end war: Asks 2% of income; sees disarmament possible. The Plain Dealer, ??(??):??,

November 30, 1961. URL http://library.ucsd.edu/dc/object/
bb6554482m.

Anonymous:1961:BUD

[Ano61b] Anonymous. The bomb under Denver. Newsweek, ??(??):70, December 25, 1961. ISSN 0028-9604.

Anonymous:1961:CIL

[Ano61c] Anonymous. Close-up: ‘I’m looking for a market for wisdom:’ Leo Szilard, scientist. Life, 51(9):75–79, September 1, 1961. URL http:

//library.ucsd.edu/dc/object/bb9762703s.

Anonymous:1961:DAB

[Ano61d] Anonymous. Developer of atom bomb proposes scientist–scholar lobby for peace. Washington Post, ??(??):??, December 2, 1961.

ISSN 0190-8286. URL http://library.ucsd.edu/dc/object/
bb6554482m.

Anonymous:1961:HNW

[Ano61e] Anonymous. Human nature: The whys of war. Newsweek, ??(??): ??, December 11, 1961. ISSN 0028-9604.

Anonymous:1961:ILM

[Ano61f] Anonymous. I’m looking for a market for wisdom: Leo Szilard, scientist. Life Magazine, ??(??):??, September 1961. URL http:/

/library.ucsd.edu/dc/object/bb9762703s. Box 3, Folder 13 in

Leo Szilard Papers archive.

Anonymous:1961:LAW

[Ano61g] Anonymous. Lobby against war, A-bomb pioneer urges: Szilard acts to rally U.S. intellectuals. Chicago Daily Tribune, ??(??):??,

December 2, 1961. URL http://library.ucsd.edu/dc/object/
bb6554482m.

Anonymous:1961:LP

[Ano61h] Anonymous. Lobby for peace. Commonweal, ??(??):305–

306, December 15, 1961. URL http://library.ucsd.edu/dc/
object/bb14008552;   http://library.ucsd.edu/dc/object/
bb6554482m.

Anonymous:1961:NYL

[Ano61i] Anonymous. Nineteen years later. Chicago Sun-Times, ??(??):??,

December 3, 1961. URL http://library.ucsd.edu/dc/object/
bb6554482m.

Anonymous:1962:AAF

[Ano62a] Anonymous. Atom age’s ‘father’ lobbies against war. San Francisco News-Call Bulletin, ??(??):??, January 10, 1962. URL http://

library.ucsd.edu/dc/object/bb6554482m.

Anonymous:1962:APB

[Ano62b] Anonymous. Atomic pioneer bares peace plan. The Oregon ????,

??(??):??, ???? 13, 1962. URL http://library.ucsd.edu/dc/
object/bb6554482m.

Anonymous:196x:ESU

[Ano62c] Anonymous. Expert says use of nuclear weapons ‘inevitable’ if victory U.S. goal. The Oregonian, ??(??):??, January ??, 1962.

URL http://library.ucsd.edu/dc/object/bb6554482m.

Anonymous:1962:NES

[Ano62d] Anonymous. Nuclear expert says world rollin along on road to war. The Oregonian, ??(??):??, January 13, 1962. URL http://

library.ucsd.edu/dc/object/bb6554482m.

Anonymous:1962:TAL

[Ano62e] Anonymous. Thinking ahead with … Leo Szilard. International Science and Technology, ??(??):33–38, May 1962. CODEN

ISCTAB. ISSN 0097-3777. URL http://library.ucsd.edu/dc/
object/bb58718856.

Anonymous:1964:LSDa

[Ano64a] Anonymous. Leo Szilard dies, A-bomb physicist. New York Times, ??(??):??, May 31, 1964. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1964:LSDb

[Ano64b] Anonymous. Leo Szilard dies, A-bomb physicist. London Times, ??(??):??, June 1, 1964.

Anonymous:1964:LSO

[Ano64c] Anonymous. Leo Szilard [obituary]. Salk Institute Files, June 13, 1964.

Anonymous:1964:SIB

[Ano64d] Anonymous. Salk Institute for Biological Studies: Dr. Leo Szilard. Nature, 202(4929):238–239, April 1964. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://adsabs.

harvard.edu/abs/1964Natur.202T.238..

Anonymous:1966:EUS

[Ano66] Anonymous. Editorial: Universities and secrecy. New York Times, ??(??):46, September 7, 1966. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://library.ucsd.edu/

dc/object/bb7851423b.

Anonymous:1973:LS

[Ano73] Anonymous. Leo Szilard. In Encyclopedia of World Biography, volume 10, pages 316–317. McGraw-Hill, New York, NY, USA,

1973. URL http://library.ucsd.edu/dc/object/bb6554482m.

Anonymous:1976:POD

[Ano76] Anonymous. The periodical observer: Dynamic duo confronts refrigerator menace: “The Einstein–Szilard Refrigerators” by gene dannen, in Scientific American (jan. 1997). The Wilson Quarterly (1976–2012), 21(2):117–142 (133–134), Spring 1976. ISSN 0363-

3276. URL http://www.jstor.org/stable/40259469.

Anonymous:1979:ELSa

[Ano79a] Anonymous. Excerpts: Leo Szilard: His version of the facts. Bulletin of the Atomic Scientists, 35(2):37–40, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:ELSb

[Ano79b] Anonymous. Excerpts: Leo Szilard: His version of the facts: Part II. Bulletin of the Atomic Scientists, 35(3):55–59, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:LSH

[Ano79c] Anonymous. Leo Szilard: His version of the facts. Bulletin of the Atomic Scientists, 35(4):28–32, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:PFL

[Ano79d] Anonymous. Part four: Leo Szilard: His version of the facts. Bulletin of the Atomic Scientists, 35(5):34–35, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:ADL

[Ano84a] Anonymous. Address by Dr. Leo Szilard. Press release from The Nation., December 3, 1984.

Anonymous:1984:LSI

[Ano84b] Anonymous. Leo Szilard: Inventor of the atomic bomb, 1933. LP Recording by George Garabedian Production, 1984. Two 12-inch 33 1/3 rpm records (about 100 minutes).

Anonymous:1985:BBN

[Ano85] Anonymous. Background books: Nuclear power. The Wilson Quarterly (1976–2012), 9(5):132–133, Winter 1985. ISSN 0363-3276.

URL http://www.jstor.org/stable/40256965.

Anonymous:1987:SLP

[Ano87a] Anonymous. Leo Szilard — the 1st Pugwash. The Scientist (Philadelphia, PA), 1(16):24, June 29, 1987. ISSN 0890-3670 (print), 1945-5127 (electronic). Reprinted from Toward a Livable World, Leo Szilard and the Crusade for Nuclear Arms Control, 1987.

Anonymous:1987:P

[Ano87b] Anonymous. Periodicals. The Wilson Quarterly (1976–2012), 11 (4):13–44, Autumn 1987. ISSN 0363-3276. URL http://www.

jstor.org/stable/40257800.

Anonymous:1990:YABe

[Ano90] Anonymous. 35 years ago in the Bulletin. Bulletin of the Atomic Scientists, 46(8):5, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBF

[Ano93a] Anonymous. Bulletin: Bernard T. Feld. Bulletin of the Atomic Scientists, 49(3):5, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:VLS

[Ano93b] Anonymous. Vignette: Leo Szilard. Science, 261(5127):1462, September 10, 1993. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1994:CPL

[Ano94] Anonymous. Correction: [picture of Leo Szilard]. Bulletin of the Atomic Scientists, 50(5):3, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Les94].

Anonymous:1995:BBF

[Ano95] Anonymous. Background books: The future that never came. The Wilson Quarterly (1976–2012), 19(2):67, Spring 1995. ISSN 0363-

3276. URL http://www.jstor.org/stable/40258978.

Anonymous:1998:OSL

[Ano98a] Anonymous. Országos Szilárd Leó fizikaverseny. (Hungarian) [The Leo Szilard Memorial Competition]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.hu/

fszemle/archivum/fsz9804/tart9804.html. Second special is-

sue on the centenary of Leo Szilard’s birth.

Anonymous:1998:PCL

[Ano98b] Anonymous. Prescience and conscience: Leó Szilárd (1898–1964). Europhysics News, 29(3):92–93, ???? 1998. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic). URL http://www.

europhysicsnews.org/.

Anonymous:1998:RLS

[Ano98c] Anonymous. The register of Leo Szilard papers 1898–1998. Mandeville Special Collections Library Geisel Library University of California, San Diego, 1998. URL http://libraries.ucsd.edu/

speccoll/testing/html/mss0032a.html.

Anonymous:19xx:BMP

[Anoxx] Anonymous. The birth of modern physics. Los Alamos National Laboratory Web document, 19xx. URL http://www.lanl.gov/

history/road/birthofmodernphysics.shtml. Section The Ein-

stein Letter discusses the Einstein–Roosevelt letter of 2 August 1939, and says “drafted mostly by Szilard.”.

Anonymous:2002:HOA

[Ano02] Anonymous. Handbuch Österreichischer Autorinnen und Autoren jüdischer Herkunft 18. bis 20. Jahrhundert. K. G. Saur, München, Germany; New Providence, NJ, USA, 2002. ISBN 3-598-11545-8. xxiii + 1818 (3 volumes) pp. LCCN ????

Anonymous:2005:LSP

[Ano05] Anonymous. Leo Szilard papers. Mandeville Special Collections Library, 9500 Gilman Drive, La Jolla, California, 92093-0175,

USA., 2005. URL http://oac.cdlib.org/findaid/ark:/13030/
tf0z09n7k3. MSS 32.

Anonymous:2008:SL

[Ano08] Anonymous. Szilard, Leo. In Complete Dictionary of Scientific Biography. Encyclopedia.com, ????, 2008. URL http://www.

encyclopedia.com/doc/1G2-2830906128.html.

Anonymous:2011:MPH

[Ano11] Anonymous. This month in physics history: December 2, 1942: First self-sustained nuclear chain reaction. APS Physics, 20(11):

??, December 2011. URL http://www.aps.org/publications/
apsnews/201112/physicshistory.cfm.

Anonymous:2012:HAT

[Ano12] Anonymous. Harold Agnew talk delights audience. Los Alamos National Laboratory News Center, May 11, 2012. URL http://www.

lanl.gov/news/stories/agnew-colloquium.html. From the ar-

ticle: “Agnew, 92, fondly recounted his long association with Enrico Fermi, starting with his work on the team that created the first controlled nuclear reaction in a graphite pile at the University of Chicago. He remembers Fermi as, “a wonderful person, but just a regular guy,” who, in a very low-key manner, advanced key research in the development of the first atomic weapons and many other breakthrough scientific innovations.

He also told a story of Leo Szilard’s quirky behavior.

“At Chicago, Fermi used to give weekly lectures on Thursday evenings. Outside the lecture hall there was a sign up sheet, on a yellow legal pad. Szilard, who I think wasn’t sure we were going to win the war, refused to put his name on the sign up sheet, so Fermi had him listen to the lectures from the hallway, through an open door.”.

Arnold:1950:CS

[Arn50a] James Arnold. The cost of suicide. Newsweek, ??(??):??, October 30, 1950. ISSN 0028-9604.

Arnold:1950:HCB

[Arn50b] James R. Arnold. The hydrogen–cobalt bomb. Bulletin of the Atomic Scientists, 6(10):290–292, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also [Arn50a].

Arnett:1955:LTR

[Arn55] John H. Arnett. Letter to the Times: Revising the U.N. charter. New York Times, ??(??):176, February 13, 1955. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/docview/113168988. Response to

[Szi55d].

Aron:1950:ABE

[Aro50] Raymond Aron. The atomic bomb and Europe. Bulletin of the Atomic Scientists, 6(4):110–114, 125, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aronowitz:1996:PSW

[Aro96] Stanley Aronowitz. The politics of the science wars. Social Text, 62 (46/47):177–197, Spring/Summer 1996. CODEN ???? ISSN 0164-2472 (print), 1527-1951 (electronic). URL http://www.jstor.

org/stable/466853.

Asimov:196x:SL

[Asi6x] Isaac Asimov. Szilard, Leo. In Encyclopedia of Men and Women of Science, pages 466–. ????, ????, 196x. URL http://library.

ucsd.edu/dc/object/bb6554482m.

Badash:1994:BRW

[Bad94] Lawrence Badash. Book review: William Lanouette and Bela Silard: Genius in the Shadows: A Biography of Leo Szilard, the Man behind the Bomb. Isis, 85(1):172–173, March 1994. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235948.

Badash:2005:APN

[Bad05] Lawrence Badash. American physicists, nuclear weapons in World War II, and social responsibility. Physics in Perspective (PIP), 7(2):138–149, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.springerlink.

com/content/u2638016081u076h/.

Baksay:1998:SLD

[Bak98] Baksay László. A Szilárd Leó díj (Hungarian) [The Leo Szilard Award]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9804/

tart9804.html. Second special issue on the centenary of Leo Szilard’s birth.

Balaban:2005:SD

[Bal05] N. Q. Balaban. Szilard’s dream. Nature Methods, 2(9):648–649, September 2005. CODEN NMAEA3. ISSN 1548-7091 (print), 1548-7105 (electronic).

Bailyn:2012:BRB

[BB12] Lotte Bailyn and Bernard Bailyn. Book review: In Defence of Learning: The Plight, Persecution, and Placement of Academic Refugees, 1933–1980s, edited by Shula Marks, Paul Weindling, and Laura Wintour. Journal of Interdisciplinary History, 43(2):301–302, Autumn 2012. CODEN ???? ISSN 0022-1953 (print), 1530-9169 (electronic). URL http://muse.jhu.

edu/journals/journal_of_interdisciplinary_history/v043/
43.2.bailyn.pdf. See [MWW11].

Bethe:1950:FAHa

[BBSS50] Hans A. Bethe, Harrison S. Brown, Frederick Seitz, and Leo Szilard. The facts about the hydrogen bomb. Bulletin of the Atomic Scientists, 6(4):106–109, 126–127, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bier:2012:SML

[BC12] Martin Bier and Francisco J. Cao. Szilard-machine-like features in a processive motor protein. Acta Physica Polonica B, 43(5):

889–908, ???? 2012. CODEN APOBBB. ISSN 0587-4254 (print), 1509-5770 (electronic).

Bedford:2002:HCH

[BD02] Joel S. Bedford and William C. Dewey. Historical and current highlights in radiation biology: Has anything important been learned by irradiating cells? Radiation Research, 158(3):251–291, September 2002. CODEN RAREAE. ISSN 0033-7587 (print), 1938-5404

(electronic). URL http://www.jstor.org/stable/3580880.

Bennett:1987:DES

[Ben87] Charles H. Bennett. Demons, engines and the Second Law. Scientific American, 257(5):108–116, November 1987. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v257/n5/pdf/scientificamerican1187-108.pdf.

Bernstein:1976:ABC

[Ber76] Barton J. Bernstein, editor. The Atomic Bomb: The Critical Issues. Critical issues in American history series. Little, Brown and Co., Boston, MA, USA, 1976. xix + 169 pp. LCCN E183 .A85.

Bernstein:1987:LSG

[Ber87] B. J. Bernstein. Leo Szilard: Giving peace a chance in the nuclear age. Physics Today, 40(9):40–47, September 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://adsabs.harvard.edu/abs/1987PhT....40i..40B.

Berger:1990:SDR

[Ber90] Jorge Berger. Szilard’s demon revisited. International Journal of Theoretical Physics, 29(9):985–995, September 1990. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic). URL

http://adsabs.harvard.edu/abs/1990IJTP...29..985B.

Bernstein:1993:RFH

[Ber93] Jeremy Bernstein. Revelations from Farm Hall: Book review: Heisenberg’s War: The Secret History of the German Bomb, by Thomas Powers. Science, 259(5103):1923–1926, March 26, 1993. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/259/5103/
1923.extract.

Klara:1998:SCJ

[Ber98] Berei Klára. A Szilárd–Chalmers-jelenség és a forróatom-kémia. (Hungarian) [Szilárd–Chalmers effect and hot atom chemistry]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://

www.kfki.hu/fszemle/archivum/fsz9804/tart9804.html. Sec-

ond special issue on the centenary of Leo Szilard’s birth.

Bernstein:2003:RAG

[Ber03] Barton J. Bernstein. Reconsidering the “Atomic General”: Leslie R. Groves. The Journal of Military History, 67(3):883–920, July 2003. ISSN 0899-3718 (print), 1543-7795 (electronic). URL http:

//www.jstor.org/stable/3397331.

Bess:1985:LSS

[Bes85] Michael Bess. Leo Szilard: scientist, activist, visionary. Bulletin of the Atomic Scientists, 41(11):11–18, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bess:1989:RAC

[Bes89] Michael Demaree Bess. Rebels against the Cold War: Four intellectuals who campaigned to recast world politics, 1945–1985: Leo Szilard (U.S.A.), E. P. Thompson (England), Louise Weiss (France), Danilo Dolci (Italy). Ph.D., University of California, Berkeley, Berkeley, CA, USA, 1989. 581 pp. URL http://search.

proquest.com/docview/303672720.

Bess:1993:RUM

[Bes93] Michael Bess. Realism, utopia, and the mushroom cloud: four activist intellectuals and their strategies for peace, 1945–1989: Louise Weiss (France), Leo Szilard (USA), E. P. Thompson (England), Danilo Dolci (Italy). University of Chicago Press, Chicago, IL, USA, 1993. ISBN 0-226-04420-3 (hardcover), 0-226-04421-1 (paperback). xxviii + 322 pp. LCCN JX1962.A2 B47 1993. URL http://

www.loc.gov/catdir/description/uchi052/93009707.html;
http://www.loc.gov/catdir/enhancements/fy0608/93009707-
t.html;   http://www.loc.gov/catdir/enhancements/fy0609/
93009707-b.html.

Bethe:1993:BRS

[Bet93] Hans Bethe. Book review: Szilard worked first to build the bomb and then to oppose it: William Lanouette and Bela Szilard, Genius in the Shadows: A Biography of Leo Szilard, the Man Behind the Bomb. Physics Today, 46(9):63–64, September 1993. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//www.physicstoday.org/resource/1/phtoad/v46/i9/p63_s1.

Bernstein:1947:ISFc

[BFAS47] S. Bernstein, B. Feld, J. Ashkin, and Leo Szilard. Inelastic scattering of Fe, Pb, and Bi. Report MonP-375, US Atomic Energy Commission, Washington, DC, USA, September 25, 1947.

Brown:1949:SGN

[BFM^{+}49] Harrison Brown, James Franck, Joseph E. Mayer, Leo Szilard, and Harold C. Urey. Scientists give new warning. Bulletin of the Atomic Scientists, 5(10):264, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:1947:ISFb

[BFS47] S. Bernstein, B. T. Feld, and Leo Szilard. Inelastic scattering of fast neutrons. Report MDDC-1292, US Atomic Energy Commission, Washington, DC, USA, September 10, 1947. Reprinted in [FS87, pages 321–322]. Original report undated.

Badash:1986:NFR

[BHT86] Lawrence Badash, Elizabeth Hodes, and Adolph Tiddens. Nuclear fission: Reaction to the discovery in 1939. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 130(2):196–231, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.

jstor.org/stable/987181.

Bothe:1941:ATN

[BJ41] W. Bothe and P. Jensen. Die Absorption thermischer Neutronen in Elektrographit. (German) [The absorption of thermal neutrons in electrographite]. Report G-71, ????, ????, January 20, 1941. Captured German report cited in [Wea76, page 29] that contained incorrect results on the neutron capture cross-section of carbon. Those results caused the German Uranium Project scientists to switch from graphite to the hard-to-get deuterium (in heavy water obtained from the Rjukan Plant in Norway) as a neutron moderator, and likely, substantially delayed their progress in achieving a working nuclear reactor as a precursor to an atomic bomb.

Broda:1981:CBJ

[BL81] Englebert Broda and Robert Jay Lifton. Commentary: On behalf of Joliot. Bulletin of the Atomic Scientists, 37(3):61, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lif80].

Bird:1996:HSW

[BL96] Kai Bird and Lawrence Lifschultz, editors. Hiroshima’s shadow: writings on the denial of history and the Smithsonian controversy. Pamphleteer’s Press, Stony Creek, CT, USA, 1996. ISBN 0-9630587-3-8 (hardcover), 0-9630587-4-6 (paperback). lxxvii + 584 pp. LCCN 9609 BOOK NOT YET IN LC.

Blakeslee:1946:AF

[Bla46] Howard W. Blakeslee. The atomic future. Chicago Sun, ??(??):??, May 3, 1946. 28-page supplement.

Blow:2004:MFP

[Blo04] D. M. Blow. Max Ferdinand Perutz, OM CH CBE, 19 May 1914–6 February 2002. Biographical Memoirs of Fellows of the Royal Society, 50(3):227–256, ???? 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL http://www.jstor. org/stable/4140521. Brief mention of Szilard’s campaign to create an international laboratory of molecular biology on the model of CERN, with reference to [LS92].

Brasch:1934:ANB

[BLW^{+}34a] A. Brasch, F. Lange, A. Waly, T. E. Banks, T. A. Chalmers, L. Szilard, and F. L. Hopwood. Auslösung von Neutronen aus Beryllium durch harte Röntgenstrahlen. Erzeugung radioaktiver Elemente. (German) [Release of neutrons from beryllium by hard X-rays. production of radioactive elements]. Naturwissenschaften, 22(50): 839, December 1934. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://adsabs.harvard.edu/abs/
1934NW.....22..839B.

Brasch:1934:LNB

[BLW^{+}34b] A. Brasch, F. Lange, A. Waly, T. E. Banks, T. A. Chalmers, Leo Szilard, and F. L. Hopwood. Liberation of neutrons from beryllium by X-rays: Radioactivity induced by means of electron tubes. Nature, 134(3397):880, December 8, 1934. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http:

//adsabs.harvard.edu/abs/1934Natur.134..880B;   http:/
/www.nature.com/nature/journal/v134/n3397/pdf/134880a0.

pdf. Reprinted in [FS87, pages 147–148].

Browne:1998:SNO

[BM98] John C. Browne and Mezei Ferenc. Szilárd nukleáris öröksége és a jövő. (Hungarian) [The future of Szilárd’s legacy]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.hu/

fszemle/archivum/fsz9804/tart9804.html. Second special is-

sue on the centenary of Leo Szilard’s birth.

Blumberg:1976:ECL

[BO76] Stanley A. Blumberg and Gwinn Owens. Energy and conflict: the life and times of Edward Teller. Putnam, New York, NY, USA, 1976. ISBN 0-399-11551-X. xvii + 492 + 4 pp. LCCN QC16.T37

B58 1976. US$12.95. URL http://alsos.wlu.edu/information.
aspx?id=367.

Bourgeois:2013:GSI

[Bou13] Suzanne Bourgeois. Genesis of the Salk Institute: the epic of its founders. University of California Press, Berkeley, CA, USA, 2013. ISBN 0-520-27607-8 (hardcover). xxxiv + 230 pp. LCCN QH322.S25 B68 2013.

Boyer:1985:BEL

[Boy85] Paul S. Boyer. By the Bomb’s Early Light: American Thought and Culture at the Dawn of the Atomic Age. Pantheon Books, New York, NY, USA, 1985. ISBN 0-394-52878-6, 0-394-74767-4. xx + 440 pp. LCCN E169.12 .B684 1985. US$22.50.

Brillouin:1956:SIT

[Bri56a] Léon Brillouin. Science and Information Theory. Academic Press, New York, USA, 1956. 320 pp. LCCN Q175 .B786.

Brillouin:1956:SPW

[Bri56b] Léon Brillouin. Szilard’s problem: The well-informed heat engine. In Science and Information Theory [Bri56a], page ?? LCCN Q175 .B786.

Brillouin:1962:SIT

[Bri62a] Léon Brillouin. Science and Information Theory. Academic Press, New York, USA, second edition, 1962. 351 pp. LCCN Q175 .B786

Brillouin:1962:SPW

[Bri62b] Léon Brillouin. Szilard’s problem: The well-informed heat engine. In Science and Information Theory [Bri62a], page ?? LCCN Q175 .B786 1962.

Brillouin:2004:SIT

[Bri04a] Léon Brillouin. Science and Information Theory. Dover phoenix editions. Dover Publications, Inc., New York, NY, USA, second edition, 2004. ISBN 0-486-43918-6. xv + 351 pp. LCCN Q175

.B786 2004. URL http://www.loc.gov/catdir/enhancements/
fy0618/2004051975-d.html.

Brillouin:2004:SPW

[Bri04b] Léon Brillouin. Szilard’s problem: The well-informed heat engine. In Science and Information Theory [Bri04a], pages 176–178. ISBN 0-486-43918-6. LCCN Q175 .B786 2004. URL http://www.loc.

gov/catdir/enhancements/fy0618/2004051975-d.html.

Brillouin:2013:SIT

[Bri13a] Léon Brillouin. Science and Information Theory. Dover Publications, Inc., New York, NY, USA, second edition, 2013. ISBN 0-486-49755-0 (paperback). xv + 351 pp. LCCN Q360

.B75 2013. URL http://www.loc.gov/catdir/enhancements/
fy1307/2012051388-d.html;   http://www.loc.gov/catdir/
enhancements/fy1318/2012051388-t.html.

Brillouin:2013:SPW

[Bri13b] Léon Brillouin. Szilard’s problem: The well-informed heat engine. In Science and Information Theory [Bri13a], page ???? ISBN 0-486-49755-0 (paperback). LCCN Q360 .B75 2013. URL http:

//www.loc.gov/catdir/enhancements/fy1307/2012051388-d.
html;   http://www.loc.gov/catdir/enhancements/fy1318/
2012051388-t.html.

Brown:1954:CMF

[Bro54] Harrison Brown. The challenge of man’s future; an inquiry concerning the condition of man during the years that lie ahead. Viking Press, New York, NY, USA, 1954. 290 pp. LCCN GF31 .B68.

Bronowski:1973:AM

[Bro73a] Jacob Bronowski. The Ascent of Man. British Broadcasting Corporation, London, UK, 1973. ISBN 0-563-10498-8. 448 pp. LCCN Q175 .B7918 1973; CB151.

Bronowski:1973:PTL

[Bro73b] Jacob Bronowski. The principle of tolerance: Leo Szilard. In The Ascent of Man [Bro73a], pages 368–374. ISBN 0-563-10498-8. LCCN Q175 .B7918 1973; CB151.

Bronowski:1978:ZGL

[Bro78] Jacob Bronowski. Zehn Gebote von Leo Szilard. (German) [Ten commandments of Leo Szilard]. Available in the Leo Szilard Papers archive and in [WS78, pages vi–vii]., 1978.

Brown:1985:ELP

[Bro85] Harrison Brown. From the Editors; linking past and future. Bulletin of the Atomic Scientists, 41(7):4–7, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bohr:1936:CNB

[BS50] Niels Bohr and Leo Szilard. [correspondence: Niels Bohr and Leo Szilard]. Five letters between the authors held in the Leo Szilard Papers archive., 1936–1950. URL http://library.ucsd.edu/dc/

object/bb25612766.

Biedenharn:1995:QMT

[BS95] L. C. Biedenharn and J. C. Solem. A quantum-mechanical treatment of Szilard’s engine: Implications for the entropy of information. Foundations of Physics, 25(8):1221–1229, August 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02055259.

Batyi:2004:SLP

[BSST04] Emese Bátyi, Leo Szilard, Orsolya Szîucs, and Jánosné Takács. A Szilárd Leó Professzori èOsztöndíj nyertesei: 2004. Oktatási Minisztérium, Budapest, Hungarian, 2004. 60 pp. LCCN MLCS 2008/45435 (Q).

Batyi:2003:SLP

[BST03] Emese Bátyi, Leo Szilard, and Jánosné Takács. A Szilárd Leó Professzori èOsztöndíj nyertesei: 2003. MF Plusz, Budapest, Hungary, 2003. 64 pp. LCCN MLCS 2008/45436 (Q).

Bullard:1979:SHB

[Bul79] Edward Bullard. Szilard and his bomb. Nature, 278(5701):285– 286, March 1979. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://adsabs.harvard.edu/abs/

1979Natur.278..285B. Book review of [SWS80].

Bush:1993:BRB

[Bus93] Jonathan A. Bush. Book review essay: Nuremberg: The modern law of war and its limitations: The Anatomy of the Nuremberg Trials: A Personal Memoir, by Telford Taylor. Alfred A. Knopf, New York, 1992. Pp. xii, 703. $35.00. Columbia Law Review, 93 (8):2022–2086, December 1993. ISSN 0010-1958 (print), 1945-2268

(electronic). URL http://www.jstor.org/stable/1123010.

Byers:2002:FS

[Bye02] N. Byers. Fermi and Szilard. ArXiv Physics e-prints, July 2002.

URL http://adsabs.harvard.edu/abs/2002physics...7094B.

Byers:2004:FS

[Bye04] Nina Byers. Fermi and Szilard. In Cronin [Cro04], pages 215–228. ISBN 0-226-12111-9 (hardcover). LCCN QC16.F46

F49 2004. URL http://www.loc.gov/catdir/description/
uchi052/2003020524.html;   http://www.loc.gov/catdir/enhancements/
fy0614/2003020524-b.html;   http://www.loc.gov/catdir/
toc/ecip049/2003020524.html.

Canaday:2000:NML

[Can00] John Canaday. The Nuclear Muse: Literature, Physics, and the First Atomic Bombs. University of Wisconsin Press, Madison, WI, USA, 2000. ISBN 0-299-16850-6 (hardcover), 0-299-16854-9 (paperback). xviii + 310 pp. LCCN QC791.96 .C36 2000. URL http:

//www.e-streams.com/es0409/es0409_1493.htm.

Canaday:2007:LS

[Can07] John Canaday. ‘Leo Szilard’. Raritan — a Quarterly Review, 27 (1):61–62, SUmmer 2007. ISSN 0275-1607.

Cassidy:1993:ACB

[Cas93] David C. Cassidy. Atomic conspiracies: Book review: Heisenberg’s war: the secret history of the German bomb, by Thomas Powers, Knopf/Cape: 1993. Pp. 610. $27.50, £20. Nature, 363(6427):311– 312, May 27, 1993. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v363/n6427/pdf/363311a0.pdf.

Caylor:1962:JYL

[Cay62] Arthur Caylor. Just 10 years left. San Francisco News-Call Bulletin, ??(??):??, January 11, 1962. URL http://library.ucsd.

edu/dc/object/bb6554482m.

Compton:1948:PSC

[CD48] Arthur H. Compton and Farrington Daniels. A poll of scientists at Chicago, July 1945. Bulletin of the Atomic Scientists, 4(2):44, 63, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cai:2012:MQSa

[CDS12] C. Y. Cai, H. Dong, and C. P. Sun. Multiparticle quantum Szilard engine with optimal cycles assisted by a Maxwell’s demon. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 85(3 (part 1)):031114, March 14, 2012. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic).

Collie:1935:CBN

[CGS35] C. H. Collie, J. H. E. Griffiths, and L. Szilard. Collisions between neutrons and diplons. Nature, 135(3422):903–904, June 1, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v135/
n3422/pdf/135903b0.pdf.

Chambers:1999:OCA

[Cha99] John Whiteclay Chambers, editor. The Oxford companion to American military history. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-507198-0. 916 pp. LCCN E181 .O94 1999.

Chan:2011:MAB

[Cha11] Keng Wai Chan. Measurement and analysis of bubble pump and Einstein–Szilard single pressure absorption refrigeration system. Thesis (D.Phil.), Mathematical, Physical & Life Sciences Division, Oxford University, Oxford, UK, 2011. ix + 237 pp. URL http:

//ora.ox.ac.uk/objects/ora:5962;   http://ora.ox.ac.uk/
objects/uuid:f10706c5-6952-4c50-9aef-065f1627a19b;   http:
//search.proquest.com/docview/1414910237.

Childs:1960:SPD

[Chi60] Marquis Childs. Somber proposal from Dr. Szilard. Washington Post, ??(??):??, February 17, 1960. ISSN 0190-8286.

Chodos:2010:MPH

[Cho10] Alan Chodos. This month in physics history: November 11, 1930: Patent granted for Einstein–Szilard refrigerator. APS News, ??(??):??, November 2010. URL http://www.aps.org/

publications/apsnews/201011/physicshistory.cfm;   http:
//www.aps.org/publications/capitolhillquarterly/201012/
physicshistory.cfm.

Cantelon:1991:AAD

[CHW91a] Philip L. (Philip Louis) Cantelon, Richard G. Hewlett, and Robert Chadwell Williams, editors. The American atom: a documentary history of nuclear policies from the discovery of fission to the present. University of Pennsylvania Press, Philadelphia, second edition, 1991. ISBN 0-8122-3096-5 (hardcover), 0-8122-1354-8 (paper). xviii + 369 pp. LCCN UA23 .A597 1991.

Cantelon:1991:LSD

[CHW91b] Philip L. (Philip Louis) Cantelon, Richard G. Hewlett, and Robert Chadwell Williams. Leo Szilard and the discovery of fission. In The American atom: a documentary history of nuclear policies from the discovery of fission to the present [CHW91a], pages 7– 8. ISBN 0-8122-3096-5 (hardcover), 0-8122-1354-8 (paper). LCCN UA23 .A597 1991.

Cook:2018:UP

[CL18] Peter Cook and William Lanouette. Uranium + peaches. Stage play. World Premiere Performance, The Salk Institute for Biological Studies, Conrad T. Prebys Auditorium, La Jolla, California., June 21, 2018.

Clemente:2008:DSS

[Cle08] Bill Clemente. The dolphin still speaks: Leo Szilard and science fiction. Hungarian Journal of English and American Studies (HJEAS), 14(2):373–387, ???? 2008. CODEN ???? ISSN 1218-

7364. URL http://www.jstor.org/stable/41274436.

Creutz:1942:TDM

[CMS42] E. Creutz, John Marshall, Jr., and Leo Szilard. Technological division and M.I.T. group. Report for month ending October 15, 1942. Report CE-301, US Atomic Energy Commission, Washington, DC, USA, 1942.

Coffin:1964:LSC

[Cof64] Tristram Coffin. Leo Szilard: The conscience of a scientist. Holiday, ??(??):64–67, 92–99, February 1964. URL http://library.ucsd.

edu/dc/object/bb26977929.

Compton:1956:AQP

[Com56] Arthur Holly Compton. Atomic quest, a personal narrative. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1956. xix + 370 pp. LCCN QC773.A1 C65.

Coser:1984:RSA

[Cos84] Lewis A. Coser. Refugee Scholars in America: Their Impact and Their Experiences. Yale University Press, New Haven, CT, USA, 1984. ISBN 0-300-03193-9. xviii + 351 pp. LCCN ????

Cronin:2004:FR

[Cro04] James W. Cronin, editor. Fermi remembered. University of Chicago Press, Chicago, IL, USA, 2004. ISBN 0-226-12111-9 (hardcover). xi + 287 pp. LCCN QC16.F46 F49 2004. URL http://www.loc.

gov/catdir/description/uchi052/2003020524.html;   http:
//www.loc.gov/catdir/enhancements/fy0614/2003020524-b.
html;   http://www.loc.gov/catdir/toc/ecip049/2003020524.
html.

Csikai:1998:SLE

[Csi98] Csikai Gyula. Szilárd Leó életműve és hatása. (Hungarian) [Leo Szilard’s life work and its impact on physics]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.hu/

fszemle/archivum/fsz9802/tart9802.html. Special issue on the

centenary of Leo Szilard’s birth.

Csoori:1998:SLE

[Cso98a] Csoóri Sándor. Szilárd Leó emléktáblájának felavatása. (Hungarian) [Leo Szilard at home again]. Fizikai Szemle (Budapest), 48(??): ??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-

0540 (electronic). URL http://www.kfki.hu/fszemle/archivum/

fsz9804/csoori.html. Second special issue on the centenary of Leo Szilard’s birth.

Csoori::1998:SLE

[Cso98b] Csoóri Sándor. Szilárd Leó emléktáblájának felavatása. (Hungarian) [Leo Szilárd at home again]. Fizikai Szemle (Budapest), 48(??):

??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-

0540 (electronic). URL http://www.kfki.hu/fszemle/archivum/

fsz9804/tart9804.html. Second special issue on the centenary of Leo Szilard’s birth.

Czinkos:2001:SLP

[CST01] Tímea Czinkos, Leo Szilard, and Jánosne Takács. A Szilárd Leó Professzori èOsztöndíj nyertesei: 2001. Oktatási Minisztérium, Budapest, Hungary, 2001. 44 pp. LCCN MLCS 2008/45433 (Q).

Czinkos:2002:SLP

[CST02] Tímea Czinkos, Leo Szilard, and Jánosné Takács. A Szilárd Leó Professzori èOsztöndíj nyertesei: 2002. MF Plusz, Budapest, Hungary, 2002. 44 pp. LCCN MLCS 200845422 (Q).

Cairns:1992:POM

[CSW92] John Cairns, Gunther S. (Gunther Siegmund) Stent, and James D. Watson, editors. Phage and the origins of molecular biology. Cold Spring Harbor Laboratory Press, Plainview, NY, USA, 1992. ISBN 0-87969-407-6. x + 366 pp. LCCN QR342 .P45 1992.

Czeizel:1998:SLC

[Cze98] Czeizel Endre. Szilárd Leó családfája. (Hungarian) [Leo Szilard’s family tree]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9802/

tart9802.html. Special issue on the centenary of Leo Szilard’s birth.

Dannen:1995:PPU

[Dan95] Gene Dannen. A petition to the President of the United States: July 17, 1945. Web site., 1995. URL http://www.dannen.com/ decision/45-07-17.html. Contains text of the petition signed by Leo Szilard, Eugene Wigner, and 68 others to avoid use of the atomic bomb on Japan.

Dannen:1996:ERA

[Dan96] Gene Dannen. Einstein to Roosevelt, August 2, 1939. Web site.,

1996. URL http://www.dannen.com/ae-fdr.html.

Dannen:1997:ESR

[Dan97a] Gene Dannen. The Einstein–Szilard refrigerators. Scientific American, 276(1):90–95, January 1997. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v276/n1/pdf/
scientificamerican0197-90.pdf;   http://www.sciam.com/
0197issue/0197currentissue.html.

Dannen:1997:SRD

[Dan97b] Gene Dannen. Story of refrigerator designed by Einstein and Szilard gets iced. Physics Today, 50(12):102, December 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://adsabs.harvard.edu/abs/1997PhT....50l.102D.

Dannen:1998:LSI

[Dan98] Gene Dannen. Leo Szilard the inventor: a slideshow. A talk delivered at the Leo Szilard Centenary Eötvös University, Budapest, Hungary., February 9, 1998. URL http://www.dannen. com/budatalk.html. Published [with many typographical errors] in [Mar98b].

Dannen:19xx:LSE

[Danxx] Gene Dannen. Leo Szilard and Ernest O. Lawrence. Web site, 19xx.

URL http://members.peak.org/~danneng/lawrence.html.

Dannen:2000:WRW

[Dan00] Gene Dannen. Are we on the road to war? Web site., 2000. URL

http://www.dannen.com/roadtowar.html.

Dannen:2001:SIP

[Dan01] Gene Dannen. Szilard’s inventions patently halted. Physics Today, 54(3):102, March 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://adsabs.harvard.

edu/abs/2001PhT....54c.102D.

Dannen:2015:PLL

[Dan15] Gene Dannen. A physicist’s lost love: Leo Szilard and Gerda Philipsborn. Web site., 2015. URL http://www.dannen.com/

lostlove/.

Delano:1998:DAE

[Del98] Andrew Douglas Delano. Design analysis of the Einstein refrigeration cycle. Ph.D. thesis, Georgia Institute of Technology, Atlanta,

GA, USA, 1998. 171 pp. URL http://search.proquest.com/
docview/304432717.

Demortier:2015:QEQ

[Dem15] Guy Demortier. Qui est qui dans ^{≪}the letter to Roosevelt ^{≫}?. (French) [Who is who in “the letter to Roosevelt”?]. Revue des Questions Scientifiques, 186(1):125–152, ???? 2015. CODEN RQS-

CAN. ISSN 0035-2160. URL https://www.unamur.be/sciences/
philosoc/revueqs/textes-en-ligne/rqs_186_1_demortier.

Devereux:2002:SEM

[Dev02] Michael Devereux. Szilard’s Engine: Measurement, information, and Maxwell’s Demon. In Daniel P. Sheehan, editor, Quantum Limits to the Second Law: First International Conference on Quantum Limits to the Second Law. San Diego, California (USA), 29–31 July 2002. AIP Conference Proceedings, volume 643 of American Institute of Physics Conference Series, pages 279–284. American Institute of Physics, Woodbury, NY, USA, November 2002. ISBN 0-

7354-0098-9. URL http://adsabs.harvard.edu/abs/2002AIPC.
.643..279D.

Devereux:2003:MSE

[Dev03] Michael Devereux. A modified Szilard engine: Measurement, information, and Maxwell’s Demon. Foundations of Physics Letters, 16(1):41–55, February 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Doering:1963:LEU

[DFF^{+}63] William Doering, Bernard T. Feld, Jr., Allan Forbes, James G. Patton, and Leo Szilard. Letter to the Editor: U.N. in Cuba. Bulletin of the Atomic Scientists, 19(4):33, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dolbier:1961:VLS

[Dol61] Maurice Dolbier. The voice of Leo Szilard. New York Herald Tribune, ??(??):??, March 14, 1961.

Doering:1957:MCC

[DS57] William Doering and Leo Szilard. Memorandum to Cass Canfield: A proposal to create two interdependent research institutes operating in the general area of public health, designated as: ‘Research Institute for Fundamental Biology and Public Health’ and ‘Institute for Problem Studies’. Technical report, ????, ????, January 11, 1957. Reprinted in [FS87, pages 505–524].

Duffus:1946:ABV

[Duf46] R. L. Duffus. Atom bomb versus ‘human race’: The scientists who fashioned it remind us of its terrible threat: Review of One World or None. A Report to the Public on the Full Meaning of the Atomic Bomb. Edited by Dexer Masters and Kathlerine Way. Foreword by Niels Bohr. Introduction by Arthur H. Compton. x + 79 pp. New York: Whittlesey House. $1. New York Times, ??(??):BR1–BR2, March 17, 1946. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-

667X, 1553-8095. URL http://search.proquest.com/docview/
107455627.

E:2001:LS

[E.01] S. E. Leo Szilard 1898–1964. Europhysics News, 32(1):31, January/February 2001. CODEN EUPNAS. ISSN 0531-7479 (print),

1432-1092 (electronic). URL https://www.europhysicsnews.
org/articles/epn/pdf/2001/01/epn2001-32-1.pdf.

Eckart:1964:LSI

[Eck64] Carl Eckart. Leo Szilard’s influence on physics. Available in the Jonas Salk Papers archive at the University of California, San Diego., June 13, 1964.

Esterer:1972:PAA

[EE72] Arnulf K. Esterer and Louise A. Esterer. Prophet of the atomic age: Leo Szilard. Julian Messner, New York, NY, USA, 1972. ISBN 0-671-32522-1 (hardcover), 0-671-32523-X (MCE). 189 pp. LCCN QC16.S95 E83.

Ehrenberg:1967:MD

[Ehr67] W. Ehrenberg. Maxwell’s demon. Scientific American, 217(5): 103–110, November 1967. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v217/n5/pdf/scientificamerican1167-
103.pdf.

Einstein:1939:AEL

[Ein39] Albert Einstein. Albert Einstein’s letter to President Franklin Delano Roosevelt. World-Wide Web document, August 2, 1939. URL

http://hypertextbook.com/eworld/einstein.shtml;   http:
//www.anl.gov/Science_and_Technology/History/Anniversary_
Frontiers/aetofdr.html;   http://www.lanl.gov/history/
road/pdf/Einstein.pdf.

Erdi-Krausz:1998:UCM

[ÉK98] Érdi-Krausz Gábor. Az utolsó csille… — a magyarországi uránbányászat története. (Hungarian) [The last miner’s truck — the history of uranium mining in Hungary]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.hu/

fszemle/archivum/fsz9802/tart9802.html. Special issue on the

centenary of Leo Szilard’s birth.

Eliot:1950:HB

[Eli50] George Fielding Eliot, editor. The H bomb. Didier, New York, NY, USA, 1950. 175 pp. LCCN UF767 .H2. Introduction by Albert Einstein and commentary by George Fielding Eliot.

Earman:1998:EXW

[EN98] John Earman and John D. Norton. Exorcist XIV: The wrath of Maxwell’s Demon. Part I. from Maxwell to Szilard. Studies in History and Philosophy of Modern Physics, 29(4):435– 471, December 1998. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219898000239.

Earman:1999:EXW

[EN99] John Earman and John D. Norton. Exorcist XIV: The wrath of Maxwell’s Demon. Part II from Szilard to Landauer and beyond. Studies in History and Philosophy of Modern Physics, 30 (1):1–40, March 1999. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219898000264.

Einstein:1930:R

[ES30] Albert Einstein and Leo Szilard. Refrigeration. US Patent 1,781,541., November 11, 1930. URL http://www.google.com/ patents?id=t0BRAAAAEBAJ. Application filed December 16, 1927 (serial number 240,566) and in Germany, December 16, 1926. See [Dan97a, Dan97b] for accounts of this invention.

Farkas:1991:NPF

[Far92] Ákos Farkas. A nuclear physicist’s foray into the realm of fiction: Leo Szilard’s The Voice of the Dolphins. Hungarian Studies, 7(1–2):197–203, ???? 1991–1992. URL http://www.academia.

edu/5522155/A_NUCLEAR_PHYSICISTS_FORAY_INTO_THE_REALM_
OF_FICTION_LEO_SZILARDS_THE_VOICE_OF_THE_DOLPHINS.

Farmelo:2013:CBHa

[Far13a] Graham Farmelo. Churchill’s bomb: a hidden history of science, war and politics. Faber and Faber, London, UK, 2013. ISBN 0-571-24978-7 (hardcover). x + 554 + 12 pp. LCCN ????

Farmelo:2013:CBHb

[Far13b] Graham Farmelo. Churchill’s bomb: how the United States overtook Britain in the first nuclear arms race. Basic Books, New York, NY, USA, 2013. ISBN 0-465-02195-6 (hardcover), 0-465-06989-4 (e-book). vi + 554 pp. LCCN UA647 .F28 2013.

Farmelo:2013:SFSa

[Far13c] Graham Farmelo. September 1933 to February 1935 Szilárd’s nuclear epiphany. In Churchill’s bomb: a hidden history of science, war and politics [Far13a], page ?? ISBN 0-571-24978-7 (hardcover). LCCN ????

Farmelo:2013:SFSb

[Far13d] Graham Farmelo. September 1933 to February 1935 Szilárd’s nuclear epiphany. In Churchill’s bomb: how the United States overtook Britain in the first nuclear arms race [Far13b], page ?? ISBN 0-465-02195-6 (hardcover), 0-465-06989-4 (e-book). LCCN UA647 .F28 2013.

Fleming:1969:IME

[FB69] Donald Fleming and Bernard Bailyn, editors. The intellectual migration: Europe and America, 1930–1960. Belknap Press of Harvard University Press, Cambridge, MA, USA; London, UK, 1969. ISBN 0-674-33411-6, 0-674-33412-4. 748 pp. LCCN E169.1 .F6.

Feis:1966:ABE

[Fei66] Herbert Feis. The atomic bomb and the end of World War II. Princeton University Press, Princeton, NJ, USA, revised edition, 1966. vi + 213 pp. LCCN D767.2 .F4 1966. Leo Szilard is mentioned on pages 50–51.

Feld:1978:DD

[Fel78] Bernard T. Feld. The dilemma of deterrence. Bulletin of the Atomic Scientists, 34(3):12–13, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:GWS

[Fel81] Bernard T. Feld. Gertrud Weiss Szilard, 1909–1981. Bulletin of the Atomic Scientists, 37(10):56, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1984:LSS

[Fel84] Bernard Feld. Leo Szilard, scientist for all seasons. Social Research, 51(3):675–690, ???? 1984. ISSN 0037-783X.

Fenly:1978:ABE

[Fen78] Leigh Fenly. The agony of the bomb, and ecstasy of life with Leo Szilard. San Diego Union, ??(??):D–1, D–8, November 19, 1978.

Fermi:1955:PCU

[Fer55] Enrico Fermi. Physics at Columbia University: The genesis of the nuclear energy project. Physics Today, 8(11):12–16, November 1955. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/?PTO/8/12/1;
http://link.aip.org/link/phtoad/v8/i11/p12/s1.

Fermi:1968:III

[Fer68] Laura Fermi. Illustrious immigrants; the intellectual migration from Europe, 1930/41. University of Chicago Press, Chicago, IL, USA, 1968. xi + 440 pp. LCCN E184.A1 F47.

Fermi:1971:III

[Fer71] Laura Fermi. Illustrious immigrants; the intellectual migration from Europe, 1930/41. University of Chicago Press, Chicago, IL, USA, second edition, 1971. ISBN 0-226-24376-1, 0-226-24378-8 (paperback). xi + 431 pp. LCCN E184.A1 F47 1971.

Feiveson:2014:UBF

[FGMvH14] Harold A. Feiveson, Alexander Glaser, Zia Mian, and Frank N. von Hippel. Unmaking the bomb: a fissile material approach to nuclear disarmament and nonproliferation. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02774-7 (hardcover), 0-262-31918-7 (e-book). x + 4 + 277 pp. LCCN JZ5675 .F45 2014.

Franck:1945:RCP

[FHN^{+}45] James Franck, Donald J. Hughes, J. J. Nickson, Eugene Rabinowitch, Glenn T. Seaborg, J. C. Stearns, and Leo Szilard. Report of the Committee on Political and Social Problems, Manhattan Project ‘Metallurgical Laboratory’, University of Chicago (The Franck Report). U.S. National Archives, Washington, DC: Record Group 77, Manhattan Engineer District Records, Harrison–Bundy File, folder #76., June 11, 1945. URL http:

//datatorch.com/Science/Scientists_Stories.aspx?id=52;
 http://en.wikipedia.org/wiki/Franck_Report;   http://
nuclearsecrecy.com/blog/2012/01/11/weekly-document-9-the-
uncensored-franck-report-1945-1946/;   http://www.atomicarchive.
com/Docs/ManhattanProject/FranckReport.shtml;   http://
www.dannen.com/decision/franck.html;   http://www.fas.
org/sgp/eprint/franck.html;   http://www.nuclearfiles.
org/menu/key-issues/ethics/issues/scientific/franck-report.
htm;   http://www.wissenschaft-und-frieden.de/seite.php?

artikelID=1092. This is the famous minority report, written one month before the first successful test of the atomic bomb at the Trinity site in New Mexico, in which seven scientists, most of who worked on the Manhattan Project, recommended against the use of the atomic bomb on Japan, and predicted a dangerous post-war nuclear arms race. According to the Web document “The Uncensored Franck Report (1945–1946)”, by Alex Wellerstein, dated 11 January 2012, the original report was censored in at least eight places; of those, seven have since been recovered, and at least one sentence exists, but has been blacked out by government censors. According to [Har10, page 153], the final version was written by Leo Szilard and signed by 68 scientists.

Franck:1946:RSW

[FHN^{+}46] James Franck, Donald J. Hughes, J. J. Nickson, Eugene Rabinowitch, Glenn T. Seaborg, J. C. Stearns, and Leo Szilard. A report to the Secretary of War. Bulletin of the Atomic Scientists, 1(10):2–4, 16, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//datatorch.com/Science/Scientists_Stories.aspx?id=52;
 http://en.wikipedia.org/wiki/Franck_Report;   http://
nuclearsecrecy.com/blog/2012/01/11/weekly-document-9-the-
uncensored-franck-report-1945-1946/;   http://www.atomicarchive.
com/Docs/ManhattanProject/FranckReport.shtml;   http://
www.dannen.com/decision/franck.html;   http://www.fas.
org/sgp/eprint/franck.html;   http://www.nuclearfiles.
org/menu/key-issues/ethics/issues/scientific/franck-report.
htm;   http://www.wissenschaft-und-frieden.de/seite.php?

artikelID=1092. The authors are members of the Committee on Political and Social Problems, Manhattan Project ‘Metallurgical Laboratory’, University of Chicago, and this report subsequently became known as ‘The Franck Report’. The report was written on 11 June 1945, one month before the first successful test of the atomic bomb at the Trinity site in New Mexico. In it, seven scientists, most of who worked on the Manhattan Project, recommend against the use of the atomic bomb on Japan, and predict a dangerous post-war nuclear arms race. This version was partially declassified in February 1946; eight fragments of text have been removed at the order of government censors. Some of that censored text was later recovered, but at least one sentence may still be missing — see the Web document “The Uncensored Franck Report (1945– 1946)”, by Alex Wellerstein, dated 11 January 2012. Reprinted in [FHN^{+}63], and in a longer form, in [Smi65b, Appendix] and [Smi70, pages 371–383].

Franck:1963:RSW

[FHN^{+}63] James Franck, Donald J. Hughes, I. I. Nickson, Eugene Rabinowitch, Glenn T. Seaborg, Joyce C. Stearns, and Leo Szilard. A report to the Secretary of War. In Grodzins and Rabinowitch [GR63], pages 19–27. LCCN D842 .B78. Reprint of [FHN^{+}46].

Fiege:2007:ASS

[Fie07] Mark Fiege. The atomic scientists, the sense of wonder, and the bomb. Environmental History, 12(3):578–613, July 2007. ISSN 1084-5453 (print), 1930-8892 (electronic). URL http://www.

jstor.org/stable/25473133.

Finney:1960:TSW

[Fin60] John W. Finney. Two scientists wary on policing of atomic-test ban: Disclose concern at capital ceremony honoring four. New York Times, ??(??):3–??, May 19, 1960. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/114981445. Report on the Atoms for

Peace Awards with shared $150,000 prize presented in Washington, DC, on 18 May 1960 to Leo Szilard, Alving M. Weinberg, Eugene Wigner, and Walter H. Zinn.

Fowler:1955:LED

[Fow55] Harold S. Fowler. Letter to the editor: [Disarmament and the problem of peace]. Bulletin of the Atomic Scientists, 11(10):380– 381, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Szi55a].

Fernandez:2013:LS

[FR13a] Bernard Fernandez and Georges Ripka. Leo Szilard. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896– 1956 [FR13b], page 371. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:UMA

[FR13b] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (ebook). xviii + 522 pp. LCCN QC773 .F47 2013.

Frank:2005:ERG

[Fra05] Tibor Frank. Ever ready to go: The multiple exiles of Leo Szilard. Physics in Perspective (PIP), 7(2):204–252, June 2005. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://adsabs.harvard.edu/abs/2005PhP.....7..204F;
 http://www.springerlink.com/content/1422-6960/.

Frank:2009:DEM

[Fra09] Tibor Frank. Double exile: migrations of Jewish–Hungarian professionals through Germany to the United States, 1919–1945, volume 7 of Exil-Studien: eine interdisziplinäre Buchreihe. Peter Lang, Oxford, UK, 2009. ISBN 3-03911-331-3. 501 pp. LCCN ????

Frisch:1974:BRC

[Fri74] O. R. Frisch. Book review: The Collected Works of Leo Szilard: Scientific Papers, by Bernard T. Feld and Gertrude Weiss Szilard (eds.), London: MIT Press 1973, pp. xxii + 737. Physics Bulletin, 25(1):27, January 1974. CODEN PHSBB4. ISSN 0031-9112.

Feld:1941:MCC

[FS41] Bernard T. Feld and Leo Szilard. Memorandum on the critical condition for a fast neutron chain reaction inside a spherical shell of uranium metal. Report CF-338, Columbia University, New York, NY, USA, December 26, 1941. Reprinted in [FS87, pages 280–287].

Feld:1942:MPL

[FS42a] B. Feld and Leo Szilard. A magnetic pump for liquid bismuth. Report CE-279, US Atomic Energy Commission, Washington, DC, USA, July 14, 1942. Reprinted in [FS87, pages 351–358].

Feld:1942:EPD

[FS42b] B. T. Feld and Leo Szilard. Examples for pressure drop calculations in parallel-flow helium cooling. Report CP-308, US Atomic Energy Commission, Washington, DC, USA, June 18, 1942. Reprinted in [FS87, pages 346–350].

Fermi:1950:PIT

[FS50] Enrico Fermi and Leo Szilard. On the possibility of initiating a thermonuclear reaction in a mass of deuterium. Report, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, Summer 1950.

Fermi:1955:NRa

[FS55a] Enrico Fermi and Leo Szilárd. Neutronic reactor. US Patent 2,708,656., May 17, 1955. URL http://www.google.com/ patents?vid=2708656. US Patent Application 568,904 filed December 19, 1944. Reprinted in [FS87, pages 691–696].

Fox:1955:DGB

[FS55b] Maurice S. Fox and Leo Szilard. A device for growing bacterial populations under steady state conditions. Journal of General Physiology, 39(2):261–266, ???? 1955. CODEN JGPLAD. ISSN 0022-1295 (print), 1540-7748 (electronic). Reprinted in [FS87, pages 441–446].

Fermi:1957:MON

[FS57a] Enrico Fermi and Leo Szilárd. Method of operating a neutronic reactor. US Patent 2,798,847., July 9, 1957. Filed December 1, 1952, serial number 323,452.

Fermi:1957:NR

[FS57b] Enrico Fermi and Leo Szilárd. Neutronic reactor. US Patent 2,807,581., September 24, 1957. URL http://www.google.com/ patents/US2807581.. US Patent Application 621,838, filed October 11, 1945.

Fermi:1958:ACN

[FS58a] Enrico Fermi and Leo Szilárd. Air cooled neutronic reactor. US Patent 2,836,554, May 27, 1958. Filed May 29, 1945, serial number 596,465.

Fermi:1958:CR

[FS58b] Enrico Fermi and Leo Szilard. Chain reactions. Canadian Patent

552312., January 28, 1958. URL http://brevets-patents.ic.
gc.ca/opic-cipo/cpd/eng/patent/552312/summary.html.

Fermi:1961:ACN

[FS61] Enrico Fermi and Leo Szilard. Air-cooled neutronic reactor. Canadian Patent 620923., May 30, 1961. URL http://brevets-

patents.ic.gc.ca/opic-cipo/cpd/eng/patent/620923/summary.
html.

Feld:1972:CWL

[FS87] Bernard T. Feld and Gertrud Weiss Szilard, editors. The collected works of Leo Szilard: Scientific papers, volume 1. MIT Press, Cambridge, MA, USA, 1972–1987. ISBN 0-262-06039-6 (vol. 1: 1972), 0-262-19168-7 (vol. 2: 1978), 0-262-19260-8 (vol. 3: 1987). xxii + 737 pp. LCCN QC3 .S97. With Kathleen R. Winsor. Foreword by Jacques Monod. Introductory essays by Carl Eckart, Bernard T. Feld, Maurice Goldhaber, Aaron Novick, and Julius Tabin.

Feld:1946:UTD

[FSS46] Bernard T. Feld, R. Scalettar, and Leo Szilard. Use of threshold detectors for fast neutron studies. Report MDDC-897, US Atomic Energy Commission, Washington, DC, USA, December 26, 1946. Published in [FSS47]. Reprinted in [FS87, pages 328–328].

Feld:1947:UTD

[FSS47] Bernard T. Feld, R. Scalettar, and Leo Szilard. Use of threshold detectors for fast neutron studies. Physical Review, 71(??):464, ???? 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Gabor:1957:EEG

[Gab57] Dennis Gabor. Die Entwicklungsgeschichte des Elektronenmikroskops. (German) [The development history of the electron microscope]. Elektrotechnische Zeitschrift, Ausgabe A, 78(15):522– 530, August 1, 1957. CODEN EKZAA3. ISSN 0302-265X.

Gabor:1973:BBC

[Gab73] Dennis Gabor. Books: The Collected Works of Leo Szilard, Scientific Papers, edited by Bernard T. Feld and Gertrud Weiss Szilard. Bulletin of the Atomic Scientists, 29(7):51–52, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gabor:1974:HEM

[Gab74] Dennis Gabor. History of the electron microscope, from ideas to achievements. In J. V. Sanders, editor, Electron Microscopy 1974: abstracts of papers presented to the Eight International Congress on Electron Microscopy, held in Canberra, Australia August 25–31, 1974, volume 1, pages 6–12. The Australian Academy of Science, Canberra, ACT, Australia, 1974. ISBN 0-85847-027-6. LCCN ????

Gea-Banacloche:2005:QVS

[GBL05] J. Gea-Banacloche and H. S. Leff. Quantum version of the Szilard one-atom engine and the cost of raising energy barriers. Fluctuation and Noise Letters, 5(4):C39–C47, December 2005. CODEN FNLLAW. ISSN 0219-4775 (print), 1793-6780 (electronic).

Garwin:2002:MMF

[GC02] Richard L. Garwin and Georges Charpak. Megawatts and megatons: the future of nuclear power and nuclear weapons. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-28427-1 (paperback). xvii + 412 pp. LCCN QC792 .C4713 2002. URL http:

//catdir.loc.gov/catdir/bios/uchi051/2002027143.html;
http://catdir.loc.gov/catdir/description/uchi051/2002027143.
html;   http://catdir.loc.gov/catdir/toc/uchi051/2002027143.
html.

Gerjuoy:1998:LST

[Ger98] E. Gerjuoy. Leo Szilard. Toward a livable world. APS April Meeting Abstracts, ??(??):405–??, April 1998. URL http://adsabs.

harvard.edu/abs/1998APS..APR..D405G. American Physical So-

ciety, APS/AAPT Joint April Meeting, April 18-21, 1998 Columbus, Ohio, abstract #D4.05.

Gerjuoy:1999:LST

[Ger99] Edward Gerjuoy. Leo Szilard. Toward a livable world. Physics and Society, 28(1):??, January 1999. URL http://flux.aps.org/

meetings/YR98/BAPSAPR98/abs/S395005.html;   http://www.
aps.org/units/fps/newsletters/1999/january/ajan99.html#
a3.

Goldhaber:1939:RIN

[GHS39] M. Goldhaber, R. D. Hill, and Leo Szilard. Radioactivity induced by nuclear excitation I. excitation by neutrons. Physical Review, 55(1):47–59, January 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v55/i1/p47_1. Reprinted in [FS87, pages 155–

Gibson:2019:SIH

[Gib19] Susannah Gibson. The Spirit of Inquiry: How One Extraordinary Society Shaped Modern Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2019. ISBN 0-19-883337-7 (hardcover). xxi + 377 pp. LCCN Q41.C194 G537 2019.

Goldberg:1974:BRB

[Gol74] Stanley Goldberg. Book review: The Collected Works of Leo Szilard. Volume I: Scientific Papers by Bernard T. Feld; Gertrud Weiss Szilard. Isis, 65(3):425–426, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228993.

Goldberg:1992:GSO

[Gol92a] Stanley Goldberg. Groves, Szilard, and Oppenheimer. Bulletin of the Atomic Scientists, 48(10):37, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1992:GTR

[Gol92b] Stanley Goldberg. Groves takes the reins. Bulletin of the Atomic Scientists, 48(10):32–39, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1992:ICO

[Gol92c] Stanley Goldberg. Inventing a climate of opinion: Vannevar Bush and the decision to build the bomb. Isis, 83(3):429–452, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233904.

Goujon:1996:HEA

[Gou96] P. Goujon. Historical and epistemological analyses of relations between entropy, neguentropy and information — from Maxwell’s and Szilard’s imaginary experiences to Brillouin’s theory. Cybernetica, 39(3):167–200, ???? 1996. ISSN 0011-4227.

Grodzins:1963:AAS

[GR63] Morton Grodzins and Eugene I. Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Basic Books, New York, NY, USA, 1963. xviii + 616 pp. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Grandy:1994:LSS

[Gra94] David Grandy. Leo Szilard: Science as a mode of being. Ph.D. thesis, Indiana University, Bloomington, IN, USA, 1994. 312 pp.

URL http://search.proquest.com/docview/304128624.

Grandy:1996:LSS

[Gra96] David Grandy. Leo Szilard: science as a mode of being. University Press of America, Lanham, MD, USA, 1996. ISBN 0-7618-0308-4 (hardcover). xiii + 189 pp. LCCN QC16.S95 G73 1996. Based on the author’s doctoral dissertation [Gra94].

Groves:1946:RMD

[Gro46] Leslie R. Groves. Recommendation for military decoration (Dr. Leo Szilard). Memorandum EIDM-WL-26, Oak Ridge Laboratory., July 8, 1946.

Groetsch:1996:TIP

[Gro96] C. W. Groetsch. Tartaglia’s inverse problem in a resistive medium. American Mathematical Monthly, 103(7):546–551, August/September 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2974665.

Gruber:1983:MPM

[Gru83] Carol S. Gruber. Manhattan Project maverick: The case of Leo Szilard. Prologue — Quarterly of the National Archives, 15(2): 73–87, Summer 1983.

Griffiths:1937:GRE

[GS37] J. H. E. Griffiths and Leo Szilard. Gamma rays excited by capture of neutrons. Nature, 139(3512):323–324, February 20, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v139/

n3512/pdf/139323b0.pdf. Reprinted in [FS87, pages 153–154].

Gerstell:1950:HSA

[GS50] Richard Gerstell and Leo Szilard. How to survive an atomic bomb, volume 845 of A Bantam book. Bantam Books, New York, NY, USA, 1950. 149 pp. LCCN UF767 .G39.

Guston:2012:PTM

[Gus12] David H. Guston. The pumpkin or the tiger? Michael Polanyi, Frederick Soddy, and anticipating emerging technologies. Minerva, 50(3):363–379, September 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11024-012-9204-8.pdf.

Gwertzman:1963:WSY

[Gwe63] Bernard Gwertzman. “The world has six years to live”: Leo Szilard, who helped invent the atomic bomb, now is trying desperately to save the world from it. Do his programs make any sense? Are his predictions to be taken seriously? Sunday: The Washington Star,

??(??):??, March 24, 1963. URL http://library.ucsd.edu/dc/
object/bb8807065d.

Hewlett:1962:B

[HA62] Richard G. Hewlett and Oscar E. Anderson, Jr. In the beginning [and other chapters]. In The New World: 1939–1946, pages 13– 29, 56, 178–180, 332–334, 341–342, 355, 366, 399, 440–442, 454, 668–669, 759. Pennsylvania State University Press, State College,

PA, USA, 1962. URL http://library.ucsd.edu/dc/object/
bb6554482m.

Haberer:1969:PCS

[Hab69] Joseph Haberer. Politics and the Community of Science. Van Nostrand Reinhold, New York, NY, USA, 1969. vi + 337 pp. LCCN Q125 .H23 1969.

Hafemeister:1996:SLS

[Haf96] D. Hafemeister. Szilard Lecture: In search of an improved science and public-policy process. APS Meeting Abstracts, ??(??):401–

??, May 1996. URL http://adsabs.harvard.edu/abs/1996APS.

.MAY..G401H. American Physical Society, APS/AAPT Joint Meeting, May 2-5, 1996, abstract #G4.01.

Hannigan:1993:BRB

[Han93] Paul Hannigan. Book review: Genius in the Shadows by William Lanouette with Bela Silard. Harvard Review, 261(4):226, Spring

1993. ISSN 1077-2901. URL http://www.jstor.org/stable/
27559870.

Hartshorne:1965:LES

[Har65] Charles Hartshorne. Letter to the Editor: Szilard and the realities of life. Bulletin of the Atomic Scientists, 21(5):35, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hargittai:2006:MSF

[Har06] István Hargittai. The Martians of Science: five physicists who changed the twentieth century. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-517845-9 (hardcover). xxiv + 313 + 32 pp. LCCN QC15 .H27 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0636/2005029427-d.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
2005029427-b.html;   http://www.loc.gov/catdir/toc/ecip061/
2005029427.html.

Hargittai:2010:JET

[Har10] István Hargittai. Judging Edward Teller: a closer look at one of the most influential scientists of the Twentieth Century. Prometheus Books, Amherst, NY, USA, 2010. ISBN 1-61614-221-9 (hardcover). 575 pp. LCCN QC16.T37 H37 2010. Afterword by Richard Garwin.

Hawkins:19xx:LSM

[Hawxx] Helen Hawkins. Leo Szilard: The man behind the bomb, a postscript with Gertrud Weiss Szilard. Undated interview broadcast on KPBS-TV, San Diego, CA, USA, and held in the Leo Szilard papers archive., 19xx.

Hawkes:2013:BEM

[Haw13] Peter W. Hawkes, editor. The Beginnings of Electron Microscopy, volume 16 of Advances in electronics and electron physics. Supplement. Academic Press, New York, USA, 2013. ISBN 0-12-014578-2, 1-4832-8465-4 (e-book). xix + 633 pp. LCCN QH212.E4.

Herbig:1976:KDA

[Her76] Jost Herbig. Kettenreaktion: das Drama der Atomphysiker. (German) [Nuclear reaction: the drama of the nuclear physicist]. Hanser-Verlag, München, Germany, 1976. ISBN ???? 514 + 8 pp. LCCN ???? Neumann János (1903–1957), Teller Ede (1908– 2003), Szilárd Leó (1898-1964), Wigner Jenő (1902–1995).

Herken:1993:GSB

[Her93] G. Herken. Genius in the shadows — a biography of Leo Szilard — the man behind the bomb. Nature, 362(6421):661, April 15, 1993. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Herschbach:1998:SDH

[Her98] Dudley Herschbach. Szilárd és delfinjei. (Hungarian) [Szilard and his dolphin]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9804/

tart9804.html. Second special issue on the centenary of Leo Szilard’s birth.

Hewlett:1994:BRB

[Hew94a] Richard G. Hewlett. Book review: Genius in the Shadows: A Biography of Leo Szilard — The Man behind the Bomb by William Lanouette; The Rickover Effect: How One Man Made a Difference by Theodore Rockwell. Technology and Culture, 35(2):442–444, April 1994. ISSN 0040-165X (print), 1097-3729 (electronic). URL

http://www.jstor.org/stable/3106333.

Hewlett:1994:BRW

[Hew94b] Richard G. Hewlett. Book review: William Lanouette, Genius in the Shadows: A Biography of Leo Szilard – the Man behind the Bomb. Technology and Culture, 35(2):442–??, ???? 1994. ISSN 0040-165X (print), 1097-3729 (electronic).

Hawkins:1987:TLW

[HGS87] Helen S. Hawkins, G. Allen Greb, and Gertrud Weiss Szilard. Toward a livable world: Leo Szilard and the crusade for nuclear arms control, volume 3 of Collected works of Leo Szilard. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-19260-8, 0-262-08162-8. lxxiv + 499 pp. LCCN QC3 .S97 vol. 3.

Halasz:1974:LSR

[HH74] Nicholas Halász and Robert Halász. Leo Szilard, the reluctant father of the atom bomb. The New Hungarian Quarterly, 15(55): 163–173, Autumn 1974. ISSN 0028-5390. URL http://library.

ucsd.edu/dc/object/bb75101231.

Howes:1999:TDS

[HHW99] Ruth (Ruth Hege) Howes, Caroline L. Herzenberg, and Ellen C. Weaver. Their day in the sun: women of the Manhattan Project. Labor and social change. Temple University Press, Philadelphia, PA, USA, 1999. ISBN 1-56639-719-7 (hardcover), 1-59213-192-1 (paperback), 0-585-38881-4 (e-book). viii + 264 pp. LCCN QC773.3.U5 H68 1999.

Higgins:1960:SBO

[Hig60] Marguerite Higgins. Szilard bars operation to work for peace. Sees 803 chance of nuclear war. New York Herald Tribune, ??(??):??, April 6, 1960.

Horvath:1997:TKE

[Hor97] T. Horvath. Theodore Kármán, [Eugene] Paul Wigner, John [von] Neumann, Leo Szilard, Edward Teller and their ideas of ultimate reality and meaning. Ultimate Reality and Meaning, 20(2–3):123– 146, June/September 1997. ISSN 0709-549X.

Hatano:1998:NSS

[HS98] T. Hatano and S. Sasa. Numerical simulations on Szilard’s engine and information erasure. Progress of Theoretical Physics, 100 (4):695–702, October 1998. CODEN PTPKAV. ISSN 0033-068x (print), 1347-4081 (electronic).

Hemmo:2011:SPM

[HS11] Meir Hemmo and Orly Shenker. Szilard’s Perpetuum Mobile. Philosophy of Science, 78(2):264–283, April 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/659231.

Hughes:1980:SLH

[Hug80] P. S. Hughes. Leo Szilard — his version of the facts. Contemporary Sociology — a Journal of Reviews, 9(4):551–552, ???? 1980. ISSN 0094-3061 (print), 1939-8638 (electronic).

Hutchins:1961:BBV

[Hut61] Robert M. Hutchins. Books: The Voice of the Dolphins, by Leo Szilard. Bulletin of the Atomic Scientists, 17(7):290, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Infeld:1970:LIB

[Inf70] Eryk Infeld. Leopold Infeld bibliography. General Relativity and Gravitation, 1(2):191–208, June 1970. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic). URL https://link.

springer.com/article/10.1007/BF00756897.

Irwin:1961:LDS

[Irw61] Theodore Irwin. The legend that is Dr. Szilard. Pageant, ??(??):

1, 53–59, December 1961. URL http://library.ucsd.edu/dc/
object/bb4916246x;   http://library.ucsd.edu/dc/object/
bb6554482m.

Jasper:1993:BRB

[Jas93] James M. Jasper. Book review: AEC and critics: Containing the Atom: Nuclear Regulation in a Changing Environment, 1963–1971, by J. Samuel Walker. Science, 259(5097):996–997, February 12, 1993. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/2880630.

Jeanloz:2009:LSL

[Jea09] R. Jeanloz. Leo Szilard Lectureship Award talk: Science and international security. APS April Meeting Abstracts, ??(??):7001–

??, May 2009. URL http://adsabs.harvard.edu/abs/2009APS.

.APR.R7001J. American Physical Society, 2009 APS April Meeting, May 2-5, 2009, abstract #R7.001.

Jenkin:2011:AEM

[Jen11] John G. Jenkin. Atomic energy is “moonshine”: What did Rutherford really mean? Physics in Perspective (PIP), 13(2):128–145, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/10.
1007/s00016-010-0038-1.

Jones:2010:THR

[Jon10] Derry W. Jones. Titanic Hungarian refugee physicists. Contemporary Physics, 51(3):267–271, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joravsky:1979:SSB

[Jor79] David Joravsky. Scientists as servants: The Physicists: The History of a Scientific Community in Modern America by Daniel J. Kevles. Scientists in Power by Spencer R. Weart. Leo Szilard: His Version of the Facts edited by Spencer R. Weart, edited by Gertrud Weiss Szilard. Science in a Free Society by Paul Feyerabend. Against Method: Outline of an Anarchistic Theory of Knowledge by Paul Feyerabend. New York Review of Books, 26(11):34–39, June 28, 1979. ISSN 0028-7504 (print), 1944-7744 (electronic).

URL http://www.nybooks.com/articles/archives/1979/jun/
28/scientists-as-servants/.

Kathren:1975:LP

[Kat75] Ronald L. Kathren. Letters: Also present. Bulletin of the Atomic Scientists, 31(1):3, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The author notes that Carl C. Gamertsfelder was omitted from the list of participants at the first test of the Chicago pile (CP-1) on 2 December 1942 given in [And74b, And74a].

Kelly:2002:LSL

[Kel02] H. Kelly. Leo Szilard Lectureship Award: The public obligation of scientists in the 21st Century. APS April Meeting Abstracts, pages

I6002–??, April 2002. URL http://adsabs.harvard.edu/abs/

2002APS..APR.I6002K. American Physical Society, April Meeting, Jointly Sponsored with the High Energy Astrophysics Division (HEAD) of the American Astronomical Society April 20 - 23, 2002 Albuquerque Convention Center Albuquerque, New Mexico Meeting ID: APR02, abstract #I6.002.

Kelly:2004:RMP

[Kel04] Cynthia C. Kelly, editor. Remembering the Manhattan Project: Perspectives on the Making of the Atomic Bomb and Its Legacy. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2004. ISBN 981-256-040-8. LCCN QC773.A1 R46 2004.

Kenworthy:1954:DHB

[Ken54] E. W. Kenworthy. The drama of the hydrogen bomb — and Dr. Oppenheimer’s key role: Security case focuses attention on disputes that preceded first successful test of H-bomb at Pacific Proving Ground. New York Times, ??(??):E5, April 18, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL https://search.proquest.com/hnpnewyorktimes/
docview/113104238/.

Kish:2012:ERC

[KG12a] L. B. Kish and C. G. Granqvist. Energy requirement of control: Comments on Szilard’s engine and Maxwell’s demon. Europhysics Letters, 98(6):68001, June 2012. CODEN EULEEJ. ISSN 0295-5075 (print), 1286-4854 (electronic).

Kish:2012:EMD

[KG12b] Laszlo Bela Kish and Claes-Goran Granqvist. Electrical Maxwell demon and Szilard engine utilizing Johnson noise, measurement, logic and control. PLoS One, 7(10):e46800, October 15, 2012. CO-DEN POLNCL. ISSN 1932-6203.

Kim:2011:ITF

[KK11] Kang-Hwan Kim and Sang Wook Kim. Information from timeforward and time-backward processes in Szilard engines. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 84(1 (part 1)):012101, July 8, 2011. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic).

Kim:2012:SIH

[KK12] Kang-Hwan Kim and Sang Wook Kim. Szilard’s information heat engines in the deep quantum regime. Journal of the Korean Physical Society, 61(8):1187–1193, October 2012. CODEN JKPSDV. ISSN 0374-4884 (print), 1976-8524 (electronic).

Klein:1987:AOH

[Kle87] Georg Klein. Ateisten och den heliga staden: möten och tankar. (Swedish) [The atheist and the holy city: encounters and reflections]. Bonnier, Stockholm, Sweden, 1987. ISBN 91-0-047194-1. 248 pp. LCCN ???? Includes two chapters on Leo Szilard for the period 1955–1964.

Klein:1990:AHC

[Kle90] George Klein. The atheist and the holy city: encounters and reflections. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-11041-5, 0-262-11155-1. xiv + 223 pp. LCCN RC265 .K5413 1990. Includes two chapters on Leo Szilard for the period 1955–1964.

Gyorgy:1998:TSL

[Kle98] Klein György. Találkozások Szilárd leóval. (Hungarian) [Szilard plays chess with death]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9804/

tart9804.html. Second special issue on the centenary of Leo Szilard’s birth.

Klein:2005:SFO

[Kle05a] Georg Klein. Skapelsens fullkomlighet och livets tragik: essäer. (Swedish) [Creation’s perfection and life’s tragedy: essays]. Bonnier, Stockholm, Sweden, 2005. ISBN 91-0-010786-7. 241 + 1 pp. LCCN ????

Kleppner:2005:LSR

[Kle05b] Daniel Kleppner. Leo Szilard and the role of physicists in countering nuclear threats. APS April Meeting Abstracts, pages 6003–??,

April 2005. URL http://adsabs.harvard.edu/abs/2005APS..

APR.T6003K. American Physical Society, APS April Meeting, April 16-19, 2005, abstract #T6.003.

Kozloff:1954:DGP

[KNSB54] L. Kozloff, Alvin Novick, Leo Szilard, and Edgar John Boell, editors. Dynamics of growth processes: Papers presented at the society’s 11th annual symposium held at Williams College in June 1952. Princeton University Press, Princeton, NJ, USA, 1954. LCCN ????

Konuma:1998:TSL

[Kon98] Michiji Konuma. Tisztelet Szilárd Leónak. (Hungarian) [Honoring Leo Szilard]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9804/

tart9804.html. Second special issue on the centenary of Leo Szilard’s birth.

Kramish:1993:PHE

[Kra93] Arnold Kramish. Powers on Heisenberg: Embellishments on the Lesart. American Scientist, 81(??):479–480, ???? 1993. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

Krakovsky:2021:NTH

[Kra21] Marina Krakovsky. News: Taking the heat: Maxwell’s demon and the high cost of erasure. Communications of the ACM, 64(6):18–20, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL https://dl.acm.org/doi/10.1145/3460214.

Kim:2011:QSE

[KSDU11] Sang Wook Kim, Takahiro Sagawa, Simone De Liberato, and Masahito Ueda. Quantum Szilard engine. Physical Review Letters, 106(7):070401, ???? 2011. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Kubbig:1998:KIK

[Kub98] Bernd W. Kubbig. Kommunikatoren im Kalten Krieg: Die Pugwash-Konferenzen, die Amerikanisch-Sowjetische Studiengruppe zur Rüstungskontrolle und die Grundlegung des ABM-Vertrages. (German) [Communicators in the Cold War: The Pugwash Conferences, the Soviet–American Study Group on Arms Control and the foundation of the ABM Treaty]. Amerikastudien / American Studies, 43(2):197–228, ???? 1998. ISSN 0340-2827.

URL http://www.jstor.org/stable/41157367.

Langer:1964:SPC

[Lan64] E. Langer. Scientists in politics: Council founded by Szilard brings cash and sophistication to lobbying. Science, 145(3632):561–563, August 7, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Lanouette:1983:DM

[Lan83] William Lanouette. Dream machine. Atlantic Monthly, ??(??): 35–52, 85–86, April 1983. ISSN 1072-7825 (print), 2151-9463 (electronic). A 20-page cover story about the history of the nuclear breeder reactor, a power plant designed to make more fuel than it consumes, including Szilard’s role both devising and naming the nuclear “chain reaction” in 1933 and the “breeder” in 1943. Letters and author’s replies in June, pages 6-7; July, pages 6 & 8; and August.

Lanouette:1985:AE

[Lan85] William Lanouette. Atomic energy, 1945–1985. The Wilson Quarterly (1976–2012), 9(5):90–131, Winter 1985. ISSN 0363-3276.

URL http://www.jstor.org/stable/40256964. A 43-page his-

tory of nuclear power’s past and potential, citing Szilard’s pivotal role in conceiving the nuclear chain reaction and creating the first nuclear reactor.

Lanouette:1986:LSA

[Lan86a] William Lanouette. Leo Szilard and the art of citizens’ diplomacy. Discussant at a session on “Citizens’ Diplomacy and International Exchanges” at the Sixth World Congress of International Physicians for the Prevention of Nuclear War. Cologne, Germany, May 31, 1986.

Lanouette:1986:BNG

[Lan86b] William Lanouette. Von des Bombes aus Netz. (German) [From the bomb to the grid]. History of nuclear power in America. WL interviewed in Washington, DC, Three Mile Island PA, and Shippingport, PA. RIAS-TV., 1986.

Lanouette:1988:WLH

[Lan88] William Lanouette. William Lanouette on his Leo Szilard biography. The New Hungarian Quarterly, XXIX(111):160–170, Autumn 1988. ISSN 0028-5390. Conversation with Gabor Pallo.

Lanouette:1989:BTB

[Lan89a] William Lanouette. Bumbling toward the bomb. Bulletin of the Atomic Scientists, 45(7):7–11, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). An article about Szilard, Einstein, and other refugee physicists alerting a suspicious U.S. Government about nuclear fission and nuclear weapons. Contains reproduction of August 2, 1939 Einstein/Szilard letter to President Franklin D. Roosevelt.

Lanouette:1989:PSL

[Lan89b] William Lanouette. Portrévázlat Szilard Leorol. Magyar Tudomány [Hungarian Science], ??(??):639–648, 1989.

Lanouette:1989:PS

[Lan89c] William Lanouette. The power and the story. Interview with WL about the development of nuclear power in America. Radio Dialogue #48., October 9, 1989.

Lanouette:1989:VCL

[Lan89d] William Lanouette. The varied career of Leo Szilard. Rockefeller Archive Center Newsletter, ??(??):7–??, Summer 1989. An article about Szilard‘s many affiliations with the Rockefeller Foundation and University, as revealed in the Rockefeller Archive Center holdings. Items included Szilard’s physics research in New York in the 1930s, settling refugee scholars, friendships with Einstein and von Neumann, information theory, microbiology, nuclear power, and arms-control initiatives.

Lanouette:1990:RMP

[Lan90] William Lanouette. Running down the Manhattan Project: Leo Szilard’s efforts to build and ban the bomb. A talk for the Seminar “Running the Manhattan Project: Conant, Groves, and Szilard” at the annual meeting of the History of Science Society. Seattle, WA, USA, October 26, 1990.

Lanouette:1992:ISP

[Lan92a] William Lanouette. Ideas by Szilard, physics by Fermi. Bulletin of the Atomic Scientists, 48(10):16–23, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special edition for the 50th Anniversary of the Chain Reaction (2 December 1942). An article about Leo Szilard’s role as co-designer of the world’s first nuclear reactor, and his invention of atomic secrecy.

Lanouette:1992:LSP

[Lan92b] William Lanouette. Leo Szilard and post-war science: From nuclear physics to molecular biology. A talk questioning if Szilard’s scientific style employs evolution, revolution, or subversion. The History of Science Society and American Historical Association, Washington, DC, USA, December 29, 1992.

Lanouette:1993:AWL

[Lan93a] William Lanouette. Author William Lanouette talks about Leo Szilard. Interview by Neal Conan on “Morning Edition,” National Public Radio, March 1, 1993.

Lanouette:1993:ALB

[Lan93b] William Lanouette. Author’s lecture and booksigning: World War II series. Presentation about Szilard’s attempts to prevent the atomic bombing of Japan, with readings from Genius in the Shadows The National Archives, Washington, DC., August 5, 1993.

Lanouette:1993:AL

[Lan93c] William Lanouette. Authors’ lunch. In a discussion among writers about Hungary with President Arpad Goncz, former US Ambassador Philip Kaiser, and others. Embassy of Hungary, Washington, DC, USA, April 23, 1993.

Lanouette:1993:CIG

[Lan93d] William Lanouette. Consultant and interviewer for “The Genius Behind the Bomb”. PBS/NOVA documentary about Leo Szilard narrated by Ed Asner. [NOVA Program #1911]., September 29, 1993.

Lanouette:1993:FSG

[Lan93e] William Lanouette. Faculty seminar on genius in the shadows. Talk at The George Washington University, Washington, DC, USA., March 2, 1993.

Lanouette:1993:GS

[Lan93f] William Lanouette. Genius in the shadows. Interview on the Voice of America about Szilard’s efforts to prevent the atomic bombing of Japan and the scientists’ petition to Truman., August 6, 1993.

Lanouette:1993:HGT

[Lan93g] William Lanouette. Hungarian geniuses and their bomb. Seminar on Science and Social Responsibility, about Szilard, Teller, von Neumann, and Wigner. Georgetown University, Washington, DC, USA., May 10, 1993.

Lanouette:1993:LSM

[Lan93h] William Lanouette. Leo Szilard: The man behind the award. American Physical Society, Szilard Award ceremony, Washington, DC, USA, April 14, 1993.

Lanouette:1993:WFM

[Lan93i] William Lanouette. Where the four Martians of the Manhattan Project landed: The national laboratory connections of Teller, Wigner, von Neumann, and Szilard. Talk at the American Physical Society, Washington, DC, USA., April 13, 1993.

Lanouette:1994:AS

[Lan94a] William Lanouette. Atomic spies. Debate with the authors of Special Tasks by Soviet spymaster Pavel Sudoplatov [SSSS95], which falsely alleged that Szilard, Niels Bohr, J. Robert Oppenheimer and Enrico Fermi were Soviet agents within the Manhattan Project. MacNeil/Lehrer NewsHour, April 26, 1994.

Lanouette:1994:LSL

[Lan94b] William Lanouette. Leo Szilard’s life. A talk to the Inter-Agency Seminar Group at the Brookings Institution, Washington, DC., March 23, 1994.

Lanouette:1994:SLS

[Lan94c] William Lanouette. Seminar on Leo Szilard. History Associates, Rockville, Maryland, May 6, 1994.

Lanouette:1995:A

[Lan95a] William Lanouette. August 1945. Woodrow Wilson Center “Dialogue” radio interview on the decision to drop A-bombs on Japan, and the Manhattan Project Scientists (including Szilard) who opposed using the weapons they had created. Radio Dialogue #295., May 1–7, 1995.

Lanouette:1995:GSB

[Lan95b] William Lanouette. Genius in the shadows: The biography of Leo Szilard. Woodrow Wilson Center “Dialogue” radio interview about Szilard’s efforts to create and control the A-bomb, and his life-long commitment to social issues. Radio Dialogue #306., July 17–23,

Lanouette:1995:HWB

[Lan95c] William Lanouette. Hiroshima: Why the bomb was dropped. Interview about Szilard’s attempts to demonstrate the A-bomb and stop its use on cities. ABC Peter Jennings Reporting., July 27, 1995.

Lanouette:1995:SSB

[Lan95d] William Lanouette. Scientists, statesmen, and the bomb. A talk about Leo Szilard’s efforts to prevent the atomic bombing of Japan, presented at the International Colloquium on the 50th Anniversary of Hiroshima, University of Mons-Hainaut, Belgium, September 15, 1995.

Lanouette:1995:WWD

[Lan95e] William Lanouette. Why we dropped the bomb. Civilization (The Magazine of the Library of Congress), 2(1):28–38, January/ February 1995. CODEN ???? ISSN 1077-9795. An article about how and why the United States decided to use atomic bombs against Japan, including the conflict between Leo Szilard (con) and Presidential advisor James F. Byrnes (pro). Reprinted in [Lan96e].

Lanouette:1996:LSP

[Lan96a] William Lanouette. Leo szilard: Physicist and peacemaker. A talk prepared for “Hungarian Emigre Scientists and the Technologies of Peace and War 1919-1989,” a workshop organized for the 23rd Symposium of the International Committee for the History of Technology, Budapest 8-9 August 1996. Published in The Martians: Hungarian Emigre Scientists and the Technology of War and Peace (Gyorgy Marx, editor). Budapest: Eötvös University, 1997, 88-95., August 8–9, 1996.

Lanouette:1996:NJP

[Lan96b] William Lanouette. A note on the July 17th Petition. In Bird and Lifschultz [BL96], pages 557–560. ISBN 0-9630587-3-8 (hardcover), 0-9630587-4-6 (paperback). LCCN 9609 BOOK NOT YET IN LC. A book chapter about Leo Szilard’s petition to President Truman by Manhattan Project scientists to prevent the atomic bombing of Japan in 1945.

Lanouette:1996:NAR

[Lan96c] William Lanouette. The nuclear arms race and the scientists behind it. Lecture about Szilard and fellow refugee scientists who created and continued the nuclear arms race. Visiting Scholar at James Madison University, February 5, 1996.

Lanouette:1996:TAS

[Lan96d] William Lanouette. Three attempts to stop the bomb. In Bird and Lifschultz [BL96], pages 99–118. ISBN 0-9630587-3-8 (hardcover), 0-9630587-4-6 (paperback). LCCN 9609 BOOK NOT YET IN LC. A book chapter about Leo Szilard’s efforts to prevent the atomic bombing of Japan in 1945.

Lanouette:1996:WWD

[Lan96e] William Lanouette. Why we dropped the bomb. In ????, editor, Annual Editions: World History, volume II, pages 159–165. Brown & Benchmark, Guilford, CT, USA, 1996. ISBN ???? LCCN ????

Lanouette:1996:WSB

[Lan96f] William Lanouette. Writing Szilard’s biography. Presentation to a Faculty Seminar, Loyola University, New Orleans, Louisiana, March 25, 1996.

Lanouette:1997:MBB

[Lan97a] William Lanouette. The man behind the bomb. In Lopez and Myers [LM97], pages 9–17. ISBN 0-8476-8594-2 (hardcover), 0-8476-8595-0 (paperback). LCCN JX1391 .P39 1997.

Lanouette:1997:NAR

[Lan97b] William Lanouette. The nuclear arms race: From Leo Szilard to stockpile stewardship. A seminar presented at the Peace Studies Program, Cornell University, Ithaca, New York., October 23, 1997.

Lanouette:1997:SLF

[Lan97c] William Lanouette. Szilard Leo: Fizikus es bekecsinalo. (Hungarian) [Leo Szilard: Physicist and peacemaker]. Fizikai Szemle (Budapest), ??(3):96.0–??, 1997. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.

hu/fszemle/archivum/fsz9703/lanoue.html. Hungarian trans-

lation of [Lan96a].

Lanouette:1997:SLZ

[Lan97d] William Lanouette. Szilard Leo: Zseni Arnyekban. (Hungarian) [Leo Szilard genius in the shadows]. Magyar Vilag Kiado, Budapest, Hungary, 1997. ISBN 963-9075-02-7. 439 pp. LCCN ???? Hungarian translation by Péter Hraskó of [LS92].

Lanouette:1998:LS

[Lan98a] William Lanouette. Leo Szilard. Interview with Veronika Sarkany for “Hungarian Journey: Leo Szilard.” Broadcast on Radio Hungary., April 25, 1998.

Lanouette:1998:LSCc

[Lan98b] William Lanouette. Leo Szilard: A comic and cosmic wit. Europhysics News, pages 92–93, 123, May/June 1998. CODEN EUP-NAS. ISSN 0531-7479 (print), 1432-1092 (electronic).

Lanouette:1998:LSCa

[Lan98c] William Lanouette. Leo Szilard: A comic and cosmic wit. A talk at the International Szilard Seminar, Leo Szilard Centenary Celebration, Eötvös University, Budapest, 9 February 1998. Published in [Lan98d]., February 9, 1998.

Lanouette:1998:LSCb

[Lan98d] William Lanouette. Leo szilard: A comic and cosmic wit. In Marx [Mar98b], pages 103–108. LCCN ????

Lanouette:1998:LSM

[Lan98e] William Lanouette. Leo Szilard memorial service and interment. Ithaca, New York, May 3, 1998.

Lanouette:1998:LSP

[Lan98f] William Lanouette. Leo Szilard: Physics, politics, and the narrow margin of hope. A talk at the Symposium in Celebration of the 100th Anniversary of Leo Szilard’s Birth. Sponsored by the Forum on International Physics, Forum on Physics and Society, and Forum on the History of Physics. American Physical Society annual meeting, Columbus, Ohio. Published in [Lan99c], April 18, 1998.

Lanouette:1998:NB

[Lan98g] William Lanouette. The Nazi bomb. Interview about Szilard’s attempts to spur the USA to develop an A-bomb before Germany could, including his drafting of a letter from Albert Einstein to President Roosevelt and his early research with Enrico Fermi. The History Channel. 10 June 1998., June 10, 1998.

Lanouette:1998:PCL

[Lan98h] William Lanouette. Prescience and conscience: Leo Szilard (1898– 1964). An adaptation of [Lan98d]., 1998.

Lanouette:1998:SGN

[Lan98i] William Lanouette. The shadow of genius in the nuclear age: Reflections on the life of Leo Szilard. Policy Forum with Nobel laureate Kenneth Wilson, Ohio State Univ., Columbus, Ohio., April 17, 1998.

Lanouette:1998:SLK

[Lan98j] William Lanouette. Szilárd Leó — komikus és kozmikus tréfamester. (Hungarian) [Leo Szilárd — a comic and cosmic wit]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://

www.kfki.hu/fszemle/archivum/fsz9804/tart9804.html. Sec-

ond special issue on the centenary of Leo Szilard’s birth.

Lanouette:1999:CUP

[Lan99a] William Lanouette. The century. ultimate power: The race to build the atomic bomb. Interview about Szilard’s efforts to stop the bomb he had initiated. ABC Peter Jennings Reporting., April 1, 1999.

Lanouette:1999:LSB

[Lan99b] William Lanouette. Leo Szilard and the A-bomb. In Chambers [Cha99], page ?? ISBN 0-19-507198-0. LCCN E181 .O94 1999.

Lanouette:1999:LSP

[Lan99c] William Lanouette. Leo Szilard: Physics, politics, and the narrow margin of hope. Physics and Society, 28(1):8–9, January 1999. CODEN ???? ISSN ???? URL http://flux.aps.org/

meetings/YR98/BAPSAPR98/abs/S395003.html;   http://www.
aps.org/units/fps/newsletters/1999/january/ajan99.html#
a2.

Lanouette:1999:RCH

[Lan99d] William Lanouette. Reason and circumstances of the Hiroshima bomb. A talk about the U.S. scientists’ struggle to control the A-bomb they had created, presented at a UNESCO Conference on World Cultural Heritage of the 20th Century “Modernity and Barbarism.” The Bauhaus, Dessau, Germany. 4 June 1999. Published in [Lan03]., June 4, 1999. ISBN 3-

936314-93-4. URL http://hsozkult.geschichte.hu-berlin.
de/rezensionen/2003-3-159.

Lanouette:1999:WFM

[Lan99e] William Lanouette. Where the four Martians of the Manhattan Project landed: The National Laboratory connections of Teller, Wigner, von Neumann, and Szilard. A Talk at a Conference on Laboratory History and Sociology, State University of New York at Stony Brook & Brookhaven National Laboratory, June 10, 1999.

Lanouette:2000:OCB

[Lan00a] William Lanouette. The odd couple and the bomb. Scientific American, 283(5):104–109, November 2000. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). An article about the Fermi–Szilard collaboration to create the world’s first nuclear reactor. Reprinted in [Lan01d, Lan01e].

Lanouette:2000:RND

[Lan00b] William Lanouette. Respondent to “New Dimensions in the Dialogue Between Science and the Humanities”. With examples of ways Szilard used science fiction and fables to reveal scientific and political insights. Gordon Research Conference on “New Frontiers in Science and Technology Policy” Plymouth, New Hampshire., August 22, 2000.

Lanouette:2001:GPS

[Lan01a] W. Lanouette. Goldhaber provided Szilard’s isotopes. Physics Today, 54(10):14, October 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Lanouette:2001:BC

[Lan01b] William Lanouette. Bomb cooperation. Scientific American (Japanese Edition), 31(2):114–121, February 2001. CODEN ???? ISSN ????

Lanouette:2001:BWP

[Lan01c] William Lanouette. Bombowa wspolpraca. (Polish) [Bomb cooperation]. ’Swiat Nauki, pages 70–75, May 2001. CODEN ???? ISSN 0867-6380 (print), 1689-0191 (electronic).

Lanouette:2001:FSE

[Lan01d] William Lanouette. Fermi, Szilard und der erste Atomreaktor. (German) [Fermi, Szilard and the first atomic reactor]. Spektrum der Wissenschaft (German translation of Scientific American), ?? (??):78–83, January 1, 2001. CODEN SPEKDI. ISSN 0170-2971.

Lanouette:2001:EPB

[Lan01e] William Lanouette. La extrana pareja y la bomba. (Spanish) [The odd couple and the bomb]. Investigacion y Ciencia, ??(??):18–23, January 2001. CODEN ???? ISSN 1665-4412.

Lanouette:2002:ELR

[Lan02a] William Lanouette. Einstein’s letter to Roosevelt. Consultant and speaker on Szilard’s role in drafting and delivering Einstein’s 1939 letter to FDR that prompted the Manhattan Project. The Learning Channel, Fall 2002.

Lanouette:2002:HV

[Lan02b] William Lanouette. A historian’s view. In Leventhal et al. [LTD02], pages 233–234. ISBN 1-57488-494-8 (hardcover), 1-57488-495-6 (paperback). LCCN TK9145 .N8297 2002. URL http://www.loc.

gov/catdir/toc/fy035/2002018416.html. A talk about the bal-

ance between nuclear power and nuclear weapons, featuring Leo Szilard’s inventive strategy for arms control.

Lanouette:2002:MMM

[Lan02c] William Lanouette. Modern marvels: Manhattan Project. Consultant and speaker on a program about Szilard’s role in prompting the U.S. A-bomb program, then opposing the weapon’s use. The History Channel, June 4, 2002.

Lanouette:2002:WP

[Lan02d] William Lanouette. War & peace. Speaker about Leo Szilard’s role rallying scientists against the weapons they created, in a film by Anand Patwardhan about the nuclear arms race and its control, 2002.

Lanouette:2003:GHB

[Lan03] William Lanouette. Grunde und Hintergrunde des Bombenabwurfs auf Hiroshima. (German) [reason and circumstances of the Hiroshima bomb]. In Prigge [Pri03], pages 187–195. ISBN 3-936314-

93-4. LCCN CB427 2003. URL http://hsozkult.geschichte.
hu-berlin.de/rezensionen/2003-3-159. A talk about the U.S.

scientists’ struggle to control the A-bomb they had created, presented at a UNESCO Conference on World Cultural Heritage of the 20th Century “Modernity and Barbarism.” The Bauhaus, Dessau, Germany. 4 June 1999. Article translated by Marie Neumullers.

Lanouette:2004:LSB

[Lan04a] William Lanouette. Leo Szilard: Baiting brass hats. In Kelly [Kel04], pages 73–77. ISBN 981-256-040-8. LCCN QC773.A1 R46 2004. A book chapter (based on a 2002 symposium talk) about Szilard’s efforts to start, accelerate, and then end the Manhattan Project; including his struggles for — then against — atomic secrecy.

Lanouette:2004:MML

[Lan04b] William Lanouette. Martian in motion: Leo Szilard. Consultant and speaker on a 2004 documentary about Leo Szilard. Arkion Documentary, Knoxville, Tennessee., 2004.

Lanouette:2004:UPH

[Lan04c] William Lanouette. Uranium + peaches in Hollywood. Interview about Leo Szilard and his portrayal by Ed Asner in a staged Hollywood reading. “Stage and Screen” KXLU-FM Los Angeles, March 22, 2004.

Lanouette:2005:AHN

[Lan05a] William Lanouette. 60th anniversary of Hiroshima & Nagasaki. Interview for TBS (Tokyo), NTV (Nagoya), and SPIEGEL-TV (Germany) programs about the atomic bombings of Hiroshima and Nagasaki and Szilard’s efforts first to start the U.S. A-bomb program — the Manhattan Project =– and then to prevent the bomb’s use against civilians., Spring/Summer 2005.

Lanouette:2005:EEL

[Lan05b] William Lanouette. Einstein’s equation of life & death. Consultant and speaker on the BBC “Horizon” science program about how a letter to the President by physicists Albert Einstein and Leo Szilard prompted the U.S. Government to develop nuclear weapons, and how the two men reacted once A-bombs were used. Broadcast in the UK 27 & 29 January 2005, on Discovery 17 April 2005., January 27 & 29, 2005.

Lanouette:2005:GVT

[Lan05c] William Lanouette. Geniĭ v teni (Russian) [Genius in the shadows]. Russian edition of Genius in the Shadows. Translated by Viktor B. Adamskii, Andrei Sakharov’s colleague in the Soviet nuclear weapons program, 2005., 2005.

Lanouette:2005:SSP

[Lan05d] William Lanouette. Szilardian science and politics: Evolution, revolution, or subversion? A paper for the Session on “The Scientist’s Approach –0 The Decision Makers’ Approach” that details Szilard’s unconventional creations in science and public policy. World Science Forum on “Knowledge, Ethics, and Responsibility,” sponsored by UNESCO and the Hungarian Academy of Sciences. Budapest, 10th November 2005. Presented in modified form at the Program on Science and Global Security, Woodrow Wilson School, Princeton University, 15 February 2006., November 2005.

Lanouette:2006:ELa

[Lan06a] William Lanouette. Einstein’s letter. “Science Talk,” the Scientific American Podcast Interview about an upcoming History Channel program on the roles of Albert Einstein and Leo Szilard in the beginning of the nuclear age. Episode 9., April 5, 2006.

Lanouette:2006:ELb

[Lan06b] William Lanouette. Einstein’s letter. Consultant & on-air speaker about how Szilard drafted for Albert Einstein a letter to President Franklin Roosevelt in 1939 warning about a German A-bomb. History Channel series “Ten Days That Unexpectedly Changed America.”, April 10, 2006.

Lanouette:2006:SPL

[Lan06c] William Lanouette. The science and politics of Leo Szilard, 1898– 1964: evolution, revolution, or subversion? Science and Public Policy, 33(8):613–617, October 2006. ISSN 0302-3427 (print),

1471-5430 (electronic). URL http://spp.oxfordjournals.org/
content/33/8/613.abstract.

Lanouette:2008:MPC

[Lan08a] William Lanouette. The Manhattan Project and its Cold War legacy. Woodrow Wilson International Center for Scholars., February 20, 2008.

Lanouette:2008:NMH

[Lan08b] William Lanouette. A narrow margin of hope: Leo Szilard in the founding days of CARA. A talk in the session “The Early Years: The Academic Assistance Council” at the conference “In Defence of Learning: The Past and the Present” celebrating the 75th anniversary of the Council for the Assistance of Academic Refugees (The Royal Society & The British Academy) London., December 4, 2008.

Lanouette:2009:BBF

[Lan09a] William Lanouette. Building the bomb, fearing its use: Nuclear scientists, social responsibility and arms control, 1946–1996. Panelist at a Library of Congress discussion on “Building the Bomb.” Described Szilard’s struggles to build and then ban the bomb. Other participants were Mary Palevsky, Hugh Gusterson, and Martin J. Sherwin, May 27, 2009.

Lanouette:2009:CCN

[Lan09b] William Lanouette. Civilian control of nuclear weapons. Arms Control Today, 39(4):45–48, May 2009. ISSN 0196-125X (print),

1943-5754 (electronic). URL http://www.armscontrol.org/act/

2009_5/Lanouette. An article about the proposal to shift nuclear weapons work from the Energy to the Defense Department, and the efforts that Szilard led in the 1940s to establish “civilian control” of nuclear weapons and research.

Lanouette:2009:SPA

[Lan09c] William Lanouette. Scientists and politicians: The art of the impossible. A paper presented for the panel on “Nuclear Weapons Policies: Perspectives from the Past” at the Symposium “Revisiting Reykjavik: Nuclear Weapons Policy for the New Century.” Carnegie Institution for Science, Washington, DC, USA, March 14, 2009.

Lanouette:2010:LSW

[Lan10] William Lanouette. Leo Szilard in Washington: Seeking a market for wisdom. Washington Magyar Club. Described Szilard’s efforts to start the US nuclear weapons program in 1939, and to curb the nuclear arms race both during and after World War II. Also detailed his research at NIH in the 1950s and his creation of the Council for a Livable World to aid arms control., January 16, 2010.

Lanouette:2011:ESP

[Lan11a] William Lanouette. Einstein and Szilard in Princeton. Historical Society of Princeton, The Lewis B. Cuyler Lecture. Described the friendship and interactions between Albert Einstein and Leo Szilard, in Berlin and in Princeton. Included their friendships and contacts with John von Neumann and Eugene Wigner. (Program included remarks by Freeman Dyson.), February 9, 2011.

Lanouette:2011:EF

[Lan11b] William Lanouette. Einstein’s fridge. Interview in a Discovery program on Inventions that Shook the World, describing the coinvention by Einstein and Szilard of a revolutionary refrigerator in the 1920s., July 27, 2011.

Lanouette:2011:EMB

[Lan11c] William Lanouette. Ethics, morality, and the bomb. Leo Szilard’s struggle with the weapon he created and then tried to control. The Woodstock Group., December 13, 2011.

Lanouette:2011:FHM

[Lan11d] William Lanouette. The Four Hungarians of the Manhattan Project. Washington Magyar Club. Described the roles of Leo Szilard, Eugene Wigner, John von Neumann, and Edward Teller in the Manhattan Project., January 15, 2011.

Lanouette:2011:NMH

[Lan11e] William Lanouette. A narrow margin of hope: Leo Szilard in the founding days of CARA. In Marks et al. [MWW11], page ?? ISBN 0-19-726481-6. LCCN HV640.4.G7 I55 2011. URL http:

//www.loc.gov/catdir/enhancements/fy1205/2011284500-b.
html;   http://www.loc.gov/catdir/enhancements/fy1205/
2011284500-d.html;   http://www.loc.gov/catdir/enhancements/

fy1205/2011284500-t.html. Published for the British Academy.

Lanouette:2012:EAS

[Lan12a] William Lanouette. European and American scientists and their creation of nuclear weapons and power. Graduate seminar presentation on Szilard’s efforts in Europe and the USA to create A-bombs and nuclear reactors, August 16, 2012.

Lanouette:2012:UP

[Lan12b] William Lanouette. Uranium + peaches. Staged dramatic reading by the Arts Council of Princeton and the Princeton Theatre Experiment at the Paul Robeson Center, Princeton, NJ, USA, March 10, 2012.

Lanouette:2012:WWD

[Lan12c] William Lanouette. Why we dropped the bomb. Talk to University of the Third Age at the University of San Diego, featuring Szilard’s efforts to stop the atomic bombing of Japan in 1945, July 19, 2012.

Lanouette:2013:GSB

[Lan13] William Lanouette. Genius in the Shadows: A Biography of Leo Szilard, the Man Behind the Bomb. Skyhorse Publishing, New York, NY, USA, 2013. ISBN 1-62636-023-5. ???? pp. LCCN ???? With Bela Silard.

Lanouette:2014:LSM

[Lan14a] William Lanouette. Leo Szilard: The man behind the bomb. Lecture at the University of California, San Diego., April 8,

2014. URL http://www.uctv.tv/shows/Leo-Szilard-The-Man-
Behind-the-Bomb-with-William-Lanouette-28013;   https:/
/www.youtube.com/watch?v=SVIIFBHhl6g.

Lanouette:2014:MWL

[Lan14b] William Lanouette. The many worlds of Leo Szilard: Physicist, biologist, and peacemaker. History of Physics Newsletter, 12(5): 4–5, Fall 2014. CODEN ???? ISSN ???? URL http://www.aps.

org/units/fhp/newsletters/fall2014/index.cfm.

Lanouette:2015:PPF

[Lan15] William Lanouette. Public policy forum ’15: Why we dropped the bomb. . .. KBUT radio Outside/In broadcast, Gunnison County, Colorado, USA., July 25, 2015. URL https:/

/soundcloud.com/kbut-radio/outsidein-william-lanouette?
in=kbut-radio/sets/local-headlines.

Lapp:1968:WC

[Lap68] Ralph Eugene Lapp. The weapons culture. W. W. Norton & Co., New York, NY, USA, 1968. 230 pp. LCCN UF767 .L273 1968.

Lapp:1969:WC

[Lap69] Ralph Eugene Lapp. The weapons culture. Pelican books. Penguin Books, Harmondsworth, UK, 1969. ISBN 0-14-021138-1. 236 pp. LCCN UF767 .L273 1969.

Lash:1972:WHY

[Las72] Joseph P. Lash. The White House years: The 1944 campaign. In Eleanor and Franklin: the story of their relationship, based on Eleanor Roosevelt’s private papers, pages 704–705, 707, 70? W. W. Norton & Co., New York, NY, USA, third edition, 1972. URL

http://library.ucsd.edu/dc/object/bb6554482m.

Laurence:1946:X

[Lau46] William L. Laurence. Dec. 2, 1942 — the birth of the Atomic Age: Story of the great experiment which first released the energy that runs the universe. New York Times, ??(??):SM6, SM60, December 1, 1946. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/107694608/.

Launius:1994:BRR

[Lau94] Roger D. Launius. Book review: Robert A. Divine: The Sputnik Challenge: Eisenhower’s Response to the Soviet Satellite. Isis, 85(1):173–174, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235949.

Lanouette:2002:UP

[LC02] William Lanouette and Peter Cook. Uranium + peaches. Play about the May 1945 meeting between Einstein’s protégé Leo Szilard and President Truman’s mentor (soon Secretary of State) James F. Byrnes, in which the scientist who made the A-bomb and wants to stop it confronts the politician eager to use it. Five staged dramatic readings in New York City and two in Hollywood, 2002–08., 2002.

URL http://www.uraniumandpeaches.com/.

Leverett:1942:DHC

[LCM^{+}42] M. C. Leverett, C. M. Cooper, T. V. Moore, E. P. Wigner, E. S. Steinbach, E. Fermi, J. A. Wheeler, S. K. Allison, and Leo Szilard. Discussion of helium cooled power plant. Report CS-267, US Atomic Energy Commission, Washington, DC, USA, September 16, 1942.

Lear:1959:THW

[Lea59] John Lear. A theory of how we age. New Scientist, ??(??): 346, February 12, 1959. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://library.ucsd.edu/

dc/object/bb2356495j.

Leff:2002:MDS

[Lef02] Harvey S. Leff. Maxwell’s Demon and the Second Law [of thermodyamics]. AIP Conference Proceedings, 643(1):408–??, 2002. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Lerner:1960:LM

[Ler60] Max Lerner. Life of a man. New York Post magazine, ??(??):

34, March 4, 1960. URL http://library.ucsd.edu/dc/object/
bb1059555z.

Leskov:1994:SFU

[Les94] Sergei Leskov. The Sudoplatov file: An unreliable witness. Bulletin of the Atomic Scientists, 50(4):33–36, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1961:SUP

[Lew61] Richard Lewis. Szilard urges people’s drive, lobby to work for peace. Chicago Sun-Times, ??(??):??, December 2, 1961.

Lewis:1994:BRG

[Lew94] H. W. Lewis. Book review: A genius among many: Genius in the Shadows: A Biography of Leo Szilard: The Man Behind the Bomb. American Scientist, 82(1):70–71, January/February 1994. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL

http://www.jstor.org/stable/29775104.

Lanouette:2002:LSB

[LHN02] William Lanouette, Gregg Herken, and Robert S. Norrisat. Leo szilard: Baiting brass hats. A talk for the panel on “The Military and Science in the Crucible of War” (with Gregg Herken and Robert S. Norrisat) a Symposium on the Manhattan Project, Carnegie Institution of Washington, Washington DC. Broadcast as C-SPAN Program # 169799-2., April 27, 2002.

Lifton:1980:PNW

[Lif80] Robert Jay Lifton. The prevention of nuclear war. Bulletin of the Atomic Scientists, 36(8):38–43, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BL81].

Livingston:1960:LSB

[Liv60] Robert Livingston. Leo Szilard — biographical notes. Available in the Leo Szilard Papers archive., February 21, 1960.

Lu:2012:PEPa

[LL12] Yao Lu and Gui Lu Long. Parity effect and phase transitions in quantum Szilard engines. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 85(1 (part 1)):

011125, ???? 2012. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic).

Lanouette:1995:MPS

[LLGS95] William Lanouette, Ralph E. Lapp, Howard Gest, and John A. Simpson. The Manhattan Project scientists petition President Truman. Lecture and Introduction to a Discussion by Three Signers of Leo Szilard’s 17 July 1945 Petition About the Decision to Use the Atom Bomb. World War II Program. The National Archives, Washington, DC. TV broadcast by C-SPAN., July 17, 1995.

Lopez:1997:PSN

[LM97] George A. Lopez and Nancy J. Myers, editors. Peace and security: the next generation. Rowman and Littlefield, Lanham, MD, USA, 1997. ISBN 0-8476-8594-2 (hardcover), 0-8476-8595-0 (paperback). xiv + 275 pp. LCCN JX1391 .P39 1997.

Logan:1993:BMB

[Log93] Jonothan L. Logan. Bomb maker or bomb breaker? Boston Globe, ??(??):B43–B45, March 7, 1993. ISSN 0743-1791. Review of [Pow93].

Lowen:1993:BRT

[Low93] Rebecca Lowen. Book review: Thinking in the bathtub: Genius in the Shadows, by William Lanouette with Bela Silard. Bulletin of the Atomic Scientists, 49(5):51–52, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1992:GSB

[LS92] William Lanouette and Bela A. Silard. Genius in the shadows: a biography of Leo Szilard: the man behind the bomb. C. Scribner’s Sons, New York, NY, USA, 1992. ISBN 0-684-19011-7. xix + 587 + 16 pp. LCCN QC16.S95 L36 1992. US$35.00 (US$44.50 Can.).

URL http://adsabs.harvard.edu/abs/1994gsbl.book.....L.

Lanouette:1994:GSB

[LS94] William Lanouette and Bela A. Silard. Genius in the shadows: a biography of Leo Szilard: the man behind the bomb. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-46888-7 (paperback). xix + 587 pp. LCCN QC16.S95 L36 1994. URL http:

//www.loc.gov/catdir/description/uchi053/94012738.html.

Leghorn:1958:MAV

[LSW58] Richard Leghorn, Leo Szilard, and Jerome B. Wiesner. Memorandum to Alexander V. Topchiev [on a proposed meeting of American and Russian scientists]. Technical report, ????, ????, April 6, 1958. Available in the Leo Szilard Papers archive.

Leventhal:2002:NPS

[LTD02] Paul Leventhal, Sharon Tanzer, and Steven Dolley, editors. Nuclear power and the spread of nuclear weapons: can we have one without the other? Brassey’s, Washington, DC, USA, 2002. ISBN 1-57488-494-8 (hardcover), 1-57488-495-6 (paperback). xxv + 340 pp.

LCCN TK9145 .N8297 2002. URL http://www.loc.gov/catdir/

toc/fy035/2002018416.html. Foreword by Edward J. Markey.

Lubkin:1987:KEM

[Lub87] E. Lubkin. Keeping the entropy of measurement — Szilard revisited. International Journal of Theoretical Physics, 26(6):523–535, June 1987. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Li:2012:RQS

[LZL^{+}12] Hai Li, Jian Zou, Jun-Gang Li, Bin Shao, and Lian-Ao Wu. Revisiting the quantum Szilard engine with fully quantum considerations. Annals of Physics, 327(12):2955–2971, December 2012. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Maas:2004:LSP

[Maa04] W. Maas. Leo Szilard: A personal remembrance. Genetics, 167 (2):555–558, June 2004. CODEN GENTAE. ISSN 0016-6731 (print), 1943-2631 (electronic). URL http://www.genetics.org/

content/167/2/555.full.

Mabley:1962:MRA

[Mab62] Jack Mabley. Mabley’s report: Atom expert’s plan to avert all-out war. Chicago’s American, ??(??):??, February 12 (or 19), 1962.

URL http://library.ucsd.edu/dc/object/bb6554482m.

Magnasco:1996:SHE

[Mag96] M. O. Magnasco. Szilard’s heat engine. Europhysics Letters, 33(8): 583–588, March 1996. CODEN EULEEJ. ISSN 0295-5075 (print),

1286-4854 (electronic). URL http://adsabs.harvard.edu/abs/
1996EL.....33..583M.

Magyar:1998:EHH

[Mag98] Magyar Bálint. Erasmus Hazatér. (Hungarian) [Erasmus returns home]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9804/

tart9804.html. Second special issue on the centenary of Leo Szilard’s birth.

Malloy:2008:BRP

[Mal08] Sean L. Malloy. Book review: P. D. Smith: Doomsday Men: The Real Dr. Strangelove and the Dream of the Superweapon. Isis, 99(4): 876–877, December 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.
1086/597738;   http://www.jstor.org/stable/pdfplus/10.
1086/597738.pdf.

Mann:1953:C

[Man53] W. B. (Wilfrid Basil) Mann. The Cyclotron. Methuen’s monographs on physical subjects. Methuen and Co. Ltd., London, UK, fourth edition, 1953. 118 pp. LCCN QC787.C8 M3 1953. URL

http://library.ucsd.edu/dc/object/bb84657622.

Mandelbrot:1965:LSU

[Man65] Benoît Mandelbrot. Leo Szilard and unique decipherability. IEEE Transactions on Information Theory, IT-11(3):455–456, ???? 1965. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Martin:1948:LEH

[Mar48] Abram V. Martin. Letter to the Editors: How to implement Dr. Szilard’s proposal to Mr. Stalin. Bulletin of the Atomic Scientists, 4(3):80, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marton:1968:EHE

[Mar68] L. (Ladislaus) Marton. Early history of the electron microscope. History of technology monographs. San Francisco Press, San Francisco, CA, USA, 1968. viii + 56 pp. LCCN QH212.E4 M35. With a preface by Dennis Gabor [Nobel Prize in Physics 1971 “for his invention and development of the holographic method”].

Marton:1994:EHE

[Mar94a] L. (Ladislaus) Marton. Early history of the electron microscope. History of technology monographs. San Francisco Press, San Francisco, CA, USA, second edition, 1994. ISBN 0-911302-73-5. viii + 55 pp. LCCN QH212.E4 M35 1994.

Marx:1994:VMQ

[Mar94b] George Marx. The voice of the Martians: the quest for nuclear chain reaction: a Hungarian version: (50 years after). Roland Eötvös Physical Society, Budapest, Hungary, 1994. ISBN 963-8051-62-0 (paperback). 213 + 16 pp. LCCN ????

Marx:1996:MMG

[Mar96] George Marx. The myth of the Martians and the golden age of Hungarian science. Science & Education (Springer), 5(3):225–234, July 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1996Sc%26Ed..
.5..225M.

Marx:1998:BLR

[Mar98a] George Marx. Beller Lecture: The roots of Leo Szilard and his interdisciplinarity. APS April Meeting Abstracts, pages 404–??,

April 1998. URL http://adsabs.harvard.edu/abs/1998APS..

APR..D404M. American Physical Society, APS/AAPT Joint April Meeting, April 18-21, 1998 Columbus, Ohio, abstract #D4.04.

Marx:1998:LSC

[Mar98b] George Marx, editor. Leo Szilárd Centenary Volume: lectures and contributions of the Conference at the Centenary of Szilard’s Birth, Budapest on 9–11 February 1998. Eötvös Physical Society, Budapest, Hungary, 1998. LCCN ????

Marx:1998:GHI

[Mar98c] Marx György. Gaia, a halandó istennő. (Hungarian) [Gaia, the mortal goddess]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9802/

tart9802.html. Special issue on the centenary of Leo Szilard’s birth.

Marx:1998:SLH

[Mar98d] Marx György. Szilárd Leó hazatért. (Hungarian) [Leo Szilard returned home]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9804/

tart9804.html. Second special issue on the centenary of Leo Szilard’s birth.

Marton:2006:GEN

[Mar06] Kati Marton. The great escape: nine Jews who fled Hitler and changed the world. Simon and Schuster, New York, NY, USA, 2006. ISBN 0-7432-6115-1 (hardcover), 0-7432-6116-X (paperback). 271 + 16 pp. LCCN DS135.H93 A153 2006. URL http://catdir.

loc.gov/catdir/enhancements/fy0665/2006049162-d.html;
http://catdir.loc.gov/catdir/enhancements/fy0668/2006049162-
s.html;   http://catdir.loc.gov/catdir/enhancements/fy0668/
2006049162-t.html;   http://catdir.loc.gov/catdir/enhancements/
fy0912/2006049162-b.html.

Michaudon:2000:FMW

[MB00] André F. Michaudon and Ileana G. Buican. A factor of millions: Why we made plutonium. Los Alamos Science, ??(26):4–9, November 2000. CODEN LASCDI. ISSN 0273-7116. URL http://www.

fas.org/sgp/othergov/doe/lanl/pubs/00818004.pdf.

McClaughry:1965:BRB

[McC65] John McClaughry. Book review: The Voice of the Dolphins. The Progressive, ??(??):26–29, April 1965.

McMillan:1994:SFF

[McM94] Priscilla Johnson McMillan. The Sudoplatov file: Flimsy memories. Bulletin of the Atomic Scientists, 50(4):30–33, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1947:BEB

[MCW47] Leslie Miller, Robert Considine, and Frank Wead. The Beginning or the End: Book of the Film. ????, ????, 1947. LCCN ???? The MGM film about the development of atomic weapons in the Manhattan Project has serious flaws; see [Rei84].

McDayter:1967:GBB

[MD67] Walt McDayter and Norman Drew. The giants: The bomb builders. Denver Post, ??(??):??, February 3, 1967. URL http://library.

ucsd.edu/dc/object/bb0103915g. This is a reasonably accurate

83-frame comic strip on the history of the building of the atomic bomb, with Leo Szilard as the central figure of the story.

Millet:2005:OPR

[Mil05] Lydia Millet. Oh pure and radiant heart. Soft Skull Press, Brooklyn, NY, USA, 2005. ISBN 1-932360-85-9. 489 pp. LCCN

PS3563.I42175 O37 2005. URL http://www.loc.gov/catdir/
toc/ecip056/2005001028.htm.

Moore:1942:ETD

[MLC^{+}42] T. V. Moore, M. C. Leverett, C. M. Cooper, E. S. Steinbach, and Leo Szilard. Engineering and technological divisions. report for month ending August 15, 1942. Report CE-236, US Atomic Energy Commission, Washington, DC, USA, 1942.

Moore:1992:ENCd

[Moo92] Mike Moore. Editor’s note: Consequences. Bulletin of the Atomic Scientists, 48(10):2, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Medawar:2001:HGT

[MP01] J. S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Marathe:2010:CCPa

[MP10] Rahul Marathe and J. M. R. Parrondo. Cooling classical particles with a microcanonical Szilard engine. Physical Review Letters, 104 (24):245704, June 18, 2010. CODEN PRLTAO. ISSN 1079-7114.

Mark:1925:EVA

[MS25] H. Mark and L. Szilard. Ein einfacher Versuch zur Auffindung eines selektiven Effektes bei der Zerstreuung von Röntgenstrahlen. (German) [A simple test for the detection of a selective effect on the scattering of X-rays]. Zeitschrift für Physik, 33(1): 688–691, December 1925. CODEN ZEPYAA. ISSN 0044-3328.

URL http://adsabs.harvard.edu/abs/1925ZPhy...33..688M;
 http://www.springerlink.com/content/h672566162427q81/.

Mark:1926:PRD

[MS26] H. Mark and L. Szilard. Die Polarisierung von Röntgenstrahlen durch Reflexion an Kristallen. (German) [The polarization of Xrays by reflection in crystals]. Zeitschrift für Physik, 35(19):743– 747, October 1926. CODEN ZEPYAA. ISSN 0044-3328. URL

http://adsabs.harvard.edu/abs/1926ZPhy...35..743M.

Marshall:1941:PRF

[MS41a] John Marshall, Jr. and Leo Szilard. Preliminary report on fission caused by fission neutrons. Report CP-316, US Atomic Energy Commission, Washington, DC, USA, November 14, 1941. Reprinted in [FS87, pages 266–275].

Marshall:1941:PRC

[MS41b] John Marshall, Jr. and Leo Szilard. Preliminary report on the capture of neutrons by uranium in the energy region of photo neutrons from radium–beryllium sources. Report CP-317, US Atomic Energy Commission, Washington, DC, USA, December 5, 1941. Reprinted in [FS87, pages 276–279].

Marshall:1942:TDR

[MS42] John Marshall, Jr. and Leo Szilard. Technological division. Report for month ending September 15, 1942. Report CE-271, US Atomic Energy Commission, Washington, DC, USA, 1942. Contents: Mg reduction of UF_{4}; melting furnaces and casting of uranium.

Marcus:2006:GFC

[MSHP06] Philip I. Marcus, Gordon H. Sato, Richard G. Ham, and David Patterson. The genesis of “feeder cells”: Concept and practice. In Vitro Cellular & Developmental Biology. Animal, 42(8/9):235– 237, September/October 2006. CODEN IVCAED. ISSN 1071-2690 (print), 1543-706X (electronic). URL http://www.jstor. org/stable/4295703. In memoriam tribute to Theodore T. Puck (24 September 1916–6 November 2005).

Murphy:1973:SPH

[Mur73] Patricia Murphy. Szilard papers in her trust. Los Angeles Times, ??(??):??, April 18, 1973.

Masters:1946:OWN

[MW46a] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb. Whittlesey House, McGraw-Hill Book Co., Inc., New York, NY, USA, 1946. x + 79 pp. LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Masters:1946:VEI

[MW46b] Dexter Masters and Katherine Way, editors. Een Verden eller ingen. (Danish) [One world or none]. ????, København, Danmark, 1946. 220 pp. LCCN ???? Danish translation of [MW46a] by Gudrun Frederiksen og Ebbe Rasmussen. Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46a].

Masters:2007:OWN

[MW07] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb. New Press, New York, NY, USA, 2007. ISBN 1-59558-227-4 (hardcover). xx + 220 pp. LCCN UG1282.A8 O54

2007. URL http://thenewpress.com/index.php?option=com_
title&task=view_title&metaproductid=1703;   http://www.
loc.gov/catdir/toc/ecip0718/2007020838.htm. Foreword by

Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46a].

Marks:2011:DLP

[MWW11] Shula Marks, Paul Weindling, and Laura Wintour, editors. In defence of learning: the plight, persecution, and placement of academic refugees, 1933-1980s, volume 169 of Proceedings of the British Academy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-726481-6. LCCN HV640.4.G7

I55 2011. URL http://www.loc.gov/catdir/enhancements/
fy1205/2011284500-b.html;   http://www.loc.gov/catdir/
enhancements/fy1205/2011284500-d.html;   http://www.loc.
gov/catdir/enhancements/fy1205/2011284500-t.html. Pub-

lished for the British Academy.

Nagy:1981:SFS

[Nag81] L. G. Nagy. Szilard fluid surface phenomena connected with research. Kemiai Kozlemenyek, 56(1–2):139–145, ???? 1981. ISSN 0022-9814.

Novick:1998:KBH

[ND98] Aaron Novick and Gene Dannen. Kalandozás a biológiában. (Hungarian) [Adventure in biology and ideas]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.hu/

fszemle/archivum/fsz9804/tart9804.html. Second special is-

sue on the centenary of Leo Szilard’s birth.

Neff:1997:LSA

[Nef97] T. Neff. Leo Szilard Award Lecture: Unwinding the cold war. APS April Meeting Abstracts, ??(??):702–??, April 1997. URL http:

//adsabs.harvard.edu/abs/1997APS..APR..G702N. American

Physical Society, APS/AAPT Joint Meeting, April 18–21, 1997, abstract #G7.02.

Neffe:2005:EBG

[Nef05] Jürgen Neffe. Einstein: eine Biographie. (German) [Einstein: a Biography]. Rowohlt, Reinbek bei Hamburg, Germany, second edition, 2005. ISBN 3-498-04685-3. 490 + 14 pp. LCCN QC16.E5 N44 2005.

Neffe:2006:EBG

[Nef06] Jürgen Neffe. Einstein: eine Biographie. (German) [Einstein: a Biography], volume 61937 of rororo. Rowohlt Taschenbuch Verlag, Reinbek bei Hamburg, Germany, 2006. ISBN 3-499-61937-7 (paperback). 490 + 8 pp. LCCN ???? EUR 9.90, sfr 17.90.

Neffe:2007:EB

[Nef07] Jürgen Neffe. Einstein: a biography. Farrar, Strauss, and Giroux, New York, NY, USA, 2007. ISBN 0-374-14664-0 (hardcover). x + 461 + 16 pp. LCCN QC16.E5 N4413 2007. URL http://www.loc.

gov/catdir/enhancements/fy0705/2006026136-b.. Translated

by Shelley Frisch from the original German edition Einstein: eine Biographie (ISBN 3-498-04685-3, 3-499-61937-7) [Nef05, Nef06].

Neffe:2011:AEA

[Nef11] Jürgen Neffe. Albert Einstein and the atomic bomb. History Reader, ??(??):??, August 6, 2011. URL http://www.

thehistoryreader.com/modern-history/albert-einstein-atomic-
bomb/.

Novick:1949:ELR

[NS49] Aaron Novick and Leo Szilard. Experiments on light-reactivation of ultra-violet inactivated bacteria. Proceedings of the National Academy of Sciences of the United States of America, 35(10):591– 600, October 1949. CODEN PNASA6. ISSN 0027-8424 (print),

1091-6490 (electronic). URL http://adsabs.harvard.edu/abs/
1949PNAS...35..591N;   http://www.pnas.org/content/36/

12/708.full.pdf+html. Reprinted in [FS87, pages 393–402].

Novick:1950:DC

[NS50a] Aaron Novick and Leo Szilard. Description of the Chemostat. Science, 112(2920):715–716, December 15, 1950. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:/

/adsabs.harvard.edu/abs/1950Sci...112..715N. Reprinted in

[FS87, pages 403–404].

Novick:1950:ECS

[NS50b] Aaron Novick and Leo Szilard. Experiments with the Chemostat on spontaneous mutations of bacteria. Proceedings of the National Academy of Sciences of the United States of America, 36(12):708–719, December 1950. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://adsabs.

harvard.edu/abs/1950PNAS...36..708N;   http://www.pnas.
org/content/36/12/708.full.pdf+html. Reprinted in [FS87,

pages 405–416].

Novick:1951:GMB

[NS51a] Aaron Novick and Leo Szilard. Genetic mechanisms in bacteria and bacterial viruses, I: Experiments on spontaneous and chemically induced mutations of bacteria growing in the Chemostat. Cold Spring Harbor Symposia on Quantitative Biology, 16(??):337–343, ???? 1951. CODEN CSHSAZ. ISSN 0091-7451 (print), 1943-4456 (electronic). Reprinted in [FS87, pages 418–424].

Novick:1951:VSI

[NS51b] Aaron Novick and Leo Szilard. Virus strains of identical phenotype but different genotype. Science, 113(2924):34–35, January 1951. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1951Sci...113.

..34N. Reprinted in [FS87, page 417].

Novick:1952:AM

[NS52] Aaron Novick and Leo Szilard. Anti-mutagens. Nature, 170(4335): 926–927, November 29, 1952. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v170/n4335/pdf/170926a0.pdf. Reprinted in

[FS87, pages 425–428].

Novick:1954:IEC

[NS54] Aaron Novick and Leo Szilard. II. Experiments with the Chemostat on the rates of amino acid synthesis in bacteria. In Kozloff et al. [KNSB54], pages 21–32. LCCN ???? Reprinted in [FS87, pages 429–440].

Naray-Szabo:1998:TTF

[NS98] Náray-Szabó Gábor. A tudós társadalmi felelőssége. (Hungarian) [The scientist’s social responsibility]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.hu/

fszemle/archivum/fsz9804/tart9804.html. Second special is-

sue on the centenary of Leo Szilard’s birth.

Ottaviani:2001:FJR

[O^{+}01] Jim Ottaviani et al. Fallout: J. Robert Oppenheimer, Leo Szilard, and the political science of the atomic bomb. G. T. Labs, Ann Arbor, MI, USA, 2001. ISBN 0-9660106-3-9. 239 pp. LCCN QC773.3.U5 O78 2001.

Olive:1962:SUP

[Oli62] Ralph Olive. Scientist urges policy changes to avert all-out war: ‘council for a liveable world’ suggested. Eugene Register-Guard, ??

(??):??, January 15, 1962. URL http://library.ucsd.edu/dc/
object/bb6554482m.

Olsen:1963:TGS

[Ols63] Arthur J. Olsen. Trackdown of the German scientist: Nazism and defeat scattered the leaders of a once-great scientific establishment. Herewith a review of where they went and what some of them did when they got there. New York Times, ??(??): 214, September 22, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/116463805/.

Osborn:1997:RCR

[Osb97] June Osborn. Report from the conference rapporteur. Health and Human Rights, 2(3):105–111, ???? 1997. URL http://www.jstor.

org/stable/4065161.

Ohlinger:1944:NEC

[OSY44] L. A. Ohlinger, Leo Szilard, and G. Young. New end closures for Al cans. Report N-866, US Atomic Energy Commission, Washington, DC, USA, March 30, 1944.

Paladi:2005:SMB

[Pal05] Vitalie Eremeev Florentin Paladi. A Szilard model-based computational study of the evolution of agents-clusters. Physica A, 348: 630–640, March 17, 2005. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0378437104012051.

Pallo:2010:BRB

[Pal10] Gábor Palló. Book review: Double exile. Migration of Jewish– Hungarian professionals through Germany to the United States, 1919–1945 by Tibor Frank. Studies in East European Thought, 62(2):241–243, June 2010. CODEN SEETE3. ISSN 0925-9392 (print), 1573-0948 (electronic). URL http://www.jstor.org/

stable/40927720.

Parrondo:2001:SER

[Par01] J. M. R. Parrondo. The Szilard engine revisited: Entropy, macroscopic randomness, and symmetry breaking phase transitions. Chaos (Woodbury, NY), 11(3):725–733, September 2001. CODEN

CHAOEH. ISSN 1054-1500. URL http://adsabs.harvard.edu/
abs/2001Chaos..11..725P.

Povh:1979:ESP

[PB79] I. L. Povh and A. D. Barinberg. From the Einstein–Szilard patent to modern magnetohydrodynamics. Impact of Science on Society, 29(1):49–60, ???? 1979. ISSN 0019-2872.

Pearson:1989:TLW

[Pea89] J. M. Pearson. Toward a livable world — Leo Szilard and the crusade for nuclear arms-control. Queens Quarterly, 96(3):756–758, Fall 1989. ISSN 0033-6041.

Peierls:1943:CRC

[Pei43] Rudolf Peierls. Comments on report A-387 (CF338) by Feld and Szilard. Report MS-75 (PRO AB4/913), Directorate of Tube Alloys, ????, UK, 1943.

Perlman:1962:PPE

[Per62a] David Perlman. ‘A people’s plan’ to end arms race. San Francisco Chronicle, ??(??):??, January 10, 1962.

Perlman:1962:NLP

[Per62b] David Perlman. A national lobby ‘a people’s plan’ to end arms race: Scientist Leo Szilard: A $20 million popular lobby is his goal. San Francisco Chronicle, ??(??):??, January 10, 1962. URL http:

//library.ucsd.edu/dc/object/bb6554482m.

Perutz:1993:IBB

[Per93] M. F. Perutz. ‘An intellectual bumblebee’: Genius in the Shadows: A Biography of Leo Szilard, The Man Behind the Bomb by William Lanouette, with Bela Silard, foreword by Jonas Salk. New York Review of Books, 40(16):17–20, October 7, 1993. ISSN 0028-7504 (print), 1944-7744 (electronic).

URL http://www.nybooks.com/articles/archives/1993/oct/
07/an-intellectual-bumblebee/.

Phillips:1980:SLH

[Phi80] M. Phillips. Leo Szilard — his version of the facts. Science & Society, 44(1):87–89, ???? 1980. ISSN 0036-8237.

Pallo:1988:LWH

[PL88] G. Pallo and W. Lanouette. William Lanouette on his Leo Szilard biography + interview. The New Hungarian Quarterly, 29(111): 160–170, Fall 1988. ISSN 0028-5390.

Platt:1958:LSH

[Pla58] John Platt. Leo Szilard and his ideas. Notes at the 1958 Conference on Biophysics, Boulder, Colorado, July 20–August 16, 1958., August 1958.

Pallo:1998:SAP

[PM98] Palló Gábor and Marx György. Szilárd átvette a parancsnokságot … — beszélgetés Zeisel (Striker) Évával és lányával. (Hungarian) [Szilard took command… — a chat with Eva Zeisel-Striker and her daughter]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9802/

tart9802.html. Special issue on the centenary of Leo Szilard’s birth.

Podvig:2008:LSL

[Pod08] P. Podvig. Leo Szilard Lectureship Award talk: Nuclear disarmament after the cold war. APS April Meeting Abstracts, pages 6005–

??, April 2008. URL http://adsabs.harvard.edu/abs/2008APS.

.APR.S6005P. American Physical Society, 2008 APS April Meeting and HEDP/HEDLA Meeting, April 11–15, 2008, abstract #S6.005.

Pohl:1962:ED

[Poh62] Frederik Pohl. The expert dreamers. Doubleday, Garden City, NY, USA, 1962. 248 pp. LCCN PZ1 .P745 Ex. Reprinted in 1968 by Avon Books, New York.

Polanyi:2019:WMD

[Pol19] John Polanyi. We must do more to prevent nuclear war. Times Higher Education, ??(??):??, December 9, 2019. URL https:/

/www.timeshighereducation.com/features/we-must-do-more-
prevent-nuclear-war.

Powers:1993:HWS

[Pow93] Thomas Powers. Heisenberg’s war: the secret history of the German bomb. Alfred A. Knopf, New York, NY, USA, 1993. ISBN 0-394-51411-4. xi + 607 pp. LCCN QC16.H35 P69 1993. US$27.50, CAN$34.50.

Price:1995:RDC

[Pri95] Matt Price. Roots of dissent: The Chicago Met Lab and the origins of the Franck Report. Isis, 86(2):222–244, June 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236323.

Prigge:2003:BBA

[Pri03] Walter Prigge, editor. Bauhaus, Brasilia, Auschwitz, Hiroshima: Weltkulturerbe des 20. Jahrhunderts: Modernität und Barbarei. (German) [Bauhaus, Brasilia, Auschwitz, Hiroshima: World Heritage Sites of the 20th Century: modernity and barbarism], volume 12 of Edition Bauhaus. Jovis-Verlag, Berlin, Germany, 2003. ISBN 3-936314-93-4. 319 pp. LCCN CB427 2003. URL http:

//hsozkult.geschichte.hu-berlin.de/rezensionen/2003-3-
159.

Pinkus:2002:API

[PT02] Binyamin Pinkus and Moshe Tlamim. Atomic power to Israel’s rescue: French–Israeli nuclear cooperation, 1949–1957. Israel Studies, 7(1):104–138, ???? 2002. ISSN 1084-9513 (print), 1527-201X

(electronic). URL http://www.jstor.org/stable/30246784.

Puck:1960:LSS

[Puc60] Theodore Thomas Puck. Leo Szilard and the science of the twentieth century. Humanist, 4(??):195–200, ???? 1960.

Rabinowitch:1946:BH

[Rab46] Eugene Rabinowitch. Before Hiroshima. Bulletin of the Atomic Scientists, 1(10):1, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:WM

[Rab47] Eugene I. Rabinowitch. Working for a miracle. Bulletin of the Atomic Scientists, 3(12):350, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Comment on [Szi47d].

Rabinowitch:1960:IBL

[Rab60] Eugene I. Rabinowitch. [introduction to How to Live with the Bomb — and Survive]. Bulletin of the Atomic Scientists, 16(2):58, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Szi60f].

Rabinowitch:1964:JFL

[Rab64] Eugene Rabinowitch. James Franck 1882–1964, Leo Szilard 1898– 1964. Bulletin of the Atomic Scientists, 20(8):16–20, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Radnai:1998:SLI

[Rad98] Radnai Gyula. Szilárd Leó iskolái. (Hungarian) [Leo Szilard’s secondary school years in Budapest]. Fizikai Szemle (Budapest), 48 (??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print),

1588-0540 (electronic). URL http://www.kfki.hu/fszemle/
archivum/fsz9802/tart9802.html. Special issue on the cente-

nary of Leo Szilard’s birth.

Raiteri:2002:FJR

[Rai02] S. Raiteri. Fallout: J. Robert Oppenheimer, Leo Szilard, and the political science of the atomic bomb. Library journal, 127(14):153, September 2002. CODEN LIBJA7. ISSN 0363-0277.

Reed:2015:ABS

[Ree15] Bruce Cameron Reed. The atomic bomb: the story of the Manhattan Project: how nuclear physics became a global geopolitical gamechanger. IOP concise physics. Morgan and Claypool Publishers and IOP Publishing, San Rafael, CA, USA and Bristol, UK, 2015. ISBN 1-62705-990-3 (print), 1-62705-991-1 (e-book), 1-62705-993-8 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). 239 (est.) pp. LCCN QC773.3.U5 R443 2015eb. URL http://iopscience.

iop.org/book/978-1-6270-5991-6.

Reingold:1984:MMA

[Rei84] Nathan Reingold. MGM meets the atomic bomb. The Wilson Quarterly (1976–2012), 8(4):154–163, Autumn 1984. ISSN 0363-

3276. URL http://www.jstor.org/stable/40256804.

Reiss:1995:FNC

[Rei95] Mitchell Reiss. The future that never came. The Wilson Quarterly (1976–2012), 19(2):46–48, 50–66, Spring 1995. ISSN 0363-3276.

URL http://www.jstor.org/stable/40258977.

Rex:2017:MDH

[Rex17] Andrew Rex. Maxwell’s Demon — a historical review. Entropy, 19 (6):240:1–240:13, May 2017. CODEN ENTRFG. ISSN 1099-4300.

Rider:1984:AOE

[Rid84] Robin E. Rider. Alarm and opportunity: Emigration of mathematicians and physicists to Britain and the United States, 1933– 1945. Historical Studies in the Physical Sciences, 15(1):107–176, ???? 1984. CODEN HSPSAS. ISSN 0073-2672. URL http://

www.jstor.org/stable/27757544.

Ripin:1996:SSI

[Rip96] Barrett H. Ripin. Szilard, Schawlow inducted into Inventor’s Hall of Fame. APS News, ??(??):??, November 1996. URL http://www.

aps.org/publications/apsnews/199611/inventors.cfm.

Ripin:1998:HFL

[Rip98] Barrett H. Ripin. Historical factal: Leo Szilard. APS News, ??(??):

??, February 1998. URL http://www.aps.org/publications/

apsnews/199802/szilard.cfm. The story begins: “Leo Szilard was born 100 years ago on February 11, 1898, in Budapest, Hungary, and emigrated to the US in 1938.”.

Roberts:196x:SOF

[Rob6x] Jack Roberts. A-scientist offers formula to avoid hot war. Reporter [Portland, Oregon], ??(??):??, ???? ??, 196x. URL http://

library.ucsd.edu/dc/object/bb6554482m.

Robbins:2014:RGP

[Rob14a] Gary Robbins. Revisiting the greatness of physicist Leo Szilard: UCSD talk set on nuclear pioneer, Salk visionary. U-T San Diego, ??(??):??, March 3, 2014. ISSN 1063-102X. URL http:/

/www.utsandiego.com/news/2014/mar/03/tp-revisiting-the-
greatness-of-physicist-leo/.

Robbins:2014:URG

[Rob14b] Gary Robbins. UCSD receives grant for physicist’s papers. U-T San Diego, ??(??):??, February 7, 2014. URL http:

//www.utsandiego.com/news/2014/feb/05/nuclear-physics-
szilard/.

Robbins:2014:URF

[Rob14c] Gary Robbins. UCSD remembers father of A-bomb. U-T San Diego, ??(??):??, February 25, 2014. URL http://www.

utsandiego.com/news/2014/feb/25/history-physics-ucsd/.

Robbins:2014:UDS

[Rob14d] Gary Robbins. UCSD to digitize Szilard papers. U-T San Diego, ??(??):??, February 5, 2014. URL http://www.utsandiego.com/

news/2014/feb/05/nuclear-physics-szilard/.

Rosenfeld:1964:WLS

[Ros64] Albert Rosenfeld. This was Leo Szilard: Remembrance of a genius. Life, ??(??):31, June 12, 1964.

Rosa:1998:AKA

[Rós98] Rósa Géza. Áttekintés a különböző atomerőműtípusokról. (Hungarian) [Various types of nuclear power plants — a survey]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://

www.kfki.hu/fszemle/archivum/fsz9802/tart9802.html. Spe-

cial issue on the centenary of Leo Szilard’s birth.

Anonymous:1947:LS

[Rot48] Anna Rothe. Leo szilard. In Anna Rothe, editor, Current Biography 1947, pages 622–625. H. W. Wilson, New York, NY,

USA, 1948. LCCN ???? URL http://library.ucsd.edu/dc/
object/bb6554482m;   http://library.ucsd.edu/dc/object/
bb7851423b.

Rotblat:1973:SFL

[Rot73] J. Rotblat. The sixteen faces of Leo Szilard. Nature, 242(5392): 67–68, March 1973. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://adsabs.harvard.edu/abs/
1973Natur.242...67R.

Rotblat:1998:SLP

[Rot98] Joseph Rotblat. Szilárd Leó, a Pugwash-mozgalom úttörője. (Hungarian) [Leo Szilard, a Pugwash movement pioneer]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.hu/

fszemle/archivum/fsz9804/tart9804.html. Second special is-

sue on the centenary of Leo Szilard’s birth.

Rupp:1931:BPE

[RS31] E. Rupp and L. Szilard. Beeinflussung ,,polarisierter” Elektronenstrahlen durch Magnetfelder. (German) [Influence on polarized electron beams by magnetic fields]. Naturwissenschaften, 19(??): 422–423, May 1931. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://adsabs.harvard.edu/abs/
1931NW.....19..422R.

Reid:1965:ESM

[RTS^{+}65] R. W. Reid, Richard Thomas, John Stockbridge, Robert Rietty, James Langahm, and Murray Kash. Einstein: The story of the man by his friends. 58-minute black-and-white film from BBC Television, London, UK., 1965. URL http://explore.bfi.org. uk/4ce2b6d1c3ab0. J. Robert Oppenheimer, Thomas Bucky, Helen Dukas, Patrick Gordon Walker, Banesh Hoffman, Otto Nathan, Leo Szilard, Edward Teller and A. R. Ubbelohde.

Reid:1969:ESM

[RTS^{+}69] R. W. Reid, Richard Thomas, John Stockbridge, Robert Rietty, James Langahm, and Murray Kash. Einstein: The story of the man by his friends. 42-minute black-and-white motion picture from Time-Life Films, Paramus, NJ, USA., 1969. J. Robert Oppenheimer, Thomas Bucky, Helen Dukas, Patrick Gordon Walker, Banesh Hoffman, Otto Nathan, Leo Szilard, Edward Teller and A. R. Ubbelohde.

Sakharov:1983:MG

[Sak83] Andrei Sakharov. A message from Gorky. Bulletin of the Atomic Scientists, 39(6):2–3, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salk:1964:LSI

[Sal64] Jonas Salk. Leo Szilard’s influence on physics. Available in the Leo Szilard Papers archive., June 13, 1964.

Sargent:1982:SLH

[Sar82] B. W. Sargent. Leo Szilard — his version of the facts. Queens Quarterly, 89(1):174–177, ???? 1982. ISSN 0033-6041.

Szilard:1947:ISF

[SBFA47] Leo Szilard, S. Bernstein, B. Feld, and J. Ashkin. Inelastic scattering of Fe, Pb, and Bi. Report MDDC-1536, US Atomic Energy Commission, Washington, DC, USA, August 27, 1947. Reprinted in [FS87, pages 323–327].

Szilard:1948:ISF

[SBFA48] L. Szilard, S. Bernstein, B. Feld, and J. Ashkin. Inelastic scattering of fast neutrons by Fe, Pb, and Bi. Physical Review, 73(11):1307–1310, June 1, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//adsabs.harvard.edu/abs/1948PhRv...73.1307S;   http://
prola.aps.org/abstract/PR/v73/i11/p1307_1.

Szilard:1960:SWSb

[SBSW60] Leo Szilard, Hans A. (Hans Albrecht) Bethe, Lewis L. Strauss, and Jerome B. (Jerome Bert) Wiesner, editors. Small world: Szilard, Bethe, Strauss, Wiesner [Motion picture]. Columbia Broadcasting System, ????, 1960. LCCN ???? 28 minute 16mm film.

Szilard:1934:CSR

[SC34a] Leo Szilard and T. A. Chalmers. Chemical separation of the radioactive element from its bombarded isotope in the Fermi effect. Nature, 134(3386):462, September 22, 1934. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http:

//adsabs.harvard.edu/abs/1934Natur.134..462S;   http:/
/www.nature.com/nature/journal/v134/n3386/pdf/134462b0.

pdf. Reprinted in [FS87, pages 143–144].

Szilard:1934:DNL

[SC34b] Leo Szilard and T. A. Chalmers. Detection of neutrons liberated from beryllium by gamma rays: a new technique for inducing radioactivity. Nature, 134(3887):494–495, September 29, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://adsabs.harvard.edu/abs/1934Natur.
134..494S;   http://www.nature.com/nature/journal/v134/

n3387/pdf/134494b0.pdf. Reprinted in [FS87, pages 145–146].

Szilard:1935:RIN

[SC35] Leo Szilard and T. A. Chalmers. Radioactivity induced by neutrons. Nature, 135(3403):98, January 19, 1935. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http:/

/www.nature.com/nature/journal/v135/n3403/pdf/135098b0.

pdf. Reprinted in [FS87, page 149].

Schweber:1993:VMB

[Sch93a] S. Schweber. Visionary and mobilizer. (book reviews: Genius in the Shadows. A Biography of Leo Szilard). Science, 261(5127):1461– 1462, September 1993. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://adsabs.harvard.edu/abs/
1993Sci...261.1461L. See [LS92].

Schweber:1993:BRB

[Sch93b] Silvan Schweber. Book review: Visionary and mobilizer: Genius in the Shadows, by William Lanouette with Bela Silard. A Biography of Leo Szilard. Science, 261(5127):1461–1462, September 10, 1993. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/stable/2881904.

Schaefer:2000:SPA

[Sch00a] Laura Atkinson Schaefer. Single pressure absorption heat pump analysis. Ph.D. thesis, Georgia Institute of Technology, Atlanta,

GA, USA, 2000. 176 pp. URL http://search.proquest.com/
docview/304619904.

Schwarzschild:2000:BHS

[Sch00b] Bertram Schwarzschild. Bohr–Heisenberg Symposium marks Broadway opening of Copenhagen. Physics Today, 53(5):51–52, May 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v53/i5/
p51/s1.

Schollmeyer:2005:MR

[Sch05] Josh Schollmeyer. Minority report. Bulletin of the Atomic Scientists, 61(1):38–39, January 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schils:2012:AE

[Sch12a] René Schils. Albert Einstein. In How James Watt invented the copier: forgotten inventions of our great scientists [Sch12b], pages 123–129. ISBN 1-4614-0859-8 (softcover), 1-4614-0860-1 (e-book).

LCCN T15 .S35513 2012. URL http://link.springer.com/
chapter/10.1007/978-1-4614-0860-4_20.

Schils:2012:HJW

[Sch12b] René Schils. How James Watt invented the copier: forgotten inventions of our great scientists. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-0859-8 (softcover), 1-4614-0860-1 (e-book). vii + 170 pp. LCCN T15 .S35513 2012.

Schwarcz:2015:RCH

[Sch15] Joe Schwarcz. The right chemistry: How the Manhattan Project and ‘the bomb’ came to be. The Montreal Gazette, ??(??):??, May

10, 2015. ISSN 0384-1294. URL http://montrealgazette.com/
technology/science/the-right-chemistry-how-the-manhattan-
project-and-the-bomb-came-to-be.

Scott:1971:PPW

[Sco71] Walter Scott. Personality parade: Who is Leo Szilard? Rocky Mountain News, ??(??):??, May 30, 1971. URL http://library.

ucsd.edu/dc/object/bb7168822g.

Scully:2007:DQP

[Scu07a] Robert J. Scully. The demon and the quantum: from the Pythagorean mystics to Maxwell’s demon and quantum mystery. Wiley-VCH, Weinheim, Germany, 2007. ISBN 3-527-40688-3 (paperback). viii + 271 pp. LCCN QC174.12 .S43 2007. with endnotes by Marlan O. (Marlan Orvil) Scully.

Scully:2007:MDS

[Scu07b] Robert J. Scully. Maxwell’s Demon and Szilard’s one atom engine: A heady mixture of entropy, information and consciousness. In The demon and the quantum: from the Pythagorean mystics to Maxwell’s demon and quantum mystery [Scu07a], chapter 5, pages 65–74. ISBN 3-527-40688-3 (paperback). LCCN QC174.12 .S43 2007. with endnotes by Marlan O. (Marlan Orvil) Scully.

Szilard:1928:EMF

[SE28] Leo Szilard and Albert Einstein. Electrodynamic movement of fluid metals particularly for refrigerating machines. British Patent No. 303,065 (filed December 24, 1928, issued May 26, 1930), December 24, 1928. Reprinted in [FS87, pages 540–542].

Szilard:1926:CBL

[SE45] Leo Szilard and Albert Einstein. [correspondence between Leo Szilard and Albert Einstein]. Personal and scientific correspondence between the authors., 1926–1945. URL http://library.ucsd.

edu/dc/object/bb24247539.

Shils:1997:PGI

[SE97] Edward Shils and Joseph Epstein. Portraits: a gallery of intellectuals. University of Chicago Press, Chicago, IL, USA, 1997. ISBN 0-226-75336-0 (hardcover), 0-226-75337-9 (paperback). 255 pp.

LCCN CT3990.A2 S55 1997. URL http://www.loc.gov/catdir/
description/uchi051/96033162.htm;   http://www.loc.gov/
catdir/toc/uchi051/96033162.htm.

Seed:1995:LSC

[See95] David Seed. Leo Szilard and the cause of rational disarmament. Nuclear Texts and Contexts, 11(??):15–26, ???? 1995. CODEN ???? ISSN ????

Seed:2003:HGW

[See03] David Seed. H. G. Wells and the liberating atom. Science Fiction Studies, 30(1):33–48, March 2003. ISSN 0091-7729 (print), 2327-

6207 (electronic). URL http://www.jstor.org/stable/4241139.

The author describes a meeting between H. G. Wells (author of The World Set Free) and Leo Szilard, who had read Wells’ book in 1932 and was impressed by it as a work of fiction. When Szilard thought of the possibility of a nuclear chain reaction in 1933, his memory of Wells’ book led him to apply immediately for a British patent on nuclear fission.

Segre:1970:EFPb

[Seg70] Emilio Segrè. Enrico Fermi: physicist. University of Chicago Press, Chicago, IL, USA, 1970. ISBN 0-226-74472-8. x + 276 pp. LCCN QC16.F46 S4.

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452(1):xv– xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June 1984.

Sen:2021:EFH

[Sen21] Paul Sen. Einstein’s fridge: how the difference between hot and cold explains the universe. Scribner, New York, NY, USA, 2021. ISBN 1-5011-8130-0 (hardcover), 1-5011-8132-7 (e-book). xii + 305 pp. LCCN QC311 .S5118 2021.

Szilard:1936:CNC

[SF44] Leo Szilard and Enrico Fermi. [correspondence on nuclear chain reactions], 1936–1944. URL http://library.ucsd.edu/dc/

object/bb6929916v.

Szilard:1943:NEF

[SFA^{+}43] Leo Szilard, B. Feld, J. Ashkin, S. Bernstein, E. Creutz, J. Kelsner, and R. Scalettar. Neutron emission in fission of U238. Report CF-1177, US Atomic Energy Commission, Washington, DC, USA, December 29, 1943. Reprinted in [FS87, pages 295–320].

Szent-Gyorgyi:1955:LTF

[SG55] Albert Szent-Gyorgyi. Letter to the Times: Frequency of wars. New York Times, ??(??):8E–8E, February 20, 1955. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/docview/113227846. Response to

[Szi55d].

Szilard:1955:CBA

[SH57] Leo Szilard and Werner Heisenberg. [correspondence between the authors]. Four letters in English and German in the Leo Szilard Papers archive., 1955–1957. URL http://library.ucsd.edu/dc/

object/bb3278005z.

Sherwin:1985:RHW

[She85] Martin J. Sherwin. Retrospectives: How well they meant. Bulletin of the Atomic Scientists, 41(7):9–15, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sheffield:1994:TSL

[She94] Roy Scott Sheffield. The tragic science of Leo Szilard. Ph.D. thesis, University of Florida, Gainesville, FL, USA, 1994. 281

pp. URL http://search.proquest.com/docview/304092770;
http://www.archive.org/details/tragicscienceofl00shef.

Szilard:1957:CBA

[SHG58] Leo Szilard, Otto Hahn, and Morton Grodzins. [correspondence between the authors]. Letters in the Leo Szilard Papers archive., 1957–1958.

Shils:1964:LSM

[Shi64] Edward Shils. Leo Szilard, a memoir. Encounter, ??(??):35–41,

December 1964. URL http://www.unz.org/Pub/Encounter;
http://www.unz.org/Pub/Encounter-1964dec-00035?View=PDF.

Silard:1998:SEH

[Sil98] John Silard. Személyes emlékeim. (Hungarian) [Personal recollections]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9802/

tart9802.html. Special issue on the centenary of Leo Szilard’s birth.

Szilard:1959:CIU

[SK63] Leo Szilard and Nikita S. Khrushchev. [correspondence on informal meetings of US and Soviet nuclear scientists]. Personal correspondence between Szilard and Soviet Premier Khrushchev, with Russian and English translations., 1959–1963. URL http://library.

ucsd.edu/dc/object/bb38582144.

Smith:1960:EDS

[Smi60] Alice Kimball Smith. The elusive Dr. Szilard. Harper’s Magazine, ??(??):77–86, July 1960. CODEN HAMAA3. ISSN 1045-7143.

Smith:1962:TPC

[Smi62] Donald O. Smith. Two per cent for peace. Solid State Design,

??(??):13–14, March 1962. URL http://library.ucsd.edu/dc/
object/bb20151957.

Smith:1965:AH

[Smi65a] Alice Kimball Smith. After Hiroshima. Chicago Today, 2(2):22–

29, Summer 1965. URL http://library.ucsd.edu/dc/object/
bb4609073r.

Smith:1965:PHS

[Smi65b] Alice Kimball Smith. A peril and a hope; the scientists’ movement in America, 1945–47. University of Chicago Press, Chicago, IL, USA, 1965. xiv + 591 pp. LCCN Q127.U6 S6.

Smith:1966:RAK

[Smi66] Alice Kimball Smith. Rabinowitch awarded Kalinga Prize. Science, 153(3744):1627, September 30, 1966. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://library.ucsd.

edu/dc/object/bb3312135f.

Smith:1970:PHS

[Smi70] Alice Kimball Smith. A peril and a hope; the scientists’ movement in America, 1945–47. MIT Press, Cambridge, MA, USA, 1970. ISBN 0-262-69026-8. xvi + 398 pp. LCCN Q127.U6 S6 1971.

Smith:2007:DMR

[Smi07] Peter D. (Peter Daniel) Smith. Doomsday men: the real Dr. Strangelove and the dream of the superweapon. St. Martin’s Press, New York, NY, USA, 2007. ISBN 0-312-37397-X. xxii + 552 pp. LCCN QC773.A1 S65 2007.

Szilard:1960:SWSa

[SMM^{+}60] Leo Szilard, Thomas Edward Murray, John B. (John Bruce) Medaris, Edward Teller, and Jerome B. (Jerome Bert) Wiesner, editors. Small world: Szilard, Murray, Medaris, Teller, Wiesner [Motion picture]. Columbia Broadcasting System, ????, 1960. LCCN ???? 28 minute 16mm film.

Szilard:1945:U

[Ss45] Leo Szilard and 64 scientists. [unknown]. Petition to the US President to forbid use of atomic weapons on Japan., 1945. Cited in [Shi64, page 40].

Sudoplatov:1995:STM

[SSSS95] Pavel Sudoplatov, Anatolii Pavlovich Sudoplatov, Jerrold L. Schecter, and Leona Schecter. Special tasks: the memoirs of an unwanted witness, a Soviet spymaster. Little, Brown and Co., Boston, MA, USA, updated edition, 1995. ISBN 0-316-82115-2. xxxi + 527 pp. LCCN JN6529.I6 S83 1995.

Szilard:1960:DDP

[ST60] Leo Szilard and Edward Teller. Debate: Is disarmament possible and desirable? NBC television broadcast., 1960.

Stachel:1999:EIS

[Sta99] John Stachel. Einstein and Infeld, seen through their correspondence. Acta Physica Polonica B, 30(10):2879–2908, October 1999. CODEN APOBBB. ISSN 0587-4254 (print), 1509-5770

(electronic). URL http://www.actaphys.uj.edu.pl/fulltext?

series=Reg&vol=30&page=2879. The appendix Scientific Writings of Leopold Infeld is an shortened version of the annotated bibliography in [Inf70].

Steiner:1975:BAS

[Ste75] A. Steiner. Baptism of atomic scientists. Bulletin of the Atomic Scientists, 31(2):21–28, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strauss:1963:MDa

[Str63] Lewis L. Strauss. Men and decisions. Doubleday, Garden City, NY, USA, 1963. viii + 468 + 16 pp. LCCN E741 .S78.

Szilard:1949:DSB

[STS^{+}49] Leo Szilard, The Franck Committee, Frederick Seitz, Hans Bethe, Harold C. Urey, Harrison Brown, Irving Langmuir, Vannevar Bush, and Leslie R. Groves. Did the Soviet bomb come sooner than expected? The atomic scientists: The wartime administrators. Bulletin of the Atomic Scientists, 5(10):262–264, October 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1968:R

[SW68] Leo Szilard and Kathleen R. Winsor. Reminiscences. Perspectives in American History, 2(??):??, ???? 1968. ISSN 0079-0990.

Seaborg:1998:SLH

[SW98] Glenn Seaborg and Bertram Wolfe. Szilárd Leó hozzájárulása a Világhoz: (Hungarian) [Szilard’s contribution to the World]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL

http://www.kfki.hu/fszemle/archivum/fsz9802/tart9802.html.

Special issue on the centenary of Leo Szilard’s birth.

Szilard:1959:JFE

[SWC59] Leo Szilard, Eugene P. Wigner, and Edward C. Creutz. Jacketed fuel elements for graphite moderated reactors. US Patent 2,886,503, May 12, 1959. Filed February 20, 1946, serial number 649,080.

Sweet:1993:BRU

[Swe93] William Sweet. Book reviews: Uncertainties: Heisenberg’s War: The Secret History of the German Bomb, by Thomas Powers; Uncertainty: The Life and Science of Werner Heisenberg, by David C. Cassidy. Bulletin of the Atomic Scientists, 49(7):50–52, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wei93].

Swing:1945:RSB

[Swi45] Raymond Swing. Raymond Swing’s broadcast[: History of atomic bomb]. Script for radio broadcast on station WMAL on ABC net-

work., September 7, 1945. URL http://library.ucsd.edu/dc/
object/bb0718253r.

Szilard:1980:LSH

[SWS80] Leo Szilard, Spencer R. Weart, and Gertrud Weiss Szilard. Leo Szilard, his version of the facts: selected recollections and correspondence, volume 1 of The Collected works of Leo Szilard. MIT Press, Cambridge, MA, USA, first paperback printing edition, 1980. ISBN 0-262-69070-5 (paperback). xxii + 244 pp. LCCN ???? URL

http://adsabs.harvard.edu/abs/1979lshv.book.....W.

Szilard:1939:IEF

[SZ39] Leo Szilard and Walter H. Zinn. Instantaneous emission of fast neutrons in the interaction of slow neutrons with uranium. Physical Review, 55(8):799–800, April 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//adsabs.harvard.edu/abs/1939PhRv...55..799S;   http:/
/prola.aps.org/abstract/PR/v55/i8/p799_1. Reprinted in

[FS87, pages 158–159].

Szilard:1941:PRI

[SZ41] Leo Szilard and W. H. Zinn. Preliminary report on inelastic collision of neutrons in uranium and other heavy elements. Report CP-285, US Atomic Energy Commission, Washington, DC, USA, December 12, 1941. Reprinted in [FS87, pages 266–275].

Szabadvary:1987:LSS

[Sza87] Ferenc Szabadvary. Leo Szilard’s studies at the Palatine Joseph Technical University of Budapest. Periodica Polytechnica, 31(187): 187–190, ???? 1987.

Szabadvary:1998:SLT

[Sza98] Szabadváry Ferenc. Szilárd Leó tanulmányai a budapesti József Nádor Műegyetemen. (Hungarian) [Leo Szilard’s studies at Budapest Technical University]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-

0540 (electronic). URL http://www.kfki.hu/fszemle/archivum/

fsz9802/tart9802.html. Special issue on the centenary of Leo Szilard’s birth.

Szilard:1922:TSG

[Szi22] L. Szilard. Über die thermodynamischen Schwankungserscheinungen. (German) [On thermodynamic fluctuation phenomena]. Ph.D. thesis, Freie Universität Berlin, Berlin, Germany, 1922. 45 pp. This dissertation was awarded the notation “exima”: the highest honor, and was published in [Szi25].

Szilard:1925:APT

[Szi25] L. Szilard. Über die Ausdehnung der phänomenologischen Thermodynamik auf die Schwankungserscheinungen. (German) [On the extension of the phenomenological thermodynamics to oscillation phenomena]. Zeitschrift für Physik, 32(1):753–788, December 1925. CODEN ZEPYAA. ISSN 0044-3328. URL http://adsabs.

harvard.edu/abs/1925ZPhy...32..753S. Publication of doctoral

thesis [Szi22]. English translation in [FS87, pages 70–102].

Szilard:1926:VGM

[Szi26] Leo Szilard. Verfahren zum Gießen von Metallen in Formen unter Anwendung elektrischer Ströme. (German) [A method for casting of metals into molds under application of electrical currents]. German Patent No. 476,812 (filed January 20, 1926, issued May 8, 1929), January 20, 1926. Reprinted in [FS87, pages 532–539].

Szilard:1928:BKG

[Szi28] Leo Szilard. Beschleunigung von Korpuskeln. (German) [Acceleration of corpuscles]. German patent application S.89 028 (filed December 17, 1928), December 17, 1928.

Szilard:1929:EET

[Szi29a] L. Szilard. Über die Entropieverminderung in einem thermodynamischen System bei Eingriffen intelligenter Wesen. (German) [On entropy reduction in a thermodynamic system by the intervention of intelligent beings]. Zeitschrift für Physik, 53(11–12):840–856, ???? 1929. CODEN ZEPYAA. ISSN 0044-3328. URL http:/

/adsabs.harvard.edu/abs/1929ZPhy...53..840S. Accepted as

Habilitationsschrift by the Universität Berlin. This paper contains Szilard’s classic analysis of Maxwell’s demon, and shows that the entropy of a unit of information is equal to k log 2, foreshadowing Claude Shannon’s famous 1948 work on “A Mathematical Theory of Communication”. See also the later English translations [Szi64h, Szi76a] and [FS87, pages 120–129].

Szilard:1929:DTa

[Szi29b] Leo Szilard. Discharge tube. US Patent 1,697,210, January 1, 1929. Filed April 20, 1925, serial number 24,575, and in Germany, September 3, 1924.

Szilard:1929:DTb

[Szi29c] Leo Szilard. Discharge tube. US Patent 1,715,874, June 4, 1929. File October 28, 1925, serial number 65,394, and in Germany November 5, 1924.

Szilard:1929:KGC

[Szi29d] Leo Szilard. Korpuskularstrahlrohre. (German) [Corpuscular beam tubes]. German patent application S.89 288 (filed January 5, 1929)., January 5, 1929. Reprinted in [FS87, pages 554–563].

Szilard:1934:AST

[Szi34a] Leo Szilard. Asynchronous and synchronous transformers for particles. British Patent Application 5730/34 (filed February 21, 1934), February 21, 1934. Reprinted in [FS87, pages 564–604].

Szilard:1934:MPI

[Szi34b] Leo Szilard. Memorandum of possible industrial applications arising out of a new branch of physics. Report ??, ????, ????, ???? 1934.

Szilard:1934:SED

[Szi34c] Leo Szilard. Suggested experiments for the detection of nuclear chain reactions and the liberation of nuclear energy. Report ??, New York University, New York, NY, USA, Summer 1934.

Szilard:1935:ARN

[Szi35a] Leo Szilard. Absorption of residual neutrons. Nature, 136(3450): 950–951, December 14, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v136/n3450/pdf/136950b0.pdf. Reprinted in

[FS87, pages 150–152].

Szilard:1935:PSN

[Szi35b] Leo Szilard. Provisional specification no. 7840 1934. improvements in or relating to the transmutation of chemical elements. British Patent 440,023., December 12, 1935. URL http://

nuclearhistory.wordpress.com/2011/02/28/szilards-chain-
reaction-patents-440023-630726-lodged-1934/. Patent ap-

plied for on March 12, 1934. Reprinted in [FS87, pages 622–638].

Szilard:1936:IRT

[Szi36] Leo Szilard. Improvements in or relating to transmutation of chemical elements. British Patent 630,726., March 30,

1936. URL http://nuclearhistory.wordpress.com/2011/02/
28/szilards-chain-reaction-patents-440023-630726-lodged-

1934/. Provisional Specification No. 19157 1934. Patent applied for (in numbers 19157/34 and 19721/34) on June 28 and July 4, 1934, but held secret until published on 28 September 1949. Reprinted in [FS87, pages 639–651].

Szilard:1939:ANT

[Szi39a] Leo Szilard. Apparatus for nuclear transmutation. US Patent application 263,017, filed 20 March 1939, but never granted., March 20, 1939. Reprinted in [FS87, pages 652–690].

Szilard:1939:MLJ

[Szi39b] Leo Szilard. Memorandum to Lyman J. Briggs: The possibility of a large-scale experiment in the immediate future. Reprinted in [FS87, pages 204–206]., October 26, 1939.

Szilard:1939:PPR

[Szi39c] Leo Szilard. Process of producing radio-active elements. US Patent 2,161,985., June 13, 1939. Filed March 11, 1935, serial number 10,500. Filed in Great Britain on 12 March 1934.

Szilard:1939:M

[Szi39d] Tekla Szilard. Memoirs. Translation by Anna Bettelheim from the original Hungarian. Unpublished manuscript held in the Bela Silard Papers archive, Pleasantville, New York., 1939.

Szilard:1941:CNU

[Szi41a] Leo Szilard. The capture of neutrons by uranium in the energy region of photo neutrons from radium–beryllium sources. Report, ????, December 6, 1941. Available in the Leo Szilard Papers archive.

Szilard:1941:CCR

[Szi41b] Leo Szilard. Control of the chain reaction. Report ??, ????, ????, July 19, 1941.

Szilard:1941:MCF

[Szi41c] Leo Szilard. Memorandum on the contribution of fast neutrons to the chain reaction in a uranium–carbon system [draft]. Technical Report ??, Columbia University, New York, NY, USA, October 28, 1941. Available in the Leo Szilard Papers archive.

Szilard:1941:MRQ

[Szi41d] Leo Szilard. Memorandum raising the question whether the action of explosive chain-reacting bodies can be based on an ‘expulsion’ method (Columbia University). Report A-56, US Atomic Energy Commission, Washington, DC, USA, October 21, 1941.

Szilard:1941:PRM

[Szi41e] Leo Szilard. Preliminary report on the melting of uranium powder. Report A-24 (CT-M), US Atomic Energy Commission, Washington, DC, USA, August 16, 1941. Reprinted in [FS87, pages 329–331].

Szilard:1941:SSE

[Szi41f] Leo Szilard. Suggestions for a search for element 94 in nature. Report A-45 (CN-G), US Atomic Energy Commission, Washington, DC, USA, September 29, 1941.

Szilard:1942:MPH

[Szi42a] Leo Szilard. Memorandum for Professor A. H. Compton. Technical report, ????, ????, January 18, 1942.

Szilard:1942:MCPa

[Szi42b] Leo Szilard. Memorandum on the cooling of the power plant. Report CE-146, US Atomic Energy Commission, Washington, DC, USA, June 24, 1942. Addendum to [Szi42d].

Szilard:1942:MCPb

[Szi42c] Leo Szilard. Memorandum on the cooling of the power plant. Report CE-150, US Atomic Energy Commission, Washington, DC, USA, June 29, 1942. Supplement to [Szi42d, Szi42b].

Szilard:1942:CPP

[Szi42d] Leo Szilard. On the cooling of the power plant. Report CE-130, US Atomic Energy Commission, Washington, DC, USA, June 15, 1942. See addendum [Szi42b] and supplement [Szi42c]. Reprinted in [FS87, pages 332–345], with addenda from reports C-146 (June 24, 1942) and Report C-150 (June 29, 1942).

Szilard:1942:PRD

[Szi42e] Leo Szilard. Progress report to Dr. Doan and Dr. Hilberry on the search for a site in the New York area. Technical report, ????, ????, January 20, 1942.

Szilard:1942:SMB

[Szi42f] Leo Szilard. Short memorandum on bismuth cooled power unit. Report CP-360, US Atomic Energy Commission, Washington, DC, USA, November 23, 1942. Reprinted in [FS87, pages 359–368].

Szilard:1942:UAP

[Szi42g] Leo Szilard. Uranium aggregates for power unit. Report CP-357, US Atomic Energy Commission, Washington, DC, USA, November 23, 1942.

Szilard:1942:WWU

[Szi42h] Leo Szilard. What’s wrong with us? Memorandum., September 1942. Cited in [Shi64, page 39] and [LS92, page 235]. The memo urges speedier action on the Manhattan Project, and warns about a post-war race for peace or war.

Szilard:1944:EPE

[Szi44a] Leo Szilard. Extrusion process for elimination of weld on slugs. Report N-1346, US Atomic Energy Commission, Washington, DC, USA, July 7, 1944.

Szilard:1944:PUT

[Szi44b] Leo Szilard. Proposal for use of thorium in poisoning slugs for W pile. Report N-962, US Atomic Energy Commission, Washington, DC, USA, April 14, 1944.

Szilard:1944:WMP

[Szi44c] Leo Szilard. Water moderated pile with P-9 core. Report N-1355, US Atomic Energy Commission, Washington, DC, USA, July 7, 1944.

Szilard:1945:ABP

[Szi45a] Leo Szilard. Atomic bombs and the postwar position of the United States. Unpublished memo to US President Franklin D. Roosevelt, who died before the memo could be delivered., March 1945. Reproduced in [WS78, pages 196–204], and with an introduction to the President by Albert Einstein reproduced in [LS92, pages 261–261].

Szilard:1945:ADP

[Szi45b] Leo Szilard. An attempt to define the platform for our conversations with members of the Senate and House of Representatives. Cited in [Smi65a, page 26]., September 7, 1945.

Szilard:1945:LMC

[Szi45c] Leo Szilard. Liquid metal cooled fast neutron breeder. Report UC-LS-60, ????, ????, March 6, 1945. Reprinted in [FS87, pages 369–375]. Some sources give the number as MUC-LS-60.

Szilard:1945:WTS

[Szi45d] Leo Szilard. We turned the switch. The Nation, 156(??):718–719, December 22, 1945. CODEN ???? ISSN 0027-8378. URL http:/

/www.thenation.com/.

Szilard:1946:CWA

[Szi46a] Leo Szilard. Can we avert an arms race by an inspection system? In Masters and Way [MW46a], pages 61–65. LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Szilard:1946:DCR

[Szi46b] Leo Szilard. Divergent chain reaction in systems composed of uranium and carbon. Report MDDC-446, US Atomic Energy Commission, Washington, DC, USA, November 21, 1946. Uranium Committee Report A-55, submitted to the Physical Review, 6 February 1940, but publication withheld at the author’s request until postwar declassification.

Szilard:1946:KVU

[Szi46c] Leo Szilard. Kan vi undgaa Kaprustning ved et Inspektionssystem? In Masters and Way [MW46b], page ?? LCCN ???? Danish translation of [MW46a] by Gudrun Frederiksen og Ebbe Rasmussen. Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46a].

Szilard:1946:MEF

[Szi46d] Leo Szilard. Market economy free from trade cycles. Cited in [LS92, page 320]; prepared for publication in the Collected Works, but dropped from publication., September 1946.

Szilard:1946:OPP

[Szi46e] Leo Szilard. Observations on the provision of the progress of sciences. Compares scientific training in various countries, endorses the proposed National Science Foundation (NSF), and proposes lifetime salaries for scientists., 1946.

Szilard:1947:ABP

[Szi47a] Leo Szilard. Atomic bombs and the postwar position of the United States in the world — 1945. Bulletin of the Atomic Scientists, 3 (12):351–353, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1947:CC

[Szi47b] Leo Szilard. Calling for a crusade. Bulletin of the Atomic Scientists, 3(4–5):102–106, 125, April/May 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1947:CED

[Szi47c] Leo Szilard. Comment to the Editors by Dr. Szilard. Bulletin of the Atomic Scientists, 3(12):350, 353, December 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Szi47d, Rab47].

Szilard:1947:LS

[Szi47d] Leo Szilard. Letter to Stalin. Bulletin of the Atomic Scientists, 3 (12):348–349, 376, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Rab47, Szi47c].

Szilard:1947:PHA

[Szi47e] Leo Szilard. A personal history of the atomic bomb. Round Table 604, University of Chicago, Chicago, IL, USA, September 25, 1947.

Szilard:1947:PIP

[Szi47f] Leo Szilard. The physicist invades politics. Saturday Review of Literature, ??(??):7–8, 31–34, May 3, 1947. URL http://www.

unz.org/Pub/SaturdayRev-1947may03-00007.

Szilard:1948:SMR

[Szi48] Leo Szilard. Science is my racket. September 1948.

Szilard:1949:AFS

[Szi49a] Leo Szilard. The AEC Fellowships: Shall we yield or fight? Bulletin of the Atomic Scientists, 5(6–7):177–178, June/July 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Szi63a].

Szilard:1949:MTW

[Szi49b] Leo Szilard. My trial as a war criminal. The University of Chicago Law Review, 17(1):79–86, Autumn 1949. URL http://lawreview.

uchicago.edu/issues.html.

Szilard:1949:RGC

[Szi49c] Leo Szilard. Report on ‘Grand Central Terminal’. Fairplay Publishers, New York, NY, USA, 1949. ???? pp. LCCN ????

Szilard:1949:SWF

[Szi49d] Leo Szilard. Shall we face the facts? Bulletin of the Atomic Scientists, 5(10):269–273, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1950:CWI

[Szi50a] Leo Szilard. Can we have international control of atomic energy? Bulletin of the Atomic Scientists, 6(1):9–12, 16, January 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1950:DDD

[Szi50b] Leo Szilard. The diary of Dr. Davis. Bulletin of the Atomic Scientists, 6(2):51–57, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1950:RLC

[Szi50c] Leo Szilard. A reply to Lilienthal’s criticism. In Eliot [Eli50], page ?? LCCN UF767 .H2. Introduction by Albert Einstein and commentary by George Fielding Eliot.

Szilard:1950:SE

[Szi50d] Leo Szilard. Security and equality. Available in the Leo Szilard Papers archive., October 28, 1950.

Szilard:1951:PPM

[Szi51] Leo Szilard. Process for producing microbial metabolites. US Patent application 264,263, filed 29 December 1951 but never granted., December 29, 1951.

Szilard:1952:CCP

[Szi52a] Leo Szilard. Caffeine containing products and method for their preparation. US Patent application 320,816, filed 15 November 1952 but never granted., November 15, 1952.

Szilard:1952:YDW

[Szi52b] Leo Szilard. You do not want war with Russia? Unpublished typescript, Box 34, Folder 19, Leo Szilard Papers archive., 1952.

Szilard:1953:EML

[Szi53a] Lajos Szilard. Erinnerungen aus meinem Leben und dem Leben meiner angestammten Familie. (German) [Memories of my life and the lives of my ancestral family]. Possibly the original German typescript of Lajos’s memoirs, with many handwritten corrections, written in Yonkers, NY., August 1953. URL http://library.

ucsd.edu/dc/object/bb6895783f.

Szilard:1953:R

[Szi53b] Lajos [Louis] Szilard. Recollections. English translation of [Szi53a], with hand-drawn cover page. Translated by Bela Silard. Unpublished manuscript held in the Leo Szilard Papers archive., August

1953. URL http://library.ucsd.edu/dc/object/bb43019011.

Szilard:1953:MM

[Szi53c] Leo Szilard. Meeting of the minds. Several drafts of this paper from June 1953 to 26 January 1954 are in the Leo Szilard Papers archive., June 1953.

Szilard:1954:NME

[Szi54a] Lajos Szilard. Nachtrag zu meinen “Erinnerungen aus meinem Leben”, geschrieben im Jahre 1954 vom 1–6 Februar. (German) [Addendum to my “Memories of my life”, written in 1954 from 1 to 6 February]. The last page ends with “Ende. Yonkers den 16.

III. 1954.”, February 1954. URL http://library.ucsd.edu/dc/
object/bb5598841t.

Szilard:1954:SR

[Szi54b] Leo Szilard. Security risk. Bulletin of the Atomic Scientists, 10(10): 384–386, 398, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1955:DPP

[Szi55a] Leo Szilard. Disarmament and the problem of peace. Bulletin of the Atomic Scientists, 11(8):297–307, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty. See comment [Fow55].

Szilard:1955:FSP

[Szi55b] Leo Szilard. The first step to peace. Bulletin of the Atomic Scientists, 11(3):104, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1955:LE

[Szi55c] Leo Szilard. Letter to the editor. Bulletin of the Research Exchange on the Prevention of War, ??(??):39–??, March 1955. URL http:/

/library.ucsd.edu/dc/object/bb80220728. This is a reprint of

portions of a letter from Szilard to the New York Times [Szi55d]. The Leo Szilard archive records exchanges between the editor and Szilard, who wanted the original published in full.

Szilard:1955:LTA

[Szi55d] Leo Szilard. Letters to the Times: Action to prevent war: Sponsorship of qualified group to formulate agreements proposed. New York Times, ??(??):E8, February 6, 1955. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/113235031. See responses [Arn55, SG55]

and partial reprint [Szi55c].

Szilard:1955:PSA

[Szi55e] Leo Szilard. Process and apparatus for slowing the aging of man and similar large mammalian organisms. Unpublished manuscript, available in Box 39, Folder 4, of the Leo Szilard Papers archive., July 29, 1955.

Szilard:1957:MUR

[Szi57a] Leo Szilard. Method for unloading reactors. US Patent 2,778,792, January 22, 1957. Filed April 19, 1946.

Szilard:1957:MIO

[Szi57b] Leo Szilard. Method of intermittently operating a neutronic reactor. US Patent 2,796,396, June 18, 1957. Filed April 16, 1946, serial number 662,512.

Szilard:1958:HWM

[Szi58a] Leo Szilard. Heavy water moderated neutronic reactor. US Patent 2,832,733., April 29, 1958. Filed April 23, 1946. Serial number 664,145.

Szilard:1958:NRF

[Szi58b] Leo Szilard. Neutronic reactor and fuel element therefor. US Patent 2,825,689., March 4, 1958. Filed April 25, 1946. Serial number 664,732.

Szilard:1959:NAP

[Szi59a] Leo Szilard. On the nature of the aging process. Proceedings of the National Academy of Sciences of the United States of America, 45(1):30–45, January 1959. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.pnas.

org/content/45/1/30.full.pdf+html. Reprinted in [FS87, pages

447–462].

Szilard:1959:RWA

[Szi59b] Leo Szilard. Report on the work of the Academy’s Operating Committee on World Security Problems raised by nuclear weapons for the year 1958–1959. Technical report, National Academy of Arts and Sciences, ????, May 11, 1959. Available in the Leo Szilard Papers archive.

Szilard:1959:TA

[Szi59c] Leo Szilard. A theory of ageing. Nature, 184(4691):957–958, September 26, 1959. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v184/n4691/pdf/184957a0.pdf. Reprinted in [FS87,

pages 463–468].

Szilard:1960:CFS

[Szi60a] L. Szilard. The control of the formation of specific proteins in bacteria and in animal cells. Proceedings of the National Academy of Sciences of the United States of America, 46(3):277–292, March 1960. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1960PNAS...46.
.277S;   http://www.pnas.org/content/46/3/277.full.pdf+

html. Reprinted in [FS87, pages 469–484].

Szilard:1960:MBA

[Szi60b] L. Szilard. The molecular basis of antibody formation. Proceedings of the National Academy of Sciences of the United States of America, 46(3):293–302, March 1960. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://adsabs.

harvard.edu/abs/1960PNAS...46..293S;   http://www.pnas.
org/content/46/3/293.full.pdf+html. Reprinted in [FS87,

pages 485–494].

Szilard:1960:ABV

[Szi60c] Leo Szilard. Adaptation of The Voice of the Dolphins — leaving out the dolphins. Outline for a play, TV show, or movie, available in the Leo Szilard Papers archive., August 7, 1960.

Szilard:1960:DSR

[Szi60d] Leo Szilard. Dependence of the sex ratio at birth on the age of the father. Nature, 186(4725):649–650, May 21, 1960. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v186/

n4725/pdf/186649a0.pdf. Reprinted in [FS87, pages 495–498].

Szilard:1960:TCA

[Szi60e] Leo Szilard. Has the time come to abrogate war? Submitted to Look magazine and Foreign Affairs, but may never have been published. Box 25, Folder 25, in the Leo Szilard Papers archive., 1960.

Szilard:1960:HLB

[Szi60f] Leo Szilard. How to live with the bomb and survive — the possibility of a Pax Russo–Americana in the long-range rocket stage of the so-called atomic stalemate. Bulletin of the Atomic Scientists, 16(2):59–73, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Szi63d].

Szilard:1960:LEB

[Szi60g] Leo Szilard. Letter to the Editor: The Berlin crisis. Bulletin of the Atomic Scientists, 16(5):144, 193, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1960:MER

[Szi60h] Leo Szilard. Memorandum on enzyme repression. May 1960.

Szilard:1960:PER

[Szi60i] Leo Szilard. The postwar events and the Russian disarmament proposals of 1960. Available in the Leo Szilard Papers archive., 1960.

Szilard:1960:SS

[Szi60j] Leo Szilard. To stop or not to stop. Bulletin of the Atomic Scientists, 16(3):82–84, 108, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Szi63h].

Szilard:1961:HSP

[Szi61a] Leo Szilard. How to secure the peace in a disarmed world. Available in Box 26, Folder 14, Leo Szilard Papers archive., 1961.

Szilard:1961:IMM

[Szi61b] Leo Szilard. Induction of mutations in mammals by ionizing radiation. Report ??, ????, ????, March 10, 1961. Listed, without source, on [FS87, pages 12–13].

Szilard:1961:LEb

[Szi61c] Leo Szilard. Letter to the Editor. Washington Post, ??(??):??, April 21, 1961. ISSN 0190-8286. Protest at US invasion of Cuba in the Bay of Pigs.

Szilard:1961:LEc

[Szi61d] Leo Szilard. Letter to the Editor. New York Herald Tribune, ?? (??):??, April 21, 1961. Protest at US invasion of Cuba in the Bay of Pigs.

Szilard:1961:LEd

[Szi61e] Leo Szilard. Letter to the Editor. St. Louis Post-Dispatch, ??(??): ??, April 21, 1961. Protest at US invasion of Cuba in the Bay of Pigs.

Szilard:1961:LEA

[Szi61f] Leo Szilard. Letter to the Editor: American Society of Newspaper Editors and President Kennedy. Bulletin of the Atomic Scientists, 17(5–6):168, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1961:MD

[Szi61g] Leo Szilard. Memo on disarmament. Available in the Leo Szilard Papers archive., July 18, 1961.

Szilard:1961:MC

[Szi61h] Leo Szilard. The mined cities. Bulletin of the Atomic Scientists, 17 (10):407–412, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1961:PSE

[Szi61i] Leo Szilard. Political settlement in Europe. Available in Box 30, Folder 13, Leo Szilard Papers archive. Draft of a paper for the Pugwash meeting in Stowe, VT, USA., 1961.

Szilard:1961:VDO

[Szi61j] Leo Szilard. The voice of the dolphins, and other stories. Simon and Schuster, New York, NY, USA, 1961. 122 pp. LCCN PZ4.S9986 Vo.

Szilard:1962:WRW

[Szi62a] Leo Szilard. Are we on the road to war? Bulletin of the Atomic Scientists, 18(4):23–30, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1962:CD

[Szi62b] Leo Szilard. The Council’s dilemma. Bulletin of the Atomic Scientists, 18(10):50, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1962:HSM

[Szi62c] Leo Szilard. Homologous suppression in man. 1962.

Szilard:1962:VDA

[Szi62d] Leo Szilard. La voce dei delfini e altri racconti. (Italian) [The Voice of the Dolphins and other stories]. Feltrinelli, Milano, Italy, 1962. 161 pp. LCCN ???? Translated by Carlo Alberto Gastecchi.

Szilard:1962:VDF

[Szi62e] Leo Szilard. La voix des dauphins. (French) [The Voice of the Dolphins]. Denoël, Paris, France, 1962. 176 pp. LCCN ???? Translated by Philippe Chardeyron.

Szilard:1962:MGF

[Szi62f] Leo Szilard. Mark Gable Foundation. In The expert dreamers [Poh62], page ?? LCCN PZ1 .P745 Ex. Reprinted in 1968 by Avon Books, New York.

Szilard:1962:ODP

[Szi62g] Leo Szilard. On the occasional dominance of the ‘perceptible phenotype’ in man. 1962.

Szilard:1962:SCM

[Szi62h] Leo Szilard. The sex chromatin in mammalian cells, “dosage compensation” in the fruit fly, and the enzyme repression in bacteria. Report ??, ????, ????, 1962. Listed, without source, on [FS87, page 13].

Szilard:1963:AFS

[Szi63a] Leo Szilard. The AEC Fellowships: Shall we yield or fight? In Grodzins and Rabinowitch [GR63], pages 410–413. LCCN D842 .B78. Reprint of [Szi49a].

Szilard:1963:APC

[Szi63b] Leo Szilard. The aging process and the “competitive strength” of spermatozoa. Report ??, ????, ????, July 25, 1963. Listed, without source, on [FS87, pages 13–14].

Szilard:1963:SDU

[Szi63c] Leo Szilard. Die Stimme der Delphine: utopische Erzählungen. (German) [The Voice of the Dolphins: Utopian Narrative]. Rowohlt, Reinbek bei Hamburg, Germany, 1963. 135 pp. LCCN ????

Szilard:1963:HLB

[Szi63d] Leo Szilard. How to live with the bomb — and survive. In Grodzins and Rabinowitch [GR63], pages 217–243. LCCN D842 .B78. Reprint of [Szi60f].

Szilard:1963:LLS

[Szi63e] Leo Szilard. Letter to Lewis L. Strauss on 25 January 1939. In Men and decisions [Str63], pages 172–173. LCCN E741 .S78.

Szilard:1963:ODP

[Szi63f] Leo Szilard. On the occasional dominance of the “perceptible phenotype” in man. Report ??, ????, ????, July 12, 1963. Listed, without source, on [FS87, pages 13–14]. Revised version dated 18 May 1964.

Szilard:1963:PPU

[Szi63g] Leo Szilard. A petition to the President of the United States. In Grodzins and Rabinowitch [GR63], pages 28–29. LCCN D842 .B78. Original version 17 July 1945.

Szilard:1963:TT

[Szi63h] Leo Szilard. To test or not to test. In Grodzins and Rabinowitch [GR63], pages 342–347. LCCN D842 .B78. Reprint of [Szi60j].

Szilard:1963:WAM

[Szi63i] Leo Szilard. Why America may come to grief. March 27, 1963.

Szilard:1964:CVI

[Szi64a] Leo Szilard. Curriculum vitae: including list of publications. The document contains several typewritten and mimeographed versions of the CV, but they are vaguely dated. The last handwritten correction is dated 1964. There are no data for patents, or for magazine and newspaper articles., 1964. URL http://library.ucsd.edu/

dc/object/bb82268527?counter=36.

Szilard:1964:DSD

[Szi64b] Leo Szilard. Delfinernes stemme. (Danish) [Voice of the dolphins]. Steen Hasselbalch, Copenhagen, Denmark, 1964. 156 pp. LCCN ???? Danish translation by Michael Tejn of [Szi61j].

Szilard:1964:MAF

[Szi64c] Leo Szilard. Memorandum on antibody formation. 1964.

Szilard:1964:MIB

[Szi64d] Leo Szilard. Memorandum on the induction of B galactosidase. 1964.

Szilard:1964:MDV

[Szi64e] Leo Szilard. “Minimal deterrent” vs. saturation parity. Bulletin of the Atomic Scientists, 20(3):6–12, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1964:MBL

[Szi64f] Leo Szilard. The molecular basis of long-term memory. March 3, 1964.

Szilard:1964:MR

[Szi64g] Leo Szilard. On memory and recall. Proceedings of the National Academy of Sciences of the United States of America, 51 (6):1092–1099, June 15, 1964. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.org/

stable/72122;   http://www.pnas.org/content/51/6/1092.

full.pdf+html. Reprinted in [FS87, pages 497–504].

Szilard:1964:DET

[Szi64h] Leo Szilard. On the decrease of entropy in a thermodynamic system by the intervention of intelligent beings. Behavioral Science, 9(4): 301–310, ???? 1964. CODEN BEHSAS. ISSN 0005-7940 (print),

1932-300X (electronic). URL http://fab.cba.mit.edu/classes/
863.18/notes/computation/Szilard-1929.pdf. Translated by

Anatol Rapoport and Mechthilde Knoller from the 1929 German original. Translation revised by Cark Eckart.

Szilard:1965:SBS

[Szi65] Leo Szilard. The “sting of the bee” in saturation parity. Bulletin of the Atomic Scientists, 21(3):8–13, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1969:R

[Szi69] Leo Szilard. Reminiscences. In Fleming and Bailyn [FB69], pages 94–151. ISBN 0-674-33411-6, 0-674-33412-4. LCCN E169.1 .F6; E171 .P4 v.2.

Szilard:1971:CLC

[Szi71] Gertrud Weiss Szilard. Communications: Lunar craters. Bulletin of the Atomic Scientists, 27(4):4, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1972:CIS

[Szi72] Leo Szilard. Creative intelligence and society: The case of atomic research, the background in fundamental science. In Feld and Szilard [FS87], pages 178–189. ISBN 0-262-06039-6 (vol. 1: 1972), 0-262-19168-7 (vol. 2: 1978), 0-262-19260-8 (vol. 3: 1987). LCCN QC3 .S97. Public lecture, University of Chicago, July 31, 1946.

Szilard:1976:ERT

[Szi76a] L. Szilard. On entropy reduction in a thermodynamic system by interference by intelligent subjects. Zhurnal Physik, 53:840–856, February 1976. CODEN ???? ISSN ???? URL http://adsabs.

harvard.edu/abs/1976ZhPhy..53..840S. Translation into En-

glish from Zeitschrift für Phys. (West Germany), v. 53, 1929 p. 840–856 [Szi29a].

Szilard:1976:PCL

[Szi76b] Leo Szilard. Petition circulated by Leo Szilard to oppose U.S. use of atomic bomb against Japan, July 17, 1945. In Energy and conflict: the life and times of Edward Teller [BO76], pages 459–460. ISBN 0-399-11551-X. LCCN QC16.T37 B58 1976. US$12.95. URL http:

//alsos.wlu.edu/information.aspx?id=367.

Szilard:1978:ABP

[Szi78a] Leo Szilard. Atomic bombs and the postwar position of the United States in the world. In Weart and Szilard [WS78], pages 196– 204. ISBN 0-262-19168-7 (hardcover). LCCN QC3 .S97 vol. 2; QC16.S95.

Szilard:1978:MPI

[Szi78b] Leo Szilard. Memorandum of possible industrial applications arising on a new branch of physics. In Weart and Szilard [WS78], pages 39–?? ISBN 0-262-19168-7 (hardcover). LCCN QC3 .S97 vol. 2; QC16.S95. The memo, written on 28 June 1934, notes “The production of energy [from artificial radioactivity] and its use for power production would be possible on such a large scale and probably with so little cost that a sort of industrial revolution could be expected.”.

Szilard:1979:SLC

[Szi79] L. Szilard. Some links in the chain. New Scientist, 82(1157):738– 740, ???? 1979. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Szilard:1981:SDS

[Szi81] Leo Szilard. Die Stimme der Delphine: Science-fiction-Erzählungen. (German) [The Voice of the Dolphins: science fiction narrative], volume 53 of Suhrkamp-Taschenbuch. Suhrkamp, Frankfurt am Main, Germany, 1981. ISBN 3-518-37203-3. 112 pp. LCCN ????

Szilard:1982:RGC

[Szi82] Leo Szilard. Report on Grand Central Terminal. In Laughing space: funny science fiction [AJ82], page ?? ISBN 0-395-30519-5. LCCN PN6231.S42 L38.

Szilard:1985:CSC

[Szi85] Leo Szilard. La coscienza si chiama Hiroshima: dossier sulla bomba atomica. (Italian) [Consciousness is called Hiroshima: file on the atomic bomb]. Editori riuniti, Roma, Italy, 1985. ISBN 88-359-2812-5. x + 278 pp. LCCN ???? Edited by Spencer Weart and Gertrud Wiess Szilard. Preface by Carlo Bernardini.

Szilard:1987:DSA

[Szi87] Leo Szilard. Draft of a statement about Edward Teller (August 23, 1963). In Toward a livable world: Leo Szilard and the crusade for nuclear arms control [HGS87], pages 405–406. ISBN 0-262-19260-8, 0-262-08162-8. LCCN QC3 .S97 vol. 3.

Szilard:1992:VDO

[Szi92] Leo Szilard. The voice of the dolphins and other stories. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, expanded edition, 1992. ISBN 0-8047-1753-2, 0-8047-1754-0. vi + 182 pp. LCCN PS3569.Z5 V65 1992. US$20.00, US$8.95.

Szilard:1998:SMA

[Szi98a] Leo Szilard. The sensitive minority among men of science. In Marx [Mar98b], pages 176–184. LCCN ???? Dinner lecture at Brandeis University, Los Angeles, December 8, 1954.

Szilard:1998:EKT

[Szi98b] Szilárd Leó. Érzékeny kisebbség a tudóstársadalomban. (Hungarian) [The sensitive minority among men of science]. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.hu/

fszemle/archivum/fsz9804/tart9804.html. Second special is-

sue on the centenary of Leo Szilard’s birth.

Szilard:1998:LSH

[Szi98c] Szilárd Leó. Levél Sztálinhoz. (Hungarian) [Letter to Stalin]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL

http://www.kfki.hu/fszemle/archivum/fsz9802/tart9802.html.

Special issue on the centenary of Leo Szilard’s birth.

Szilard:2003:GCT

[Szi03] Leo Szilard. Grand central terminal: rapporto da un pianeta estinto. (Italian) [Grand Central Terminal: report from an extinct planet]. Orecchio acerbo, Roma, Italy, 2003. ISBN 88-900693-9-2. ???? pp. LCCN ????

Szilard:2004:VDI

[Szi04] Leo Szilard. La voce dei delfini. (Italian) [The Voice of the Dolphins]. L’ancora del mediterraneo, Napoli, Italy, 2004. ISBN 88-8325-145-8. 127 pp. LCCN ????

Szilard:2005:DHH

[Szi05] Szilárd Leó. A delfinek hangja. (Hungarian) [The voice of dolphins]. Kairosz, Budapest, Hungary, 2005. ISBN 963-9568-95-3. 124 pp. LCCN PS3569.Z5 V6515 2005.

Szilard:2007:CWA

[Szi07] Leo Szilard. Can we avert an arms race by an inspection system? In Masters and Way [MW07], page ?? ISBN 1-59558-227-4 (hardcover). LCCN UG1282.A8 O54 2007. URL http:

//thenewpress.com/index.php?option=com_title&task=view_
title&metaproductid=1703;   http://www.loc.gov/catdir/
toc/ecip0718/2007020838.htm. Foreword by Niels Bohr. Intro-

duction by Arthur H. Compton. Reprint of [MW46a].

Geszti:1998:SID

[Tam98] Geszti Tamás. Szilárd és az információ démona. (Hungarian) [Szilard and the demon of information]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.hu/

fszemle/archivum/fsz9802/tart9802.html. Special issue on the

centenary of Leo Szilard’s birth.

Teller:1983:SHR

[Tel83] Edward Teller. Seven hours of reminiscences. Los Alamos Science, 7:190–195, Winter/Spring 1983. CODEN LASCDI. ISSN

0273-7116. URL http://library.lanl.gov/cgi-bin/getfile?
07-22.pdf.

Telegdi:1998:SIA

[Tel98a] Valentine L. Telegdi. Szilard as an inventor: accelerators and more. APS April Meeting Abstracts, ??(??):402–??, April

1998. URL http://adsabs.harvard.edu/abs/1998APS..APR..

D402T. American Physical Society, APS/AAPT Joint April Meeting, April 18-21, 1998 Columbus, Ohio, abstract #D4.02.

Teller:1998:RMH

[Tel98b] Teller Ede. Rövid megemlékezés. (Hungarian) [A brief commemoration]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9802/

tart9802.html. Special issue on the centenary of Leo Szilard’s birth.

Telegdi:2000:SIA

[Tel00] V. L. Telegdi. Szilard as inventor: Accelerators and more. Physics Today, 53(10):25–28, October 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://adsabs.

harvard.edu/abs/2000PhT....53j..25T.

Telegdi:2002:EFA

[Tel02] Valentine L. Telegdi. Enrico Fermi in America. Physics Today, 55(6):38–43, June 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v55/i6/p38_s1.

Teresi:1993:GSB

[Ter93] D. Teresi. Genius in the shadows — a biography of Leo Szilard, the man behind the bomb. New York Times Book Review, ??(??): 1, January 24, 1993. ISSN 0028-7806.

Thompson:2004:TML

[Tho04] P. M. Thompson. Thomas Merton and Leo Szilard: The parallel paths of a monk and a nuclear physicist. Zygon Journal of Religion and Science, 39(4):979–986, December 2004. ISSN 0591-2385 (print), 1467-9744 (electronic).

Tibell:1998:AFL

[Tib98] Gunnar Tibell. Arvet från Leo Szilard: hur står det till med den etiska reflexionen i dagens forskning? (Swedish) [The heritage of Leo Szilard: ethical reflections on today’s research]. Vår lösen: kristen kulturtidskrift, ??(??):116–120, ???? 1998. ISSN 0346-4679.

Turner:1946:AE

[Tur46] Louis A. Turner. Atomic energy from U^{238}. Physical Review, 69(7–8):366, April 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/doi/

10.1103/PhysRev.69.366.

Turchetti:2006:SNS

[Tur06] Simone Turchetti. “For slow neutrons, slow pay”: Enrico Fermi’s patent and the U.S. Atomic Energy Program, 1938–1953. Isis, 97 (1):1–27, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.
1086/501097;   http://www.jstor.org/stable/pdfplus/10.
1086/501097.pdf.

Valdevit:2006:OFS

[Val06] Giampaolo Valdevit. Oppenheimer fra scienza e potere. una storia americana. (Italian) [Oppenheimer between science and power. An American story]. Studi Storici, 47(1):115–142, January/March 2006. ISSN 0039-3037 (print), 2036-458X (electronic). URL http:

//www.jstor.org/stable/20567340.

VanDeMark:2003:PKN

[Van03] Brian VanDeMark. Pandora’s keepers: nine men and the atomic bomb. Little, Brown and Co., Boston, MA, USA, 2003. ISBN 0-316-73833-6. xii + 399 + 16 pp. LCCN QC774.A2 V36 2003. URL

http://www.loc.gov/catdir/toc/fy045/2002043646.html.

Villani:2015:RLQ

[VB15] Cédric Villani and Edmond Baudoin. Les rêveurs lunaires: quatre génies qui ont changé l’histoire. (French) [The lunar dreamers: four geniuses who changed history]. Gallimard, Paris, France, 2015. ISBN 2-07-066593-3 (paperback). 179 + 12 pp. LCCN ????

vonBorstel:1996:MAA

[vBDL96] R. C. von Borstel, J. W. Drake, and L. H. Loeb. Mechanisms of antimutagenesis and anticarcinogenesis. Mutation Research, 350 (1):294, 1996. CODEN MUREAV. ISSN 0027-5107 (print), 1873-135X (electronic). Special issue: mechanisms of antimutagenesis and carcinogenesis. Dedicated to Aaron Novik and the memory of Leo Szilard (1898–1964), the discoverers of antimutagenesis.

Vizi:1998:SC

[Viz98] Vizi E. Szilveszter. Science and conscience. Fizikai Szemle (Budapest), 48(??):??, April 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.hu/

fszemle/archivum/fsz9804/tart9804.html. Second special is-

sue on the centenary of Leo Szilard’s birth.

Vaikuntanathan:2011:MMDa

[VJ11] Suriyanarayanan Vaikuntanathan and Christopher Jarzynski. Modeling Maxwell’s demon with a microcanonical Szilard engine. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 83(6 (part 1)):061120, June 15, 2011. CO-DEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic).

Volodin:1963:FPR

[Vol63] Y. Volodin. The foreign press in review: A scientist’s responsibility (u.s. atomic scientists’ journal on war and peace). International Affairs (Royal Institute of International Affairs 1944–), ??(??):78– 81, May ??, 1963. ISSN 0020-5850 (print), 1468-2346 (electronic).

URL http://library.ucsd.edu/dc/object/bb6554482m.

Wachhorst:2006:MRB

[Wac06] Wyn Wachhorst. Movie review: Ten Days That Unexpectedly Changed America: Einstein’s Letter, Directed and produced by Barak Goodman and John Maggio. History Channel, 2006. 60 mins. (A&E Home Video, P.O. Box 2284, South Burlington, VT 05407; 888-423-1212; http://www.store.aetv.com/). The Journal of American History, 93(3):983–984, 2006. ISSN 0021-8723 (print),

1945-2314 (electronic). URL http://jah.oxfordjournals.org/
content/93/3/983.short.

Walker:1997:PUD

[Wal97] J. Samuel Walker. Prompt and utter destruction: Truman and the use of atomic bombs against Japan. University of North Carolina Press, Chapel Hill, NC, USA, 1997. ISBN 0-8078-2361-9 (hardcover), 0-8078-4662-7 (paperback), 0-8078-6618-0 (e-book). xiii + 142 pp. LCCN D767.25.H6 W355 1997. URL http://www.h-

net.org/review/hrev-a0b3w6-aa;   http://www.jstor.org/
stable/1408325; http://www.jstor.org/stable/25054041;
http://www.jstor.org/stable/27551824;   http://www.jstor.
org/stable/3642279.

Wald:2004:FLP

[Wal04] Chelsea Wald. Focus: Landmarks: The Physical Review’s explosive secret. Physical Review Focus, 14(??):17–??, ???? 2004. CO-DEN PRFHAQ. ISSN 1539-0748. URL http://physics.aps.

org/story/print/v14/st17.

Wattenberg:1993:BNA

[Wat93] Albert Wattenberg. The birth of the nuclear age. Physics Today, 46(1):44–51, January 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/pdf/10.1063/1.881378.

Waugh:1962:SSS

[Wau62] John C. Waugh. Szilard stumps for strong peace movement. The Christian Science Monitor, ??(??):??, January 24, 1962. ISSN 0882-7729 (print), 1540-4617 (electronic). URL http://library.ucsd.

edu/dc/object/bb6554482m.

Weart:1976:SS

[Wea76] Spencer R. Weart. Scientists with a secret. Physics Today, 29 (2):23–30, February 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v29/i2/p23_s1. Reprinted in [WP85,

pages 123–129].

Wehrwein:1961:SWF

[Weh61] Austin C. Wehrwein. Scientist would form council to lobby for abolishing war: Szilard wants groupp to have a full-time staff for pressure on capital. New York Times, ??(??):??, December 3, 1961. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

URL http://library.ucsd.edu/dc/object/bb6554482m.

Weidenthal:1961:FBB

[Wei61] Bud Weidenthal. First A-bomb builder here for peace talk. Cleveland Press and News, ??(??):??, November 29, 1961. URL http:/

/library.ucsd.edu/dc/object/bb6554482m.

Weinberg:1974:BRB

[Wei74] Alvin M. Weinberg. Book review: The Collected Works of Leo Szilard: Scientific Papers, edited by Bernard T. Feld and Gertrude Weiss Szilard. American Journal of Physics, 42(3):261–263, March 1974. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://adsabs.harvard.edu/abs/1974AmJPh.
.42..261F;   http://ajp.aapt.org/resource/1/ajpias/v42/
i3/p261_s1.

Weinberg:1979:NES

[Wei79] Alvin M. Weinberg. Nuclear energy: Salvaging the atomic age. The Wilson Quarterly (1976–2012), 3(3):88–112, Summer 1979. ISSN

0363-3276. URL http://www.jstor.org/stable/40255664.

Weisskopf:1993:LGH

[Wei93] Victor Weisskopf. Letter: Giving Heisenberg his due. Bulletin of the Atomic Scientists, 49(9):53, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Swe93].

Weiner:1994:LFD

[Wei94] Nella Fermi Weiner. Letter: Fermi: Didn’t even tell his wife. Bulletin of the Atomic Scientists, 50(4):3, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1998:SLB

[Wei98] Alvin M. Weinberg. Szilárd Leó, a bomba és a világbéke jövője. (Hungarian) [Leo Szilard, the bomb and the future of world peace]. Fizikai Szemle (Budapest), 48(??):??, February 1998. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://www.kfki.hu/fszemle/archivum/fsz9802/

tart9802.html. Special issue on the centenary of Leo Szilard’s birth.

Winterberg:1996:MHN

[WHF^{+}96] Friedwardt Winterberg, Günter Herrmann, Igor Fodor, Lincoln Wolfenstein, and Mark E. Singer. More on how Nazi Germany failed to develop the atomic bomb. Physics Today, 49(1):11–83, January 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wiesner:1965:WSP

[Wie65] Jerome B. (Jerome Bert) Wiesner. Where Science and Politics Meet. McGraw-Hill, New York, NY, USA, 1965. viii + 302 pp. LCCN Q127.U6 W5. URL http://calteches.

library.caltech.edu/2395/1/books.pdf;   http://library.
ucsd.edu/dc/object/bb6554482m. Pages 176–179 and 226–227

discuss Leo Szilard’s proposals for nuclear disarmament.

Wigner:1969:LSB

[Wig69] E. P. Wigner. Leo Szilard 1898–1964: A biographical memoir. Biographical memoirs — National Academy of Sciences of the United States of America, 40:335–347, 1969. CODEN BMNSAC.

ISSN 0077-2933. URL http://books.nap.edu/html/biomems/
lszilard.pdf;   http://library.ucsd.edu/dc/object/bb6827526c.

Wigner:1992:SLA

[Wig92] Wigner Jenő. Szilárd Leó [Angolból] ford. Tóth Eszter. (Hungarian) [Leo Szilard [English] rpm. Eszter Tóth]. Fizikai Szemle (Budapest), 42(11):406–407, ???? 1992. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1996:LS

[Wig96a] E. P. Wigner. Leo Szilard (1898–1964). In Wigner [Wig96b], pages 139–149. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CWE

[Wig96b] Eugene Paul Wigner, editor. The collected works of Eugene Paul Wigner. Part A. The scientific papers. Vol. II. Nuclear physics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-56972-3. x + 574 pp. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Winkler:1998:RUM

[Win98] A. M. Winkler. Realism, utopia, and the mushroom cloud: Four activist intellectuals and their strategies for peace, 1945–1989; Louise Weiss (France), Leo Szilard (USA), E. P. Thompson (England), Danilo Dolci (Italy). The American Historical Review, 103(2):488, April 1998. ISSN 0002-8762 (print), 1937-5239 (electronic).

Wolff:1967:TFY

[Wol67] Anthony Wolff. Twenty-five years with the bomb: A conversation with Nobel Prize-winner Eugene P. Wigner. Look Magazine, ?? (??):58, 60–61, December 26, 1967. URL http://library.ucsd.

edu/dc/object/bb5564714p.

Wood:1962:COS

[Woo62] William B. Wood. Campus opinion: Szilard suggests a new plan of political action. The Stanford Daily, ??(??):??, January 16, 1962.

URL http://library.ucsd.edu/dc/object/bb6554482m.

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Weart:1978:LSH

[WS78] Spencer R. Weart and Gertrud Weiss Szilard, editors. Leo Szilard, his version of the facts: selected recollections and correspondence, volume 2 of The Collected works of Leo Szilard. MIT Press, Cambridge, MA, USA, 1978. ISBN 0-262-19168-7 (hardcover). xxii + 244 pp. LCCN QC3 .S97 vol. 2; QC16.S95.

Wigner:1992:REP

[WS92] Eugene Paul Wigner and Andrew Szanton. The recollections of Eugene P. Wigner as told to Andrew Szanton. Plenum Press, New York, NY, USA; London, UK, 1992. ISBN 0-306-44326-0. xxiv + 335 pp. LCCN QC16.W52 A3 1992.

Wigner:2003:REP

[WS03] Eugene Paul Wigner and Andrew Szanton. The recollections of Eugene P. Wigner as told to Andrew Szanton. Basic Books, New York, NY, USA, 2003. ISBN 0-7382-0886-8. xxiv + 335 pp. LCCN QC16.W52.

Wigner:1959:JFE

[WSC59] Eugene P. Wigner, Leo Szilard, and Edward C. Creutz. Jacketed fuel element. US Patent 2,872,401., February 3, 1959. Filed May 8, 1946, Serial number 668,110.

Wigner:1961:MLI

[WSCF61] Eugene P. Wigner, Leo Szilard, Robert F. Christy, and Francis Lee Friedman. Massive leakage irradiator. US Patent 2,986,510, May 30, 1961. Filed May 14, 1946, serial number 669,524.

Szilard:1942:ABC

[WWCS42] A. M. Weinberg, E. P. Wigner, R. F. Christy, and Leo Szilard. Approximate boundary conditions for diffusion equation at interface between two media. Report CP-189, US Atomic Energy Commission, Washington, DC, USA, July 10, 1942. Reprinted in [FS87, pages 288–291].

Yavenditti:1978:ASH

[Yav78] Michael J. Yavenditti. Atomic scientists and Hollywood: The beginning or the end? Film and History, 8(4):??, December 1978.

York:1975:SAT

[Yor75] Herbert F. York. Sounders of the alarm: Thirty years of attempts to bring the world to grips with the perpetual menace to human security. Bulletin of the Atomic Scientists, 31(10):43–45, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zipser:1969:APP

[ZB69] David Zipser and Jonathan Beckwith, editors. Abstracts of Papers Presented at the Lactose Operon Meeting, Cold Spring Harbor, New York, September 1–5, 1969. Cold Spring Harbor Laboratory of Quantitative Biology, Cold Spring Harbor, NY 11724,

1969. URL http://library.ucsd.edu/dc/object/bb18104153/
_1.pdf/download.

Zetterberg:2009:SHN

[ZBZ^{+}09] Henrik Zetterberg, Magnus Bath, Madeleine Zetterberg, Peter Bernhardt, and Ola Hammarsten. The Szilard hypothesis on the nature of aging revisited. Genetics, 182(1):3–9, May 2009. CODEN GENTAE. ISSN 0016-6731 (print), 1943-2631 (electronic).

Zeisel:1970:CS

[Zei70] Hans Zeisel. Communications: On Szilard. Bulletin of the Atomic Scientists, 26(7):41, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zinn:1939:ENU

[ZS39] W. H. Zinn and Leo Szilard. Emission of neutrons by uranium. Physical Review, 56(7):619–624, October 1, 1939. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v56/i7/

p619_1. Reprinted in [FS87, pages 163–168].

Zsadanyi:1976:PWJ

[Zsa76] Zsadanyi Oszkar. Párbeszéd Wigner Jenő Nobel-díjas professzorral. (Hungarian) [Dialogue with Nobel laureate Professor Eugene Wigner]. ????, ??(??):??, September 1, 1976. URL http://

library.ucsd.edu/dc/object/bb7851423b.

Zuckerman:1988:NWB

[Zuc88] Lord Zuckerman. Nuclear wizards: Rabi: Scientist and Citizen by John S. Rigden. Toward a Livable World: Leo Szilard and the Crusade for Nuclear Arms Control edited by Helen S. Hawkins, edited by G. Allen Greb, edited by Gertrud Weiss Szilard. Better a Shield than a Sword: Perspectives on Defense and Technology by Edward Teller. New York Review of Books, 35(5):26–31, March 31, 1988. ISSN 0028-7504 (print), 1944-7744 (electronic).

URL http://www.nybooks.com/articles/archives/1988/mar/
31/nuclear-wizards/.

Zurek:2003:MDS

[Zur03] Wojciech Hubert Zurek. Maxwell’s Demon, Szilard’s Engine and quantum measurements. eprint arXiv:quant-ph/0301076, January

2003. URL http://adsabs.harvard.edu/abs/2003quant.ph..
1076Z.